### v2

# GRAND: Graph Release with Assured Node Differential Privacy

Abstract
Differential privacy is a well-established framework for safeguarding sensitive information in data. While extensively applied across various domains, its application to network data — particularly at the node level — remains underexplored. Existing methods for node-level privacy either focus exclusively on query-based approaches, which restrict output to pre-specified network statistics, or fail to preserve key structural properties of the network. In this work, we propose GRAND (Graph Release with Assured Node Differential privacy), which is, to the best of our knowledge, the first network release mechanism that releases entire networks while ensuring node-level differential privacy and preserving structural properties. Under a broad class of latent space models, we show that the released network asymptotically follows the same distribution as the original network. The effectiveness of the approach is evaluated through extensive experiments on both synthetic and real-world datasets.
1 Introduction
Rapid technological advances and explosive social connectivity have made complex networks more common than ever. From the Internet to transportation topologies for traffic planning and international trade agreements, networks are widely used to depict intensive interactions within complex systems. However, nestled within these marvels of modernity are layers of privacy concerns that cannot be overlooked. For example, in studies of organ donation networks, many of the donors and recipients are not willing to disclose their identities (Hizo-Abes et al., 2010; Marcus et al., 2023). In studying financial system crisis, firms are generally reluctant to share records of their transactions with other firms. Similarly, in studying the sexual relationship between HIV patients, the identity of subjects has to be protected (Little et al., 2014; Abadie et al., 2021).
These concerns manifest at various levels – ranging from personal specter of private information leakage, at the macroscopic level, affecting entire donation or financial relations with the potential to precipitate systemic crises. At the forefront are safeguarding or preventive actions that intend to avoid potential risks preemptively, which focus on the fortification of complex network data against privacy breaches. In comparison to conventional (non-network) data-sharing scenarios, enhancing privacy within complex networks presents unique difficulties. These mainly arise from network intricate structures, the prevalent noise, and the vast scale of associated issues, which complicate the use of simple and well-established mathematical models for privacy protection (Karwa et al., 2017; Hehir et al., 2022). Further, algorithms are usually applied to large-scale networks, including social networks on the Internet or computer clusters; these complex networks carry higher requirements regarding algorithms’ computational and communication efficiency (Guo et al., 2023).
Differential privacy (DP; (Dwork, 2006)) has recently become a common practice for privacy protection. It intends to enforce strict privacy protection for all individuals in a dataset, including those with extreme values that are the most vulnerable to privacy leakages and deanonymization. A variety of privacy-preserving mechanisms have been proposed to guarantee differential privacy, such as the Laplace mechanism (Dwork et al., 2006, 2014) and the exponential mechanism (McSherry and Talwar, 2007). These methods have been widely applied in various learning tasks (Lei et al., 2018; Soto et al., 2022; Lin et al., 2023; Xue et al., 2024; Ma et al., 2025). And variations of differential privacy have also been thoroughly studied, including relaxed or approximate differential privacy (Abadi et al., 2016; Cai et al., 2019), local differential privacy (Rohde and Steinberger, 2018), random differential privacy (Hall et al., 2012), Kullback-Leibler privacy (Wang et al., 2016) and Gaussian differential privacy (Dong et al., 2022). Meanwhile, differential privacy has been applied across many fields, both in academia and across industries (Kaissis et al., 2020; Hie et al., 2018; Han et al., 2020; Santos-Lozada et al., 2020; Kenthapadi and Tran, 2018; Nayak, 2020).
Differential privacy in complex network data analysis has also been carefully explored (e.g., Karwa and Slavković, 2016; Rohde and Steinberger, 2018; Chang et al., 2024). However, as network data typically show structures distinct from that of Euclidean data, the preservation of differential privacy for network data entails unique challenges. First, differential privacy preservation may come at the cost of altering a network’s underlying utilities. This includes, but not limited to, crucial network properties, such as degree distributions, centrality and potentially community structures. Many existing methods may inadvertently alter the underlying distribution of edges and degrees. Meanwhile, edge-level modifications, such as edge addition or deletion, may also change node-level properties. In order to preserve a certain network property, therefore, many existing methods focus on statistic release rather than network release – that is, the release of certain summary statistics (e.g., degree distribution or the number of triangles) with privacy guarantee, rather than an entire differentially private network that has all properties preserved.
Second, the adoption of differential privacy may appear at different levels. On the one hand, many existing works consider the concept of edge differential privacy (Nissim et al., 2007), which protects privacy of individual edges. The adoption of edge differential privacy is natural and shows good theoretical properties (Mülle et al., 2015; Karwa et al., 2017; Fan et al., 2020; Hehir et al., 2022; Chang et al., 2024). On the other hand, the investigation and development of node differential privacy (Hay et al., 2009) is relatively rare. By definition, node differential privacy guarantees that a network (or its certain properties) should remain robust with the change of any node. However, the alteration of a node usually entails the alteration of all associated edges (and potentially other information). The preservation of network properties given node differential privacy is therefore considered substantially more challenging (Kasiviswanathan et al., 2013; Hehir et al., 2022). In particular, the release of entire networks that follow node differential privacy has rarely been investigated in previous literature.
In this work, we propose a novel mechanism, GRAND (Graph Release with Assured Node Differential privacy), to release networks that guarantees node differential privacy. Moreover, the proposed method is shown to preserve statistical network properties under the general latent space models, which include many widely used statistical network models as special cases. To the best of our knowledge, this is the first method that can release entire networks (rather than a summary statistic) with node differential privacy that is computationally feasible with utility preservation guarantees.
The rest of this article is organized as follows. Section 2 reviews the background of differential privacy in network data, identifies gaps in the existing literature, and outlines our primary contributions. Section 3 formally defines node differential privacy and also the statistical model class that motivates our design of the GRAND mechanism. Section 4 presents a detailed description of our method. The theoretical properties of GRAND is studied in Section 5. In Section 6, we evaluate GRAND’s ability to preserve network properties through extensive simulation examples. Section 7 further showcases the application of our method in privatizing real-world social networks. We conclude the paper with additional discussions in Section 8.
2 Differential Privacy of Network Data and Our Contributions
Node differential privacy vs. edge differential privacy
Broadly speaking, existing works on network differential privacy can be categorized into two branches (Abawajy et al., 2016; Li et al., 2023c). One branch is edge differential privacy (EDP; Nissim et al., 2007). Intuitively, an EDP mechanism ensures that one cannot infer the pairwise relation between any two nodes in the network. Many methods have been proposed and discussed regarding the protection of EDP, which protect the privacy of individual edges. This can be achieved through mechanisms such as randomized responses (Mülle et al., 2015; Hehir et al., 2022), method of moments (Chang et al., 2024), or noise injection (Karwa and Slavković, 2016; Fan et al., 2020). The other branch is node differential privacy (NDP; Hay et al., 2009). An NDP mechanism (to be formally defined later) ensures that one cannot infer an individual node in the network.
Notice that, information associated with a node includes its pairwise relations to other nodes, and hence node differential privacy is more strict than edge differential privacy. As a result, node-level privacy protection is considerably more challenging than edge-level privacy. Applying edge-level DP methods to node-level settings usually leads to degenerate results (Kasiviswanathan et al., 2013; Hehir et al., 2022; Chang et al., 2024).
However, we advocate that preserving node differential privacy can be critical in many scenarios. This is because nodes can represent individual users, customers, and patients, whose identities are sensitive and vulnerable to de-anonymization. Encoding node differential privacy, while preserving the underlying network’s properties is the key motivation of this paper.
Network release vs. statistic release
The adoption of differential privacy in network analysis, either at the node level or at the edge level, involves introducing random noise, which may lead to unintended consequences. First, existing differentially private mechanisms that are designed for Euclidean data may inadvertently change the underlying structure or properties of networks. For example, adding or deleting edges may lead to the alteration of the degree distribution, transitivity, and the number of components. In general, releasing a network that is differentially private while preserving all original properties can be very challenging.
One viable approach is statistic release. That is, rather than releasing a network, one releases a particular network summary statistic that satisfies differential privacy, while preserving the accuracy of this particular statistic. This line of works include, but not limited to, the release of triangle counts (Liu et al., 2022), node degrees (Sivasubramaniam et al., 2020; Yan, 2021; Wang et al., 2022), and network centralities (Laeuchli et al., 2022) with the edge differential privacy guarantee, and the release of triangle counts (Blocki et al., 2013), the number of connected components (Kalemaj et al., 2023; Jain et al., 2024), edge density (Ullman and Sealfon, 2019), and degree distributions (Day et al., 2016; Raskhodnikova and Smith, 2016; Macwan and Patel, 2018) with the node differential privacy guarantee.
Nevertheless, statistic release has many critical drawbacks. For example, one has to design different methods for different statistics, and each such task can be sophisticated. Moreover, for many commonly used network statistics, such as centralities and closeness, no release mechanisms are currently available. More importantly, statistic release completely eliminates the possibility of analyzing multiple network statistics simultaneously. This is because when each statistic is processed separately, the dependence between statistics would be lost. And it has been widely known that, given the complexity of network data, one usually needs inference of multiple network statistics jointly for informative analysis (Qi et al., 2024).
With the aforementioned limitations, it is then evident that directly releasing a network would be substantially more flexible for subsequent use. Ideally, if the released network is almost the same as the original network, known as an informative release (Wasserman and Zhou, 2010), we can use the privatized network in an arbitrary way in downstream tasks. In particular, when the released network has most network properties maintained the same, it essentially subsumes most statistic release. Along this path, many have explored the novel ways of generating synthetic networks under edge differential privacy such as Karwa et al. (2017); Qin et al. (2017); Hehir et al. (2022); Guo et al. (2023); Chang et al. (2024). In contrast, for node differential privacy, the only known methods in the literature (Borgs et al., 2015, 2018; Chen et al., 2024) are NP-hard, thus are only of theoretical interest.
Our contributions
The preceding discussion motivates the need for a mechanism to release entire networks (rather than statistics) with node-level differential privacy, which is expected to be a crucial tool in practice. while such a tool is not yet available in the literature. In this paper, we will introduce a computationally feasible method to release a network under node differential privacy, while preserving key network properties. To the best of our knowledge, we are the first to achieve this goal in the literature.
3 Node Differential Privacy and Latent Space Models
3.1 Notations
For any positive integer $K$, we denote the set $\{1,\cdots,K\}$ by $[K]$. In particular, we use $[N]$ to index the nodes in the network. Consider a network of $N$ individuals (nodes), where the edges represent interactions or relationships between nodes, such as friendship, collaboration, or transactions (Harris, 2009; Ji and Jin, 2016; Newman, 2018). We assume the network is unweighted and undirected. The network can be uniquely represented by an $N\times N$ adjacency matrix $A$ where $A_{ij}=1$ if individuals $i$ and $j$ are connected, and $A_{ij}=0$ otherwise. Denote the node and edge sets of network $A$ by $\mathcal{V}(A)$ and $\mathcal{E}(A)$, respectively. [^1] The proposed work can be applied to weighted or directed networks as well, as long as a general latent space model defined in Section 3.3 is applicable to the corresponding adjacency matrices. Details will be provided in Section 4.2. For any vector $z\in{\bf R}^{d}$, we use $\|{z}\|$ to denote its Euclidean norm.
3.2 Node Differential Privacy
Suppose that the network information is sensitive (e.g., a cancer patient network) and many nodes in $A$ may not be willing to release their identity or even their presence in the network. Private information may still be vulnerable to leakage if one simply releases the network to the public, even with de-anonymization (Narayanan and Shmatikov, 2008). In many cases, one may infer a node’s identity according to certain connectivity patterns of a de-anonymized network, such as when one individual is connected with all the rest of the individuals. Alternatively, a node’s identity may be revealed due to adversarial attacks.
We adopt differential privacy (Dwork, 2006) as the standard for privacy protection. One commonly used definition of is $\varepsilon$-differential privacy, where a parameter $\varepsilon$, referred to as the privacy budget, is prespecified, such that a small $\varepsilon$ indicates a tight budget, associated with a stringent privacy protection protocol. Technically, $\varepsilon$-differential privacy requires that the alteration of any entry in the original dataset only leads to a small change of the output data’s distribution. This is quantified by the ratio of distributions before and after the alteration being bounded by $e^{\varepsilon}$.
On network data, however, the definition of one data entry may be construed differently (e.g., as a node or as an edge). There have been multiple ways to define differential privacy on network data, and a detailed review can be found in Jiang et al. (2021). Specifically, in this paper, we focus on node differential privacy (NDP), formally defined as below.
Definition 1 (Node differential privacy).
Let $\varepsilon>0$ be the privacy budget. A network data releasing mechanism $\mathcal{M}$ satisfies node $\varepsilon$-differential privacy if for any measurable set $\Psi$ of the network sample space, we have
$\mathbb{P}(\mathcal{M}(A)\in\Psi)\leq e^{\varepsilon}P(\mathcal{M}(A^{\prime})\in\Psi)$ (1)
for any two networks $A$ and $A^{\prime}$ with adjacency node sets $\mathcal{V}(A)$ and $\mathcal{V}(A^{\prime})$: $A$ and $A^{\prime}$ are only different in one row and the corresponding column.
Analogously, the edge level differential privacy is defined by requiring (1) for any two networks $A$ and $A^{\prime}$ with only one difference in $\mathcal{E}(A)$ and $\mathcal{E}(A^{\prime})$. The above definitions have been also discussed in several previous frameworks (Hay et al., 2009; Kasiviswanathan et al., 2013; Karwa and Slavković, 2016; Imola et al., 2021; Hehir et al., 2022; Guo et al., 2023; Chang et al., 2024). Intuitively, one should not be able to infer an individual’s identity from the released network data under the definition of node level differential privacy. In contrast, edge level privacy refers to protection against inferring the existence of an edge. It is not difficult to see that node-level DP provides strictly stronger privacy guarantees than edge-level DP (Kasiviswanathan et al., 2013). Technically, for an individual $i\in\mathcal{V}(A)$, NDP protects $i$’s relationship with all other individuals $j$’s in the network, $j\in\mathcal{V}(A)$, $j\neq i$. In other words, the output $\mathcal{M}(A)$ should remain “similar” when an entire row and column of $A$ has been changed (to an arbitrary degree). Our goal is to design a network-releasing mechanism $\mathcal{M}$ satisfying the NDP definition above. Meanwhile, the released network $\mathcal{M}(A)$ would provably preserve network properties of the original $A$ under the privacy budget.
3.3 General latent space models
We now proceed to introduce a class of statistical models for network data, referred to as the latent space model. We first introduce the design of our GRAND mechanism based on this model. And later we will demonstrate that our NDP guarantee still holds without the model.
The idea of latent space has been widely used in random networks, and was first formally introduced by Hoff et al. (2002). In this work, we adopt a slightly more general form than that in Hoff et al. (2002), though the model class is defined based on the same idea.
Definition 2 (General latent space model).
We say that $A$ is a network generated from the general latent space model if there exists a distribution $F$ (unknown) on $\mathbb{R}^{d}$ and a known symmetric generative function $W:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1]$ such that $A$ can be generated through the procedure below:
$Z_{1},...,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(W(Z_{i},Z_{j})),\quad i>j.$
Here $Z_{1},...,Z_{N}$ are latent vectors corresponding to nodes $1,\ldots,N$, respectively, and $d$ denotes the dimension of the latent space. The above model encapsulates a series of popular models as special cases. We illustrate a few examples here.
Example 1 (Inner product latent space model (Hoff et al., 2002)).
We say that $A$ is a network generated from an inner-product latent space model if there exist distributions $F_{X}$ on $\mathbb{R}^{d}$ and $F_{\alpha}$ on $\mathbb{R}$, such that $A$ is generated as follows:
$(X_{1},\alpha_{1}),...,(X_{N},\alpha_{N})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F_{X}\times F_{\alpha},\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))),\quad i>j$
where
$W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))=\sigma\big{(}X_{i}^{T}X_{j}+\alpha_{i}+\alpha_{j}\big{)}$ (2)
and $\sigma$ is the sigmoid function. Taking $Z_{i}=(X_{i},\alpha_{i})$, this model is a special case of the general latent space model.
If one is willing to introduce additional constraint bounds on the parameter space, the model can be further simplified as below.
Example 2 (Random dot product graph (RDPG) (Young and Scheinerman, 2007)).
We say $A$ is a network generated from a generalized random dot product model if there exists a distribution $F$ on $\mathbb{R}^{d}$, such that $A$ follows the generative procedure below [^2] There is an additional constraint on the domain of $F$ such that any two random vectors $Z_{1},Z_{2}$ from this distribution must have $Z_{1}^{T}Z_{2}\in[0,1]$.:
$Z_{1},...,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(Z_{i}^{T}Z_{j}),\quad i>j$
where the support of $F$ guarantees the validity of the generalized inner product as a probability.
The RDPG model can be further modified following Rubin-Delanchy et al. (2022) resulting in the so-called generalized RDPG (gRDPG). Many other widely used random network models in literature are special cases of the above models. For example, the stochastic block model (SBM) (Holland et al., 1983) and its variants (Karrer and Newman, 2011; Airoldi et al., 2008; Sengupta and Chen, 2018; Li et al., 2022; Jin et al., 2021; Li et al., 2023b), as well as various $\beta$-models (Chatterjee et al., 2011; Chen et al., 2021) can be seen as special cases of the inner product latent space model.
Remark 1.
Depending on the specific model within the latent space model family, several standard model estimation procedures are available to estimate $Z_{i}$’s with a given $A$. These include the gradient descent for the inner product latent space model (Ma et al., 2020) and the adjacency spectral embedding for the RDPG family (Athreya et al., 2018; Rubin-Delanchy et al., 2022).
Remark 2.
In most of the latent space models, the latent vectors are identifiable only up to some operator $\mathcal{O}^{d}$, depending on the specific format of $W$. For example, under the inner product latent space model or the RDPG, $\mathcal{O}^{d}$ can be any $d\times d$ orthogonal transformation. Such operators are usually not crucial when using these models, and would not result in difficulties in our analysis. For notational simplicity, throughout this paper, when we discuss the recovery of latent vectors or their distributions, the recovery can be up to such an unidentifiable operator.
Remark 3.
Another class of random network models, the graphon model (Bickel and Chen, 2009) can also be written in the form of the latent space model in Definition 2. However, there is one key difference: As stated in the definition, the latent space model typically assumes the link function $W$ to be known while the distribution of the latent vectors, $F$, can be unknown. In contrast, a graphon model has one dimensional latent variables $Z_{i}$’s with known $F$ (the uniform distribution), while the link function (graphon) $W$ is unknown. In this paper, we focus on the scenario of the latent space model with known $W$ but unknown $F$. We leave the study of node DP under the graphon model for future work.
Remark 4.
In the literature, some models introduce an additional parameter, which formulates the density of the network. However, the definition of this parameter differs by models; for instance, it appears as a scaling parameter in the RDPG and gRDPG frameworks (Athreya et al., 2021; Rubin-Delanchy et al., 2022) and as the intercept term in the inner product latent space model (Li et al., 2023a). To maintain a uniform format of our framework and focus only on the privacy aspect of networks, we will not introduce such a parameter. Our theory can be readily extended to incorporate such a parameter, if needed.
4 The Proposed Method
4.1 The Prototype Node DP Mechanism: An Oracle Scenario
We introduce the principle of our design regarding how it may preserve network properties. The prototypical idea is to (i) acquire a latent vector for each node, (ii) privatize the latent vectors while preserving their distribution, and (iii) generate a private network through using the private latent vectors. The private network is expected to have the same network properties as the original one, since they are generated from latent vectors of the same distribution.
Under the latent space model in Definition 2, recall that the latent vectors $Z_{i}=(Z_{i1},\ldots,Z_{id})^{T}\in\mathbb{R}^{d},i\in[N]$ are an i.i.d sample from $F$. To introduce our prototype design, we assume that we know two pieces of oracle information.
  • The joint cumulative distribution function (CDF) $F$ of the true latent vectors: From $F$, we can also derive all conditional distributions. Let $F^{\,l\mid 1:(l-1)}(\cdot|Z_{i1},\ldots,Z_{i,l-1})$ be the conditional CDF of $Z_{il}$ given $(Z_{i1},\ldots,Z_{i,l-1})$. When the context is clear, we write $F^{l|1:(l-1)}$ as $F^{l}$ for simplicity. In particular, let the marginal CDF of the first coordinate be $F^{1}$.
  • The true latent vectors $Z_{i},i\in[N]$. Furthermore, let $\mathbf{Z}_{l}=(Z_{1l},\ldots,Z_{Nl})^{T}$ be the $l$th latent vector of all nodes for $l\in[d]$. For each $l$, we can treat $\mathbf{Z}_{l}$ as a sequence of univariate observations.
We next introduce a distribution-invariant privacy mechanism (DIP) proposed in Bi and Shen (2023) to perturb the univariate data to achieve differential privacy, which will then be applied to all coordinates by conditioning. Specifically, applying $F^{1}$ to $\mathbf{Z}_{1}$, we note that $F^{1}(Z_{i1})$ follows Uniform(0,1). We then perturb $F^{1}(Z_{i1})$ by adding an independent noise $e_{i1}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Laplace}(0,1/\varepsilon)$ for $i\in[N]$. Following the standard Laplace mechanism, we know that $F^{1}(Z_{i1})+e_{i1}$ satisfies $\varepsilon$-differential privacy (Dwork, 2006). Let $G$ be the CDF of $F^{1}(Z_{i1})+e_{i1}$ whose expression can be exactly obtained via convolution. Applying $G$ to $F^{1}(Z_{i1})+e_{i1}$ also results in a uniform random variable. Then the privatization mechanism $\mathfrak{m}^{1}$ for $\mathbf{Z}_{1}$ can be expressed as
$\tilde{Z}_{i1}\equiv\mathfrak{m}_{1}(Z_{i1})=(F^{1})^{-1}\circ G\big{(}F^{1}(Z_{i1})+e_{i1}\big{)},\quad i=1,\ldots,N,$
where $\circ$ denotes function composition. We can see that $\tilde{Z}_{i1}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F^{1}$. Meanwhile, $\tilde{Z}_{i1}$ is differentially private (Dwork, 2006).
For $l\geq 2$, we apply the same strategy to privatize each latent vector sequentially, using probability chain rule (Schum, 2001). Let $F^{l}(\cdot):=F^{l}(\cdot|Z_{i1},\ldots,Z_{i,l-1})$ and $\tilde{F}^{l}(\cdot):=F^{l}(\cdot|\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})$. Then
$\tilde{Z}_{il}\equiv\mathfrak{m}_{l}(Z_{il}|\tilde{Z}_{i1},\ldots,\tilde{Z}_{i,l-1})=(\tilde{F}^{l})^{-1}\circ G\big{(}F_{l}(Z_{il})+e_{il}\big{)},$ (3)
where $\mathfrak{m}_{l}(\cdot)$ denotes the privatization process for $\mathbf{Z}_{l}$, $l=2,\ldots,d$. Notice that, since $Z_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F$, we have $F^{l}(Z_{il})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Uniform(0,1)}$. On the other hand, the use of $(\tilde{F}^{l})^{-1}$ in (3) is to guarantee differential privacy of the preceding variables and preserve that $\tilde{Z}_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F$. For notational simplicity, we write
$\tilde{Z}_{i}=\mathfrak{m}(Z_{i};\mathbf{e}_{i},F),$ (4)
where $\tilde{Z}_{i}=(\tilde{Z}_{i1},\ldots,\tilde{Z}_{id})^{T}$, $\mathbf{e}_{i}=(e_{i1},\ldots,e_{id})^{T}$, and $\mathfrak{m}()$ denotes the sequential application of $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{d}$. Here $\tilde{Z}_{i}$ is the differentially private perturbation of $Z_{i}$.
Subsequently, we can follow Definition 2 and use the same generative function $W$ to generate the private adjacency matrix $\tilde{A}$. Then given the fact that $\tilde{Z}_{i}$ follows $F$ and that $W$ is known, $\tilde{A}$ should follow the same distribution as $A$. This ensures that network properties of $\tilde{A}$, especially those that can be presented as summary statistics, remains the same. Meanwhile, due to the fact that $\tilde{Z}_{i}$ satisfies $\varepsilon$-differential privacy, $\tilde{A}$ follows NDP.
The above procedure describes the high-level idea of the proposed method. However, in practice, we do not know either $F$ or $\{{Z}_{i}\}$, thus these have to be estimated from data. The crux of this lies in the privacy requirement. This privacy requirement becomes nontrivial for network data, which requires special designs to be introduced next.
4.2 The Proposed Network Release with Node-wise Estimation
Our goal is to estimate the latent distribution CDF $F$ and the latent vectors $Z_{i}$’s for the prototype procedure. It seems natural to consider the following procedure: we estimate the latent vectors by a standard network estimation method and then estimate the CDF from the estimated latent vectors, and these will be used as the “plug-in” objects in the prototype procedure. Nevertheless, such a straightforward approach will not meet the DP requirement for the following three reasons.
  1. 1.
    Privacy spill-over effect on networks: In contrast to standard multivariate data where instances are assumed to be i.i.d. and many processes can be applied to each instance independently, estimating network models involves the information of pairwise relations. That means, the estimation of one node $i$ relies on the information of its relations with all other nodes $i^{\prime}\neq i$. Thus, when a standard estimation method is used, the resulting estimator of $Z_{i}$ will also contain (private) information of other nodes, for which the DP mechanisms may fail. In other words, the alteration of any node $i$ will lead to the alteration of not only $Z_{i}$, but also other $Z_{i^{\prime}}$’s, such that the released output may not satisfy the probability bound (1) in Definition 1. This is, in our view, the major challenge in handling node-level DP in network data.
  2. 2.
    In our prototype procedure (e.g., see (3)), the outermost layer of operation depends on the inverse CDF. This is the final layer of operation without additional privacy protection. If the estimated CDF is based on the nodes to be released, the privacy of these nodes cannot be protected.
  3. 3.
    In practice, fitting any network models almost always requires additional tuning or model selection. Rigorously speaking, once such procedure involves the to-be-released data, DP cannot be guaranteed unless a specialized tuning procedure is designed.
To rectify these challenges, we assume that we have a network $A$ between $N=m+n$ nodes available. We would like to hold out the network of $m$ nodes while privatizing and releasing only the network between the remaining $n$ nodes. By doing this, we will be able to use the hold-out network as the reference to obtain estimates about $F$ and $Z_{i}$’s and also for tuning if needed while ensuring the node DP for the released network between the $n$ nodes. Figure 1 gives a high-level flowchart demonstrating the proposed method, whose details will be introduced soon. Our data splitting practice is also seen in the graph neural network literature (e.g., Yang et al., 2016; Kipf and Welling, 2017), where a sub-network is used for model training, and another sub-network is used for testing/evaluation. In statistical literature, such a data splitting strategy is also used by Chen and Lei (2018) for community detection problems.
Consider a network of size $N=n+m$. Without loss of generality, we assume the first $n$ nodes are designated for release. The adjacency matrix $A\in\{0,1\}^{(n+m)\times(n+m)}$ can be partitioned into blocks $A^{kr}$ for $k,r\in\{1,2\}$, where $A^{11}$ is an $n\times n$ matrix, $A^{12}$ is an $n\times m$ matrix, $A^{21}$ is an $m\times n$ matrix, and $A^{22}$ is an $m\times m$ matrix, as shown in Figure 1. The proposed scheme will then have the entire adjacency matrix $A$ as the input, and a private version of $A^{11}$, namely $\tilde{A}^{11}$, as the output.
Figure 1: The proposed privacy-preserving scheme illustrated: An $N\times N$ adjacency matrix AA is used as the input, and a private version of a $n\times n$ subnetwork $A^{11}$, namely $\tilde{A}^{11}$, is generated as the output. All privacy vulnerable quantities are colored in red. Each of the latent vectors of nodes $i\leq n$ is separately estimated by a node-wise estimation with the help of hold-out estimates $\{\hat{Z}_{i}\}_{i>n}$. Then $\{\hat{Z}_{i}\}_{i\leq n}$ goes through the DIP mechanism to encode differential privacy. Subsequently, the privatized subnetwork $\tilde{A}^{11}$ is generated from the private latent vectors $\{\tilde{Z}_{i}\}_{i\leq n}$.
Specifically, assuming the network is generated from a general latent space model, we apply one of the standard model estimation procedures discussed in Remark 1 to $A^{22}$. This step would result in an estimate of the latent positions for the hold-out nodes, denoted by $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. From this step, we can estimate the corresponding CDF $F$ for the latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$. This can be achieved through following standard practices such as empirical CDFs, kernel smoothing, splines, or $K$-nearest neighbors (Wasserman, 2006). Denote this estimated $F$ by $\hat{F}$.
Next, we introduce a node-wise estimation procedure to obtain an estimate for the latent vectors of the to-be-released nodes. This is a crucial step to ensure valid node DP for our method. Specifically, for each node $i\in[n]$, we use the information of the $i$th row of $A^{12}$, denoted by $A^{12}_{i*}$, following the latent space model. The estimation can be done by solving the following optimization problem involving the negative loglikelihood:
$\hat{Z}_{i}=\mathop{\arg\min}_{Z_{i}\in{\bf R}^{d}}\sum_{j=n+1}^{n+m}\big{\{}-A_{ij}\log W(Z_{i},\hat{Z}_{j})-(1-A_{ij})\log(1-W(Z_{i},\hat{Z}_{j}))\big{\}},1\leq i\leq n.$ (5)
Problem (5) is simply a pseudo-likelihood estimation procedure, by focusing only on one node at a time, while fixing the estimated latent vectors of the hold-out nodes. Note that the above estimation procedure ensures that, after fixing the estimation on the hold-out set, the estimation of $\hat{Z}_{i}$’s is separable for each $i\in[n]$. That is, the estimation of each $\hat{Z}_{i}$ in (5) uses information of its own row, thus it avoids the privacy spill-over effect.
For most commonly used $W$, this problem is easy to solve using standard algorithms. In many particular cases, the problem can be significantly simplified. For example, under the inner-product latent space model (Example 1), (5) becomes a simple logistic regression problem:
$\hat{Z}_{i}=\mathop{\arg\min}_{(X_{i},\alpha_{i})\in{\bf R}^{d}}\sum_{j=n+1}^{n+m}\big{\{}-A_{ij}\log\sigma(\alpha_{i}+\hat{\alpha}_{j}+X_{i}^{T}\hat{X}_{j})-(1-A_{ij})\log(1-\sigma(\alpha_{i}+\hat{\alpha}_{j}+X_{i}^{T}\hat{X}_{j}))\big{\}}.$ (6)
Other reasonable objectives can be used as alternatives to the negative log-likelihood in (5), and these alternatives may further simplify the algorithm. For example, under the RDPG (Example 2), it is easy to use the sum of squared errors for moment matching as the M-estimation criterion rather than the negative log-likelihood. This is because the estimation becomes a node-wise linear regression procedure.
$\hat{Z}_{i}=\arg\min_{Z_{i}\in{\bf R}^{d}}\sum_{j=n+1}^{n+m}(A_{ij}-Z_{i}^{T}\hat{Z}_{j})^{2}.$ (7)
With the estimated $\hat{F}$ and $\{\hat{Z}_{i}\}_{i\in[n]}$, we can apply the prototypical procedure in Section 4.1 for privatization. In particular, $\hat{F}$ does not utilize any information from the to-be-released nodes, thus can be used to replace $F$. Calling the formula in (4), for each $i\in[n]$, we have
[…]
Subsequently, we generate the private network $\tilde{A}^{11}$ based on $\{\tilde{Z}_{i}\}_{i\in[n]}$ following the same general latent space model. Algorithm 1 provides a summary of the proposed method.
Algorithm 1 The proposed GRAND mechanism by node-wise estimation
0:  A network $A$ of size $N=n+m$ in which the first $n$ nodes are to be privatized, and the latent dimension $d$.
1: Partition AA as in Figure 1.
2: Specify a general latent space model and the corresponding generative function $W$.
3: Estimate the hold-out latent vectors $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$ based on $A^{22}$ using a standard network model estimation procedure as provided in Remark 1.
4: Estimate the CDF $\hat{F}$ based on $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$.
5: Estimate the latent vectors $\{\hat{Z}_{i}\}_{i=1}^{n}$ by the node-wise estimation procedure (5) using $\{\hat{Z}_{i}\}_{i=n+1}^{n+m}$ and $A^{12}$.
6: Apply procedure (8) to privatize latent vectors of the to-be-released nodes, producing $\{\tilde{Z}_{i}\}_{i=1}^{n}$.
7: Generate the private adjacency matrix $\tilde{A}^{11}$ from $\{\tilde{Z}_{i}\}_{i=1}^{n}$ using $W$.
7:  A privatized $\tilde{A}^{11}$ of size $n$.
Similar to the requirements on the holdout dataset in DIP (Bi and Shen, 2023), the holdout set should not be accessed, altered, queried, or released, and should be deleted immediately once Algorithm 1 is implemented. Alternatively, additional approaches beyond the discussion provided in Bi and Shen (2023) may also be viable. First, we note that $\hat{F}$ is the estimated CDF of the latent vectors. With additional regularity conditions, one can apply some particular mechanisms, including the Private Signed Measure Mechanism (He et al., 2023), which generate $\hat{F}$ that are both differentially private and asymptotically consistent.
Second, it is also possible in the social network context that different nodes have different preferences in privacy protection. In such a case, one can construct the holdout set using nodes that are less sensitive to data release. On the one hand, one can apply differential privacy for functions (Hall et al., 2013) and achieves $(\varepsilon,\delta)$-differential privacy for a kernel density built on the holdout set. On the other hand, to accommodate the possible distributional discrepancy between the holdout and the release sets, a transfer learning model as discussed in Pan and Yang (2010) and Zhuang et al. (2020) can be used to replace (5). Nevertheless, the adoption of both a $(\varepsilon,\delta)$-differentially private holdout set and a transfer learning model entails additional theoretical establishment, which is acknowledged in future work in Section 8.
Remark 5.
In the estimation procedure, it can be seen that $A^{11}$ is not used. This is to avoid the privacy spill-over issue. Therefore, from a statistical estimation perspective, the estimation is essentially based on a size-$(m+1)$ network rather than a size-$N$ network. which will be explicitly verified in our theory and numerical results. The statistical efficiency reduction from size $N$ to size $m$ is the price we pay for the node-level privacy.
Remark 6.
The computational bottleneck of the algorithm is on the standard model estimation procedure in Step 3, which depends on $m$. The node-wise estimation is typically trivial. Moreover, it can be done completely in parallel, which further improves the computational feasibility.
Remark 7 (Selection of the private network size).
Intuitively, we hope to maximize the released size $n$, subject to m=Nnm=N-n being large enough for model estimation. In practice, this choice may also depend on other factors, such as the target size of the released networks, computational resources and differences in privacy levels for subjects in the network. In our numerical studies, we use $n=0.5N$ for simplicity.
Remark 8.
While we focus on undirected networks in this paper, it is not difficult to see that the procedure is ready to be extended to directed networks. In the directed networks, the corresponding latent space model would associate with each node a “sending” latent vector and a “receiving” latent vector. We will have to estimate both latent vectors for each node, using both $A^{12}$ and $A^{21}$, and then combine the “sending” and “receiving” vectors together in the privatization step.
5 Theoretical Guarantees
In this section, we study the theoretical guarantees for the released network from our method. The very basic result is that our method gives differential privacy at the node level.
Theorem 1.
Suppose Algorithm 1 is used to process the input network under the privacy budget $\varepsilon$. The release network $\tilde{A}^{11}$ is node $\varepsilon$-differentially private for each node $i\in[n]$.
Note that, even though we use the latent space model to motivate our method, the node DP guarantee does not rely on the latent space model. That means, even if the network data is not generated by latent space models, applying our method still protects privacy.
However, the latent space model assumption is needed to guarantee the preservation of network properties. As implemented in Algorithm 1, the estimation procedure for all $\tilde{Z}_{i}$’s involves the common $A^{22}$. That is, $\tilde{Z}_{i}$’s may have marginal dependence and can no longer be assumed to be i.i.d. This is in contrast to the true $Z_{i}$’s. Therefore, in the rest of this section, none of the theoretical results rely on the $\tilde{Z}_{i}$’s being i.i.d., which is one of the major theoretical contribution of this work.
As the focus of this section, we will show that the privatized $\tilde{Z}_{i}$’s are guaranteed to asymptotically maintain the original distribution of $Z_{i}$’s. Such consistency indicates that the distributional properties of the original network are asymptotically preserved. In particular, as an indication of latent space consistency, we will show that the released private network $\tilde{A}$ can preserve motif counts of the original network for any given motif.
5.1 Preservation of Network Properties: The Latent Distribution Consistency
As discussed, to study the preservation of network properties, we will assume that the true network data is generated from a latent space model.
Assumption A1 (Latent space model).
Suppose the network $A$ is generated from the general latent space model (Definition 2) with a fixed dimension $d$.
We then introduce the following regularity assumptions for the model.
Assumption A2 (Latent distribution).
The true latent space distribution satisfies the following properties.
  • The joint CDF $F$ is a continuous distribution on a compact support $S$. The joint density $f(x_{1},\dots,x_{d})$ of $F$ is continuous and bounded by a constant $C_{\mathrm{up}}>0$.
  • If $d>1$, for each $l=2,\dots,d$:
    • The conditional CDF $F_{l|1:(l-1)}(x|u)$ is Lipschitz continuous in $(x,u)$ and strictly increasing in $x$.
    • The marginal density $f_{1:{(l-1)}}(u)$ is continuous, bounded and lower bounded by a positive constant.
Assumption A3 (Kernel smoothing for CDF estimation).
For each $1\leq j\leq d$, the estimation of the conditional CDF $F^{l|1:(l-1)}$ is constructed from $\{\hat{Z}_{i}\}_{i>n}$ by a kernel estimation using a bounded, Lipschitz continuous kernel $K:{\bf R}^{d-1}\to[0,\infty)$ (with a constant $L_{K}$) such that $\int_{{\bf R}^{d-1}}K(u)\,du=1$ and a properly chosen bandwidth $h=(\log m)^{-c}$ for a constant c>0c>0333This choice does not necessarily give the optimal rate in estimation. The optimal choice is hinged on the rate of A4 in a more complicated way, and the detailed study is not be the focus of our current work. We choose this rate for simple interpretations., defined as
$\widehat{F}^{\,l\mid 1:(l-1)}(x\,\big{|}\,x_{1},\dots,x_{l-1})=\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Z}_{n+i,l}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{x_{1}-\hat{Z}_{i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{x_{1}-\hat{Z}_{n+i,1}}{h},\dots,\tfrac{x_{l-1}-\hat{Z}_{n+i,l-1}}{h}\Bigr{)}}.$ (9)
Assumption A4 (Latent embedding error).
Suppose $\hat{Z}_{i},i\in[N]$ are the estimated latent vectors in Algorithm 1. There exists a constant $c>0$, such that (subject to the unidentifiable operator $\mathcal{O}^{d}$)
$\max_{1\leq i\leq N}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}=o((\log m)^{-c})$
with probability approaching 1 as $m\to\infty$.
Recall that we use $\{\tilde{Z}_{i}\}_{i\in[n]}$ to generate the released network $\tilde{A}^{11}$. Given the latent vectors, the generating mechanism is always independent Bernoulli following Definition 2. Therefore, to preserve network properties, the primary guarantee we need is that the resulting distribution of the $\tilde{Z}_{i}$’s should be roughly the same as the true distribution in a proper sense. Theorem 2 expresses this property.
Theorem 2 (Individual latent distribution consistency).
Suppose assumptions A1A4 hold. Let $\tilde{F}^{(i)}$ be the joint CDF of the random vector $\tilde{Z}_{i}$. For any $1\leq i\leq n$,
$\sup_{x\in{\bf R}^{d}}|\tilde{F}^{(i)}(x)-F(x)|\to 0$
in probability as $m\to\infty$.
On the one hand, we can say that the privatized $\tilde{Z}_{i}$ has the asymptotically correct distribution individually. On the other hand, due to the fact that $\{\tilde{Z}_{i}\}_{i\in[n]}$ cannot be treated as an i.i.d sample from $F$ collectively, the latent distribution consistency is for each individual node $i$, rather than a uniform convergence that holds simultaneously for all $i\in[n]$. Fortunately, although the uniform convergence is not achievable, the privatized latent vectors still give a similar collective behavior for the resulting network. Specifically, we now show that the empirical CDF of the privatized latent vectors is asymptotically consistent, that is, having behaviors similar to that of i.i.d true latent vectors.
Theorem 3 (Latent CDF consistency).
Suppose assumptions A1A4 hold. Let $\tilde{F}_{n}$ be the empirical CDF of the privatized latent vectors based on $\tilde{Z}_{i}\in{\bf R}^{d}$, $i=1,\cdots,n$ where
$\tilde{F}_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\dots,\tilde{Z}_{id}\leq x_{d}\}}.$
Then $\tilde{F}_{n}$ converges uniformly to $F$ in probability. That is, as $n,m\to\infty$,
$\sup_{x\in{\bf R}^{d}}|\tilde{F}_{n}(x)-F(x)|\xrightarrow{P}0.$
Assumption A4 for Theorems 2 and 3 requires that the latent vectors can be accurately recovered for all nodes $i\in[N]$. For nodes in the hold-out set, this is not difficult by using the standard network estimation methods. However, for nodes $i\in[n]$, such a requirement needs additional study of the node-wise estimation procedure. Next, we show that under both the inner product latent space model and the RDPG, the proposed estimation algorithm satisfies the requirement of A4.
Theorem 4 (Error bound for node-wise estimation).
Suppose $n=o(e^{m})$. For a constant $c>0$, consider the following models.
  1. 1.
    The network is generated from the inner product latent space model and the latent vector set $\{\hat{Z}_{i}\}_{i\in[n]}$ is estimated by (6).
  2. 2.
    The network is generated from the RDPG model and the latent vector set $\{\hat{Z}_{i}\}_{i\in[n]}$ is estimated by (7).
If the true latent distribution $F$ has a bounded domain and a positive definite second-order moment: $\Sigma=\mathbb{E}[ZZ^{\top}]\succeq\mu I_{d}$ for some constant $\mu>0$, while the hold-out estimates (up to an unidentifiable transformation) satisfy
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=\delta_{m}=o(1)$
with probability approaching 1, the estimates $\{\hat{Z}_{i}\}_{i\in[N]}$ from Algorithm 1 satisfy
$\max_{i\in[n]}\|{\hat{Z}_{i}-Z_{i}}\|\leq C\max(\delta_{m},\frac{\sqrt{\log{m}}+\sqrt{\log n}}{\sqrt{m}})$
for a constant $C>0$ with probability approaching 1.
As specific examples, under both the inner product latent space model and RDPG, the assumption A4 in the latent distribution consistency results can be satisfied using commonly used estimation approaches.
Corollary 1.
Suppose Assumptions A1A3 hold. In either of the following two cases:
  1. 1.
    Under the inner product latent space model, suppose that the estimator of Li et al. (2023a) is used to estimate the hold-out latent vectors in Step 3 of Algorithm 1, and that the model satisfies the regularity conditions in Li et al. (2023a).
  2. 2.
    Under the RDPG model, the method of Rubin-Delanchy et al. (2022) is used to estimate the hold-out latent vectors in Step 3 of Algorithm 1.
The conclusions of Theorems 2 and 3 hold.
Subsequently, the established consistency of the latent distribution for the privatized latent vectors indicate correct recovery of many network properties.
5.2 Consistency of Privatized Network Moments
In this subsection, we show that, as a special application of the latent distribution consistency, the privatized network $\tilde{A}$ maintains the moments/motif counts of the original network.
Network moments are commonly used to measure various aspects of network structures for inference and comparison tasks (Cook, 1971; Milo et al., 2002; Bickel et al., 2011; Bhattacharyya and Bickel, 2015; Maugis et al., 2020; Levin and Levina, 2019; Zhang and Xia, 2022; Qi et al., 2024), as they allow comparison across networks of different sizes and node sets. More broadly, many other useful network statistics, though not directly expressed as network moments, can be written as functions of network moments of multiple motifs, such as the clustering coefficient. Therefore, the guarantee of network moment recovery can indicate the valid of many comparable analyses based on privatized data.
Recall that $\mathcal{V}(A)$ and $\mathcal{E}(A)$ are the node set and edge set of network $A$, respectively. Following Qi et al. (2024), a graph $R$ is a subgraph of $A$, written as $R\subset A$, if $\mathcal{V}(R)\subset\mathcal{V}(A)$ and $\mathcal{E}(R)\subset\mathcal{E}(A)$. Two graphs $R$ and $A$ are isomorphic, denoted by $R\cong A$, when there exists a bijective function $\phi$: $\mathcal{V}(R)\to\mathcal{V}(A)$ such that $(v_{i},v_{j})\in\mathcal{E}(R)$ if and only if edge $\phi(v_{i}),\phi(v_{j})]\in\mathcal{E}(A)$.
A motif refers to a (usually simple) graph, such as an edge (), a 2-star/V-shape (), a triangle (), or a 3-star (), which forms the building blocks of larger graphs. Here we denote a motif by $R$, with $|V(R)|=r$ representing the number of nodes. We focus exclusively on connected motifs. For a network $A$ and a motif $R$, the motif density of $R$ in $A$ is defined as the normalized number of subgraphs of $A$ that are isomorphic to $R$:
$X_{R}(A)=\big{|}\{S:S\subset A,S\cong R\}\big{|}/\binom{n}{r}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}{\mathbb{I}}(A_{[i_{1},\cdots,i_{r}]}\cong R)/\binom{n}{r}$, (10)
in which $A_{[i_{1},\cdots,i_{r}]}$ is the subgraph of $A$ induced by nodes $\{i_{1},\cdots,i_{r}\}$. We call $X_{R}(A)$ the network moment of $A$ with respect to motif $R$. For example, when $R$ is an edge, then $X_{R}(A)$ is the edge density of $A$. Network moments are summary statistics of the whole network structures. Intuitively, under the latent space model, they are determined by the collective behavior of the latent vectors. And therefore, preservation of latent vector distributions should also indicate the preservation of network moments. This is formally stated in the following result.
Theorem 5 (Consistency of network moments).
Let $R$ be a fixed motif on $r$ nodes that does not depend on $n$. Suppose $X_{R}(A^{11})$ is the motif density of $R$ in network $A^{11}$, as defined in (10). Similarly, let $X_{R}(\tilde{A}^{11})$ be motif density of $R$ in the privatized network $\tilde{A}^{11}$ from Algorithm 1. When Assumptions A1A4 hold, we have, as $n,m\to\infty$,
$X_{R}(A^{11})-X_{R}(\tilde{A}^{11})\to 0$
in probability.
It is also trivial to see that the claim can be extended to any function of network motif counts such as the clustering coefficient. In the meantime, Theorem 5 provides crucial convenience in applications. For example, as demonstrated in Qi et al. (2024), one could compare multiple unmatchable networks by comparing the distribution of their moments. Theorem 5 indicates that we can conduct the same comparison inference using networks privatized by GRAND.
6 Simulation Experiments
In this section, we introduce simulation experiments to evaluate the performance of the proposed method. In the experiments, we fix the node DP budget, and assess the extent to which the privatized network can preserve the structural properties of the original network. We consider the two commonly used network models, inner-product latent space model and RDPG, as described in Section 3.3. For both methods, we generate adjacency matrices according to their respective models, and use them as the respective original datasets.
Experiment configuration
Under both models, we set certain key parameters as follows. We vary the released size and hold-out size to be $n=m\in\{2000,4000\}$, and network density $\rho\in\{0.025,0.05,0.1\}$. For example, for a network of size 2000, the expected average degree is 50 and 100, respectively. Also, we set the dimension $d\in\{3,6\}$ and the privacy budget $\varepsilon\in\{1,2,5,10\}$. Typically, $\varepsilon=1$ corresponds to a strong privacy protection in practice, while $\varepsilon=10$ gives much weaker protection. The latent distribution $F$ is a truncated Gaussian mixture model, as a generalization of the mechanism of Li et al. (2023a) for more heterogeneity.
Methods in comparison
In addition to our proposed method, we include two additional benchmark methods for comparison.
  1. 1.
    The first one is naively applying the Laplace mechanism of Dwork (2006) to $\{\hat{Z}_{i}\}_{i\in[n]}$ and then generating the network from the resulting latent vectors (“Laplace”). Following a similar argument as in Theorem 1, it also satisfies node DP, but offers no guarantee for structural preservation. The comparison with this benchmark illustrates how important it is to consider the latent distribution consistency in the design of the proposed method in practice.
  2. 2.
    We also consider a non-private method: We directly estimate the network models from the size $m$ network using a standard network estimation, rather than the node-wise estimation method. Specifically, we use the gradient descent model fitting of Ma et al. (2020); Li et al. (2023a) for the inner product model and the adjacency spectral embedding (Sussman et al., 2014) for the RDPG. After that, we generate a new size-$n$ network from the estimated model without introducing any privatization step as the result (in this case, we have $m=n$). We call this the “Hat” network. The “Hat” method corresponds to the scenario where we do not impose DP at all and only use a standard approach to estimate the network model and generate new data. Comparing the “Hat” network with the original network gives us a non-private benchmark. As discussed in Section 4, our model estimation before the privatization is essentially based on $m+1\approx m$ nodes. Therefore, the “Hat” estimate can be seen as a comparable one with our method when $\varepsilon\to\infty$. The comparison between our method and “Hat” can show the price of structural preservation we pay for the privacy.
Performance evaluation metrics
We measure the structural preservation by examining the distributions of five node-level statistics across the $n$ nodes in each released network, and compare these distributions with those of the original network. Specifically, for each local statistic, denote its value at node $i$ by $T_{i}(A^{11})$ on network $A^{11}$, and its value at node $i$ on the released network $\tilde{A}^{11}$ as $T_{i}(\tilde{A}^{11})$. We want to compare the distribution of $\{T_{i}(\tilde{A}^{11})\}_{1\leq i\leq n}$ with that of $\{T_{i}(A^{11})\}_{1\leq i\leq n}$. We focus on the following five local statistics as representative of network properties.
  1. 1.
    Node degree: $T_{i}(A)=\sum_{j}A_{ij}$.
  2. 2.
    V-shape count: $T_{i}(A)=\binom{\sum_{j}A_{ij}}{2}$. It measures how many V-shape motifs involve node $i$ as the center. To be consistent with the motif count definition, this also includes triangles.
  3. 3.
    Triangle count: $T_{i}(A)=\frac{1}{2}\sum_{j,k}A_{ij}A_{jk}A_{ki}$. It measures how many triangles involve node $i$.
  4. 4.
    Eigen-centrality: $T_{i}(A)=v(A)_{i}$ where $v(A)$ is the eigenvector corresponding to the largest eigenvalue of $A$. It is a popular measure of how central node $i$ is in the network based on the spectral structure.
  5. 5.
    Harmonic centrality: $T_{i}(A)=\frac{1}{\sum_{j\neq i}d_{A}(i,j)}$ where $d_{A}(i,j)$ is the geodesic distance between node $i$ and node $j$ in the network.
Therefore, for each pair of $A^{11}$ and $\tilde{A}^{11}$, we can visualize the resulting distributions of the five statistics and compare them. However, to aggregate the results in a meaningful way, we also have to introduce the metrics to measure the difference between a pair of distributions. In our study, we use the Wasserstein distance metric. For the highly skewed distributions (such as V-shape, triangle counts), directly calculating the distances becomes numerically unstable, so we applied a logarithm transformation to the data and evaluate the recovery of the log-transformed distributions. Under each configuration, we independently repeat the experiments 100 times and take the average distances as the final results.
6.1 Results under the inner product latent space model
For the inner-product latent space model, we generalize the simulation setup of Li and Le (2023) to incorporate more heterogeneity. Following the model definition in Example 1, we sample latent vectors $X_{i}$’s from a mixture of truncated Gaussian distributions and generate $\alpha_{i}$’s from a uniform distribution. We then rescale all $\alpha_{i}$’s together so the resulting network from the model would have the desired density according to our previous configuration. We randomly sample half of the nodes to be the hold-out data, and privatize the network between the other half of the nodes.
Tables 1 and 2 give the distribution preservation errors for $n=2000$ and $n=4000$, respectively. In all experimental configurations, the distance is measured by the Wasserstein distance, and the network is generated from the inner product model. The results are summarized as follows.
Table 1: Node-level statistic distribution preservation results with $n=2000$ under the inner product latent space model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.047 (<0.001<0.001) 0.068 (0.003) 3.033 (0.002) 0.017 (<0.001<0.001) 0.05 (0.002) 2.334 (0.001) 0.008 (<0.001<0.001) 0.031 (0.002) 1.632 (0.001)
2 0.047 (<0.001<0.001) 0.063 (0.002) 2.95 (0.002) 0.017 (<0.001<0.001) 0.047 (0.002) 2.264 (0.002) 0.008 (<0.001<0.001) 0.031 (0.002) 1.582 (0.001)
5 0.047 (<0.001<0.001) 0.046 (0.001) 2.554 (0.004) 0.017 (<0.001<0.001) 0.036 (0.002) 1.924 (0.004) 0.008 (<0.001<0.001) 0.026 (0.001) 1.333 (0.003)
10 0.047 (<0.001<0.001) 0.047 (0.001) 1.804 (0.006) 0.017 (<0.001<0.001) 0.03 (0.001) 1.3 (0.005) 0.008 (<0.001<0.001) 0.024 (0.001) 0.881 (0.004)
6 1 0.051 (<0.001<0.001) 0.074 (0.002) 3.043 (0.001) 0.019 (<0.001<0.001) 0.051 (0.002) 2.347 (0.001) 0.008 (<0.001<0.001) 0.029 (0.002) 1.642 (0.001)
2 0.051 (<0.001<0.001) 0.068 (0.002) 3.022 (0.001) 0.019 (<0.001<0.001) 0.046 (0.002) 2.332 (0.001) 0.008 (<0.001<0.001) 0.03 (0.001) 1.63 (0.001)
5 0.051 (<0.001<0.001) 0.053 (0.001) 2.895 (0.002) 0.019 (<0.001<0.001) 0.038 (0.002) 2.227 (0.002) 0.008 (<0.001<0.001) 0.029 (0.001) 1.555 (0.001)
10 0.051 (<0.001<0.001) 0.055 (0.002) 2.534 (0.004) 0.019 (<0.001<0.001) 0.033 (0.001) 1.928 (0.003) 0.008 (<0.001<0.001) 0.021 (0.001) 1.336 (0.002)
V-shape 3 1 0.099 (0.001) 0.142 (0.005) 6.134 (0.003) 0.034 (0.001) 0.102 (0.005) 4.7 (0.003) 0.016 (<0.001<0.001) 0.063 (0.003) 3.277 (0.002)
2 0.099 (0.001) 0.132 (0.005) 5.968 (0.004) 0.034 (0.001) 0.097 (0.005) 4.559 (0.004) 0.016 (<0.001<0.001) 0.063 (0.003) 3.178 (0.003)
5 0.099 (0.001) 0.097 (0.003) 5.171 (0.008) 0.034 (0.001) 0.074 (0.003) 3.877 (0.008) 0.016 (<0.001<0.001) 0.053 (0.003) 2.677 (0.006)
10 0.099 (0.001) 0.097 (0.003) 3.662 (0.012) 0.034 (0.001) 0.061 (0.003) 2.624 (0.01) 0.016 (<0.001<0.001) 0.049 (0.002) 1.772 (0.008)
6 1 0.108 (0.001) 0.155 (0.004) 6.151 (0.003) 0.04 (0.001) 0.103 (0.004) 4.724 (0.002) 0.017 (<0.001<0.001) 0.058 (0.003) 3.298 (0.002)
2 0.108 (0.001) 0.142 (0.004) 6.11 (0.003) 0.04 (0.001) 0.094 (0.003) 4.693 (0.002) 0.017 (<0.001<0.001) 0.06 (0.003) 3.274 (0.002)
5 0.108 (0.001) 0.112 (0.002) 5.855 (0.004) 0.04 (0.001) 0.078 (0.003) 4.484 (0.003) 0.017 (<0.001<0.001) 0.058 (0.003) 3.123 (0.003)
10 0.108 (0.001) 0.115 (0.003) 5.131 (0.008) 0.04 (0.001) 0.066 (0.002) 3.885 (0.006) 0.017 (<0.001<0.001) 0.043 (0.002) 2.683 (0.005)
Triangle 3 1 0.055 (0.001) 0.127 (0.006) 8.774 (0.005) 0.029 (0.001) 0.132 (0.007) 6.768 (0.004) 0.015 (<0.001<0.001) 0.094 (0.005) 4.799 (0.003)
2 0.055 (0.001) 0.12 (0.005) 8.608 (0.006) 0.029 (0.001) 0.126 (0.008) 6.622 (0.005) 0.015 (<0.001<0.001) 0.092 (0.005) 4.694 (0.004)
5 0.055 (0.001) 0.122 (0.006) 7.789 (0.009) 0.029 (0.001) 0.1 (0.006) 5.899 (0.009) 0.015 (<0.001<0.001) 0.08 (0.005) 4.142 (0.007)
10 0.055 (0.001) 0.142 (0.006) 6.137 (0.015) 0.029 (0.001) 0.09 (0.005) 4.449 (0.013) 0.015 (<0.001<0.001) 0.077 (0.004) 3.025 (0.01)
6 1 0.066 (0.001) 0.129 (0.005) 8.794 (0.005) 0.034 (0.001) 0.102 (0.006) 6.781 (0.003) 0.017 (<0.001<0.001) 0.073 (0.005) 4.78 (0.003)
2 0.066 (0.001) 0.139 (0.006) 8.754 (0.005) 0.034 (0.001) 0.092 (0.005) 6.75 (0.003) 0.017 (<0.001<0.001) 0.077 (0.005) 4.754 (0.003)
5 0.066 (0.001) 0.163 (0.007) 8.501 (0.005) 0.034 (0.001) 0.085 (0.005) 6.531 (0.005) 0.017 (<0.001<0.001) 0.075 (0.004) 4.592 (0.004)
10 0.066 (0.001) 0.224 (0.009) 7.754 (0.009) 0.034 (0.001) 0.084 (0.004) 5.89 (0.007) 0.017 (<0.001<0.001) 0.057 (0.003) 4.101 (0.006)
Eigen Centrality 3 1 0.012 (0.001) 0.042 (0.003) 0.333 (0.005) 0.01 (0.001) 0.041 (0.003) 0.25 (0.005) 0.011 (0.001) 0.031 (0.002) 0.164 (0.004)
2 0.012 (0.001) 0.039 (0.003) 0.288 (0.005) 0.01 (0.001) 0.04 (0.003) 0.199 (0.005) 0.011 (0.001) 0.028 (0.002) 0.106 (0.003)
5 0.012 (0.001) 0.042 (0.003) 0.198 (0.005) 0.01 (0.001) 0.041 (0.003) 0.108 (0.004) 0.011 (0.001) 0.031 (0.002) 0.043 (0.001)
10 0.012 (0.001) 0.039 (0.004) 0.069 (0.004) 0.01 (0.001) 0.039 (0.003) 0.048 (0.002) 0.011 (0.001) 0.028 (0.002) 0.097 (0.003)
6 1 0.016 (0.001) 0.045 (0.003) 0.46 (0.007) 0.012 (0.001) 0.04 (0.003) 0.408 (0.007) 0.012 (0.001) 0.032 (0.002) 0.332 (0.006)
2 0.016 (0.001) 0.048 (0.003) 0.361 (0.006) 0.012 (0.001) 0.04 (0.003) 0.293 (0.006) 0.012 (0.001) 0.033 (0.002) 0.226 (0.005)
5 0.016 (0.001) 0.048 (0.003) 0.242 (0.005) 0.012 (0.001) 0.042 (0.003) 0.168 (0.005) 0.012 (0.001) 0.031 (0.002) 0.101 (0.004)
10 0.016 (0.001) 0.053 (0.004) 0.145 (0.005) 0.012 (0.001) 0.036 (0.003) 0.066 (0.003) 0.012 (0.001) 0.032 (0.002) 0.045 (0.002)
Harmonic Centrality 3 1 13.597 (0.107) 14.552 (0.604) 579.684 (0.429) 3.616 (0.047) 6.631 (0.292) 453.333 (0.282) 0.691 (0.011) 2.933 (0.151) 392.163 (0.241)
2 13.597 (0.107) 13.014 (0.594) 551.887 (0.617) 3.616 (0.047) 6.335 (0.304) 427.719 (0.508) 0.691 (0.011) 2.927 (0.153) 372.992 (0.463)
5 13.597 (0.107) 7.125 (0.394) 433.659 (0.939) 3.616 (0.047) 4.681 (0.21) 318.755 (0.98) 0.691 (0.011) 2.466 (0.144) 284.812 (0.781)
10 13.597 (0.107) 6.481 (0.264) 276.721 (1.028) 3.616 (0.047) 3.376 (0.17) 171.327 (0.916) 0.691 (0.011) 2.299 (0.116) 157.138 (0.854)
6 1 15.681 (0.1) 15.535 (0.573) 581.975 (0.328) 3.736 (0.038) 6.732 (0.253) 458.477 (0.148) 0.716 (0.01) 2.575 (0.148) 398.107 (0.118)
2 15.681 (0.1) 13.531 (0.593) 574.048 (0.364) 3.736 (0.038) 6.134 (0.212) 452.14 (0.198) 0.716 (0.01) 2.683 (0.14) 392.887 (0.169)
5 15.681 (0.1) 8.566 (0.413) 528.079 (0.618) 3.736 (0.038) 4.812 (0.207) 411.398 (0.502) 0.716 (0.01) 2.589 (0.118) 361.959 (0.389)
10 15.681 (0.1) 7.184 (0.321) 419.364 (0.942) 3.736 (0.038) 3.472 (0.123) 313.355 (0.816) 0.716 (0.01) 1.94 (0.084) 282.385 (0.741)
  • When focusing on the proposed GRAND method, it can be seen that as $\varepsilon$ increases, the performance becomes better for all statistics, which is expected due to the potential privacy-utility tradeoff.
  • Comparing our method with the Hat method, we can see that it indeed yields worse preservation. This is also expected since this reflects the cost of achieving privacy. Noticeably, for weak privacy protection, such as $\varepsilon=10$, the performance of our method becomes comparable to the non-private Hat method. Recall that the Hat method is based on an estimation of a size-$m$ network. This verifies our previous statement that the node-wise estimation basically gives the accuracy as a standard estimation of size-$m$ network. It is also noted that for several metrics, our method gives even a slightly better preservation than the Hat method for $\varepsilon=10$. This is indeed because in our method, we also use the information that the latent vectors are i.i.d following $F$, in processing the network. However, for the Hat method using the estimation of Ma et al. (2020), this distributional information is not used. In other words, when $\varepsilon$ becomes large, the advantage of using the distributional information outweighs the small loss due to the privacy protection.
  • Comparing our method with the Laplace mechanism, where both methods satisfy the same strictness of node DP, it is evident that our method achieves much better preservation of network properties. The Laplace mechanism blindly introduces the noises but fails to retain the network structures.
  • The aforementioned patterns are consistent across different configurations of latent dimensions, network sizes and densities. This demonstrates the effectiveness and robustness of our method across a wide range of scenarios.
Table 2: Node-level statistic distribution preservation results with $n=4000$ under the inner product latent space model, measured by the average Wasserstein distance over 100 replications. The numbers in parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.033 (<0.001<0.001) 0.049 (0.002) 3.04 (0.001) 0.01 (<0.001<0.001) 0.03 (0.001) 2.34 (0.001) 0.004 (<0.001<0.001) 0.02 (0.001) 1.635 (0.001)
2 0.033 (<0.001<0.001) 0.042 (0.002) 2.956 (0.002) 0.01 (<0.001<0.001) 0.03 (0.001) 2.276 (0.001) 0.004 (<0.001<0.001) 0.019 (0.001) 1.588 (0.001)
5 0.033 (<0.001<0.001) 0.031 (0.001) 2.552 (0.003) 0.01 (<0.001<0.001) 0.023 (0.001) 1.949 (0.003) 0.004 (<0.001<0.001) 0.017 (0.001) 1.341 (0.003)
10 0.033 (<0.001<0.001) 0.035 (0.001) 1.787 (0.005) 0.01 (<0.001<0.001) 0.018 (0.001) 1.327 (0.005) 0.004 (<0.001<0.001) 0.016 (0.001) 0.895 (0.004)
6 1 0.04 (<0.001<0.001) 0.053 (0.002) 3.05 (0.001) 0.012 (<0.001<0.001) 0.034 (0.001) 2.348 (0.001) 0.004 (<0.001<0.001) 0.018 (0.001) 1.644 (0.001)
2 0.04 (<0.001<0.001) 0.048 (0.001) 3.031 (0.001) 0.012 (<0.001<0.001) 0.03 (0.001) 2.334 (0.001) 0.004 (<0.001<0.001) 0.017 (0.001) 1.634 (0.001)
5 0.04 (<0.001<0.001) 0.036 (0.001) 2.91 (0.001) 0.012 (<0.001<0.001) 0.026 (0.001) 2.238 (0.001) 0.004 (<0.001<0.001) 0.015 (0.001) 1.565 (0.001)
10 0.04 (<0.001<0.001) 0.041 (0.001) 2.555 (0.003) 0.012 (<0.001<0.001) 0.022 (0.001) 1.957 (0.003) 0.004 (<0.001<0.001) 0.014 (0.001) 1.358 (0.002)
V-shape 3 1 0.068 (0.001) 0.1 (0.004) 6.113 (0.002) 0.02 (<0.001<0.001) 0.061 (0.003) 4.696 (0.002) 0.008 (<0.001<0.001) 0.041 (0.002) 3.276 (0.002)
2 0.068 (0.001) 0.085 (0.003) 5.945 (0.003) 0.02 (<0.001<0.001) 0.06 (0.003) 4.568 (0.002) 0.008 (<0.001<0.001) 0.039 (0.002) 3.182 (0.002)
5 0.068 (0.001) 0.063 (0.002) 5.137 (0.007) 0.02 (<0.001<0.001) 0.047 (0.002) 3.913 (0.007) 0.008 (<0.001<0.001) 0.035 (0.002) 2.689 (0.005)
10 0.068 (0.001) 0.071 (0.002) 3.603 (0.011) 0.02 (<0.001<0.001) 0.036 (0.001) 2.666 (0.009) 0.008 (<0.001<0.001) 0.032 (0.002) 1.796 (0.007)
6 1 0.082 (<0.001<0.001) 0.108 (0.003) 6.132 (0.002) 0.024 (<0.001<0.001) 0.069 (0.003) 4.712 (0.002) 0.009 (<0.001<0.001) 0.037 (0.002) 3.295 (0.001)
2 0.082 (<0.001<0.001) 0.098 (0.003) 6.095 (0.002) 0.024 (<0.001<0.001) 0.061 (0.003) 4.682 (0.002) 0.009 (<0.001<0.001) 0.035 (0.002) 3.274 (0.001)
5 0.082 (<0.001<0.001) 0.073 (0.002) 5.852 (0.003) 0.024 (<0.001<0.001) 0.054 (0.002) 4.491 (0.003) 0.009 (<0.001<0.001) 0.03 (0.002) 3.136 (0.002)
10 0.082 (<0.001<0.001) 0.084 (0.002) 5.142 (0.006) 0.024 (<0.001<0.001) 0.045 (0.002) 3.929 (0.006) 0.009 (<0.001<0.001) 0.028 (0.001) 2.722 (0.004)
Triangle 3 1 0.024 (<0.001<0.001) 0.097 (0.005) 8.759 (0.004) 0.013 (<0.001<0.001) 0.081 (0.005) 6.768 (0.003) 0.008 (<0.001<0.001) 0.062 (0.004) 4.796 (0.003)
2 0.024 (<0.001<0.001) 0.088 (0.004) 8.587 (0.005) 0.013 (<0.001<0.001) 0.079 (0.004) 6.633 (0.003) 0.008 (<0.001<0.001) 0.06 (0.004) 4.695 (0.003)
5 0.024 (<0.001<0.001) 0.094 (0.004) 7.747 (0.007) 0.013 (<0.001<0.001) 0.066 (0.004) 5.933 (0.007) 0.008 (<0.001<0.001) 0.055 (0.003) 4.148 (0.006)
10 0.024 (<0.001<0.001) 0.116 (0.005) 6.052 (0.012) 0.013 (<0.001<0.001) 0.054 (0.003) 4.491 (0.011) 0.008 (<0.001<0.001) 0.052 (0.003) 3.05 (0.01)
6 1 0.024 (<0.001<0.001) 0.091 (0.004) 8.783 (0.004) 0.015 (<0.001<0.001) 0.073 (0.004) 6.765 (0.003) 0.01 (<0.001<0.001) 0.049 (0.003) 4.777 (0.002)
2 0.024 (<0.001<0.001) 0.1 (0.005) 8.746 (0.004) 0.015 (<0.001<0.001) 0.068 (0.004) 6.734 (0.003) 0.01 (<0.001<0.001) 0.043 (0.003) 4.754 (0.002)
5 0.024 (<0.001<0.001) 0.12 (0.005) 8.496 (0.004) 0.015 (<0.001<0.001) 0.064 (0.003) 6.531 (0.004) 0.01 (<0.001<0.001) 0.037 (0.002) 4.604 (0.002)
10 0.024 (<0.001<0.001) 0.169 (0.006) 7.749 (0.007) 0.015 (<0.001<0.001) 0.059 (0.003) 5.921 (0.007) 0.01 (<0.001<0.001) 0.037 (0.002) 4.139 (0.004)
Eigen Centrality 3 1 0.009 (0.001) 0.031 (0.002) 0.331 (0.005) 0.007 (<0.001<0.001) 0.028 (0.002) 0.257 (0.005) 0.007 (0.001) 0.028 (0.002) 0.164 (0.004)
2 0.009 (0.001) 0.031 (0.002) 0.292 (0.004) 0.007 (<0.001<0.001) 0.029 (0.002) 0.208 (0.004) 0.007 (0.001) 0.025 (0.002) 0.103 (0.002)
5 0.009 (0.001) 0.036 (0.002) 0.193 (0.004) 0.007 (<0.001<0.001) 0.028 (0.002) 0.115 (0.004) 0.007 (0.001) 0.023 (0.002) 0.044 (0.001)
10 0.009 (0.001) 0.037 (0.003) 0.063 (0.004) 0.007 (<0.001<0.001) 0.029 (0.002) 0.041 (0.002) 0.007 (0.001) 0.025 (0.002) 0.1 (0.003)
6 1 0.012 (0.001) 0.037 (0.003) 0.466 (0.007) 0.009 (0.001) 0.028 (0.002) 0.409 (0.007) 0.008 (0.001) 0.022 (0.002) 0.328 (0.004)
2 0.012 (0.001) 0.037 (0.003) 0.366 (0.006) 0.009 (0.001) 0.032 (0.002) 0.308 (0.006) 0.008 (0.001) 0.024 (0.002) 0.228 (0.005)
5 0.012 (0.001) 0.04 (0.003) 0.247 (0.005) 0.009 (0.001) 0.031 (0.002) 0.177 (0.005) 0.008 (0.001) 0.025 (0.002) 0.099 (0.003)
10 0.012 (0.001) 0.041 (0.003) 0.144 (0.005) 0.009 (0.001) 0.035 (0.002) 0.08 (0.004) 0.008 (0.001) 0.022 (0.002) 0.048 (0.001)
Harmonic Centrality 3 1 21.91 (0.115) 23.201 (0.742) 1036.187 (0.638) 2.046 (0.025) 4.445 (0.192) 884.79 (0.37) 0.738 (0.009) 3.886 (0.23) 784.981 (0.335)
2 21.91 (0.115) 20.024 (0.664) 976.434 (0.83) 2.046 (0.025) 4.301 (0.18) 836.624 (0.756) 0.738 (0.009) 3.722 (0.21) 747.901 (0.597)
5 21.91 (0.115) 10.857 (0.554) 732.359 (1.592) 2.046 (0.025) 3.298 (0.155) 621.66 (1.79) 0.738 (0.009) 3.381 (0.189) 571.464 (1.537)
10 21.91 (0.115) 6.869 (0.295) 417.292 (1.569) 2.046 (0.025) 2.374 (0.101) 324.391 (1.682) 0.738 (0.009) 3.109 (0.165) 317.947 (1.688)
6 1 23.68 (0.118) 24.162 (0.659) 1041.142 (0.376) 1.953 (0.018) 4.435 (0.177) 898.208 (0.168) 0.765 (0.009) 3.337 (0.176) 796.937 (0.139)
2 23.68 (0.118) 21.104 (0.646) 1026.115 (0.419) 1.953 (0.018) 3.904 (0.162) 885.7 (0.286) 0.765 (0.009) 3.116 (0.149) 787.725 (0.226)
5 23.68 (0.118) 12.597 (0.542) 933.849 (0.746) 1.953 (0.018) 3.362 (0.139) 808.732 (0.812) 0.765 (0.009) 2.683 (0.143) 729.535 (0.517)
10 23.68 (0.118) 7.174 (0.265) 713.174 (1.418) 1.953 (0.018) 2.785 (0.095) 618.093 (1.537) 0.765 (0.009) 2.524 (0.124) 575.932 (1.051)
6.2 Results under the RDPG model
For the RDPG model, we sample latent vectors $Z_{1},\ldots,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Uniform}[0,1]^{d}$. The $Z_{i}$’s are then rescaled to ensure the desired network density. The other steps remain the same as in the previous experiment. Tables 3 and 4 present the experimental results under the RDPG model measured by the Wasserstein distance, corresponding to $n=2000$ and $n=4000$, respectively. The observed patterns align closely with those seen under the inner product models. Our proposed GRAND method clearly outperforms the naive Laplace mechanism across all settings. For larger privacy budgets $\varepsilon$, the preservation of network properties becomes comparable to that of the Hat mechanism. It is worth noting that, overall, model estimation under the RDPG framework becomes noticeably noisier (than the inner product model) in very sparse networks, resulting in less discernible trends with respect to $\varepsilon$ in many cases. This is likely due to the fact that the RDPG estimation is based on first-order information rather than the likelihood (Athreya et al., 2018). However, in denser scenarios, we still observe the expected improvement in property preservation as the privacy budget increases.
Table 3: Node-level statistic distribution preservation results with $n=2000$ under the RDPG model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.032 (<0.001<0.001) 0.042 (0.001) 2.559 (0.004) 0.017 (<0.001<0.001) 0.031 (0.001) 2.037 (0.003) 0.009 (<0.001<0.001) 0.026 (0.001) 1.511 (0.002)
2 0.032 (<0.001<0.001) 0.043 (0.001) 1.715 (0.007) 0.017 (<0.001<0.001) 0.03 (0.001) 1.408 (0.005) 0.009 (<0.001<0.001) 0.027 (0.001) 1.126 (0.003)
5 0.032 (<0.001<0.001) 0.043 (0.001) 0.521 (0.003) 0.017 (<0.001<0.001) 0.028 (0.001) 0.554 (0.003) 0.009 (<0.001<0.001) 0.023 (0.001) 0.518 (0.003)
10 0.032 (<0.001<0.001) 0.047 (0.001) 0.47 (0.003) 0.017 (<0.001<0.001) 0.028 (0.001) 0.508 (0.002) 0.009 (<0.001<0.001) 0.021 (0.001) 0.484 (0.002)
6 1 0.036 (<0.001<0.001) 0.038 (0.001) 2.921 (0.001) 0.018 (<0.001<0.001) 0.025 (0.001) 2.269 (0.001) 0.01 (<0.001<0.001) 0.021 (0.001) 1.601 (0.001)
2 0.036 (<0.001<0.001) 0.039 (0.001) 2.631 (0.003) 0.018 (<0.001<0.001) 0.024 (0.001) 2.082 (0.002) 0.01 (<0.001<0.001) 0.019 (0.001) 1.486 (0.001)
5 0.036 (<0.001<0.001) 0.039 (0.001) 1.444 (0.006) 0.018 (<0.001<0.001) 0.023 (0.001) 1.179 (0.005) 0.01 (<0.001<0.001) 0.018 (0.001) 0.869 (0.004)
10 0.036 (<0.001<0.001) 0.04 (0.001) 0.516 (0.003) 0.018 (<0.001<0.001) 0.023 (0.001) 0.524 (0.003) 0.01 (<0.001<0.001) 0.017 (0.001) 0.567 (0.002)
V-shape 3 1 0.068 (0.001) 0.088 (0.003) 5.183 (0.008) 0.034 (<0.001<0.001) 0.065 (0.003) 4.106 (0.006) 0.019 (<0.001<0.001) 0.053 (0.002) 3.036 (0.003)
2 0.068 (0.001) 0.091 (0.003) 3.486 (0.014) 0.034 (<0.001<0.001) 0.062 (0.003) 2.841 (0.01) 0.019 (<0.001<0.001) 0.055 (0.002) 2.263 (0.006)
5 0.068 (0.001) 0.091 (0.003) 1.069 (0.006) 0.034 (<0.001<0.001) 0.058 (0.002) 1.129 (0.006) 0.019 (<0.001<0.001) 0.047 (0.002) 1.047 (0.006)
10 0.068 (0.001) 0.098 (0.003) 1.005 (0.006) 0.034 (<0.001<0.001) 0.057 (0.002) 1.069 (0.005) 0.019 (<0.001<0.001) 0.043 (0.002) 1 (0.004)
6 1 0.074 (0.001) 0.078 (0.002) 5.902 (0.002) 0.037 (<0.001<0.001) 0.051 (0.002) 4.567 (0.002) 0.019 (<0.001<0.001) 0.042 (0.002) 3.215 (0.002)
2 0.074 (0.001) 0.08 (0.002) 5.321 (0.005) 0.037 (<0.001<0.001) 0.048 (0.002) 4.191 (0.004) 0.019 (<0.001<0.001) 0.039 (0.002) 2.984 (0.003)
5 0.074 (0.001) 0.081 (0.002) 2.935 (0.013) 0.037 (<0.001<0.001) 0.047 (0.002) 2.378 (0.011) 0.019 (<0.001<0.001) 0.037 (0.002) 1.747 (0.008)
10 0.074 (0.001) 0.084 (0.002) 1.055 (0.007) 0.037 (<0.001<0.001) 0.048 (0.002) 1.063 (0.006) 0.019 (<0.001<0.001) 0.035 (0.002) 1.145 (0.005)
Triangle 3 1 0.075 (0.001) 0.095 (0.004) 8.209 (0.008) 0.037 (0.001) 0.078 (0.004) 6.53 (0.006) 0.019 (<0.001<0.001) 0.066 (0.003) 4.788 (0.004)
2 0.075 (0.001) 0.091 (0.004) 6.208 (0.017) 0.037 (0.001) 0.076 (0.004) 5.105 (0.012) 0.019 (<0.001<0.001) 0.067 (0.003) 3.944 (0.007)
5 0.075 (0.001) 0.085 (0.003) 2.366 (0.016) 0.037 (0.001) 0.066 (0.003) 2.135 (0.01) 0.019 (<0.001<0.001) 0.057 (0.003) 1.853 (0.008)
10 0.075 (0.001) 0.08 (0.003) 1.199 (0.008) 0.037 (0.001) 0.058 (0.003) 1.373 (0.008) 0.019 (<0.001<0.001) 0.051 (0.002) 1.371 (0.006)
6 1 0.127 (0.001) 0.1 (0.003) 8.981 (0.003) 0.052 (0.001) 0.067 (0.003) 7.005 (0.003) 0.023 (<0.001<0.001) 0.055 (0.003) 5.005 (0.003)
2 0.127 (0.001) 0.102 (0.003) 8.312 (0.006) 0.052 (0.001) 0.062 (0.003) 6.583 (0.005) 0.023 (<0.001<0.001) 0.05 (0.003) 4.753 (0.003)
5 0.127 (0.001) 0.092 (0.003) 5.332 (0.018) 0.052 (0.001) 0.058 (0.003) 4.442 (0.014) 0.023 (<0.001<0.001) 0.046 (0.002) 3.37 (0.01)
10 0.127 (0.001) 0.085 (0.002) 2.314 (0.017) 0.052 (0.001) 0.056 (0.002) 2.049 (0.011) 0.023 (<0.001<0.001) 0.042 (0.002) 1.908 (0.008)
Eigen Centrality 3 1 0.051 (0.002) 0.049 (0.002) 0.115 (0.002) 0.031 (0.002) 0.033 (0.002) 0.14 (0.002) 0.019 (0.001) 0.023 (0.002) 0.137 (0.001)
2 0.051 (0.002) 0.05 (0.003) 0.187 (0.002) 0.031 (0.002) 0.033 (0.002) 0.18 (0.001) 0.019 (0.001) 0.021 (0.001) 0.164 (0.001)
5 0.051 (0.002) 0.047 (0.003) 0.299 (0.003) 0.031 (0.002) 0.034 (0.002) 0.284 (0.002) 0.019 (0.001) 0.02 (0.001) 0.209 (0.002)
10 0.051 (0.002) 0.052 (0.003) 0.263 (0.003) 0.031 (0.002) 0.032 (0.002) 0.286 (0.003) 0.019 (0.001) 0.022 (0.002) 0.268 (0.002)
6 1 0.075 (0.003) 0.054 (0.003) 0.303 (0.003) 0.042 (0.002) 0.036 (0.002) 0.272 (0.004) 0.021 (0.001) 0.029 (0.002) 0.237 (0.004)
2 0.075 (0.003) 0.053 (0.003) 0.08 (0.002) 0.042 (0.002) 0.036 (0.002) 0.092 (0.002) 0.021 (0.001) 0.027 (0.002) 0.12 (0.002)
5 0.075 (0.003) 0.054 (0.003) 0.264 (0.002) 0.042 (0.002) 0.036 (0.002) 0.233 (0.002) 0.021 (0.001) 0.028 (0.002) 0.202 (0.001)
10 0.075 (0.003) 0.05 (0.003) 0.337 (0.003) 0.042 (0.002) 0.037 (0.002) 0.339 (0.003) 0.021 (0.001) 0.026 (0.002) 0.278 (0.002)
Harmonic Centrality 3 1 3.548 (0.059) 6.465 (0.34) 394.221 (1.054) 1.628 (0.03) 3.133 (0.157) 332.865 (0.911) 0.83 (0.012) 2.388 (0.111) 331.579 (0.533)
2 3.548 (0.059) 6.846 (0.336) 207.534 (1.267) 1.628 (0.03) 2.99 (0.163) 169.035 (1.21) 0.83 (0.012) 2.484 (0.115) 213.128 (0.867)
5 3.548 (0.059) 6.853 (0.328) 44.747 (0.388) 1.628 (0.03) 2.866 (0.129) 76.293 (0.556) 0.83 (0.012) 2.103 (0.093) 92.078 (0.485)
10 3.548 (0.059) 7.361 (0.323) 52.71 (0.551) 1.628 (0.03) 2.891 (0.145) 81.309 (0.75) 0.83 (0.012) 1.887 (0.081) 67.14 (0.378)
6 1 4.03 (0.037) 5.486 (0.206) 526.587 (0.342) 1.563 (0.02) 2.144 (0.087) 427.069 (0.2) 0.877 (0.012) 1.978 (0.099) 382.348 (0.19)
2 4.03 (0.037) 5.613 (0.229) 419.844 (0.834) 1.563 (0.02) 2.027 (0.096) 352.837 (0.646) 0.877 (0.012) 1.805 (0.08) 335.148 (0.505)
5 4.03 (0.037) 5.68 (0.21) 182.552 (0.893) 1.563 (0.02) 2.018 (0.083) 136.861 (0.998) 0.877 (0.012) 1.687 (0.079) 165.261 (0.983)
10 4.03 (0.037) 6.044 (0.214) 56.276 (0.413) 1.563 (0.02) 2.025 (0.081) 58.349 (0.389) 0.877 (0.012) 1.603 (0.077) 88.231 (0.481)
7 Examples of real-world data sets
In this section, we demonstrate our method on privatizing real-world social network data sets. Similar to the simulation experiments, we focus on demonstrating how well the released network preserves the network properties, while fixing the privacy budget. In addition to the Hat method and the Laplace mechanism, we also use the true local network statistics of the original network as an addition benchmark.
Caltech Facebook network
The first example data set is the Facebook social network between students in California Institute of Technology (Caltech), collected by Traud et al. (2012). Each node in the network is a student and the edges are the Facebook connections between students. We process the data set following Wang and Rohe (2016) and Li et al. (2022) using the 2-core algorithm. The resulting data set contains 734 students with an average degree of 45.29. Similar to our previous experiments, we randomly select half of the students to hold out. The network of the other half of the students ($n=367$) is then privatized and released. The inner product latent space model is used for model fitting, and the latent dimension $d=6$ is selected by the edge cross-validation method of Li et al. (2020a) on the hold-out network. We enforce a strong privacy requirement at the level of $\varepsilon=1$. The distributions of the five local statistics in the released network are calculated and compared with the distributions in the true network.
Table 4: Node-level statistic distribution preservation results with $n=4000$ under the RDPG model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.017 (<0.001<0.001) 0.026 (0.001) 2.496 (0.004) 0.009 (<0.001<0.001) 0.019 (0.001) 2.053 (0.002) 0.005 (<0.001<0.001) 0.02 (0.001) 1.516 (0.001)
2 0.017 (<0.001<0.001) 0.025 (0.001) 1.613 (0.006) 0.009 (<0.001<0.001) 0.021 (0.001) 1.43 (0.004) 0.005 (<0.001<0.001) 0.019 (0.001) 1.138 (0.003)
5 0.017 (<0.001<0.001) 0.023 (0.001) 0.544 (0.002) 0.009 (<0.001<0.001) 0.018 (0.001) 0.535 (0.002) 0.005 (<0.001<0.001) 0.016 (0.001) 0.511 (0.002)
10 0.017 (<0.001<0.001) 0.025 (0.001) 0.504 (0.002) 0.009 (<0.001<0.001) 0.016 (0.001) 0.487 (0.002) 0.005 (<0.001<0.001) 0.015 (0.001) 0.473 (0.001)
6 1 0.019 (<0.001<0.001) 0.021 (0.001) 2.908 (0.001) 0.01 (<0.001<0.001) 0.016 (0.001) 2.259 (0.001) 0.005 (<0.001<0.001) 0.013 (0.001) 1.598 (0.001)
2 0.019 (<0.001<0.001) 0.021 (0.001) 2.562 (0.003) 0.01 (<0.001<0.001) 0.015 (0.001) 2.038 (0.002) 0.005 (<0.001<0.001) 0.013 (0.001) 1.471 (0.001)
5 0.019 (<0.001<0.001) 0.022 (0.001) 1.308 (0.005) 0.01 (<0.001<0.001) 0.015 (0.001) 1.074 (0.004) 0.005 (<0.001<0.001) 0.013 (0.001) 0.819 (0.003)
10 0.019 (<0.001<0.001) 0.021 (0.001) 0.523 (0.002) 0.01 (<0.001<0.001) 0.015 (0.001) 0.562 (0.002) 0.005 (<0.001<0.001) 0.011 (0.001) 0.593 (0.002)
V-shape 3 1 0.034 (<0.001<0.001) 0.054 (0.002) 5.025 (0.008) 0.018 (<0.001<0.001) 0.038 (0.002) 4.121 (0.004) 0.01 (<0.001<0.001) 0.04 (0.002) 3.039 (0.003)
2 0.034 (<0.001<0.001) 0.05 (0.002) 3.253 (0.012) 0.018 (<0.001<0.001) 0.042 (0.002) 2.872 (0.007) 0.01 (<0.001<0.001) 0.038 (0.002) 2.283 (0.005)
5 0.034 (<0.001<0.001) 0.048 (0.002) 1.105 (0.005) 0.018 (<0.001<0.001) 0.037 (0.002) 1.08 (0.004) 0.01 (<0.001<0.001) 0.032 (0.001) 1.027 (0.004)
10 0.034 (<0.001<0.001) 0.051 (0.002) 1.06 (0.004) 0.018 (<0.001<0.001) 0.031 (0.001) 1.005 (0.003) 0.01 (<0.001<0.001) 0.031 (0.001) 0.962 (0.003)
6 1 0.039 (<0.001<0.001) 0.043 (0.001) 5.846 (0.002) 0.02 (<0.001<0.001) 0.032 (0.001) 4.532 (0.001) 0.01 (<0.001<0.001) 0.027 (0.001) 3.202 (0.001)
2 0.039 (<0.001<0.001) 0.042 (0.001) 5.153 (0.005) 0.02 (<0.001<0.001) 0.031 (0.001) 4.09 (0.004) 0.01 (<0.001<0.001) 0.026 (0.001) 2.948 (0.002)
5 0.039 (<0.001<0.001) 0.044 (0.001) 2.638 (0.01) 0.02 (<0.001<0.001) 0.031 (0.001) 2.158 (0.008) 0.01 (<0.001<0.001) 0.026 (0.001) 1.643 (0.006)
10 0.039 (<0.001<0.001) 0.043 (0.001) 1.061 (0.004) 0.02 (<0.001<0.001) 0.03 (0.001) 1.134 (0.003) 0.01 (<0.001<0.001) 0.023 (0.001) 1.192 (0.004)
Triangle 3 1 0.037 (<0.001<0.001) 0.063 (0.003) 8.075 (0.008) 0.019 (<0.001<0.001) 0.047 (0.003) 6.532 (0.004) 0.01 (<0.001<0.001) 0.05 (0.003) 4.788 (0.003)
2 0.037 (<0.001<0.001) 0.058 (0.002) 5.977 (0.014) 0.019 (<0.001<0.001) 0.05 (0.003) 5.117 (0.008) 0.01 (<0.001<0.001) 0.048 (0.003) 3.961 (0.006)
5 0.037 (<0.001<0.001) 0.051 (0.002) 2.254 (0.012) 0.019 (<0.001<0.001) 0.042 (0.002) 2.091 (0.009) 0.01 (<0.001<0.001) 0.04 (0.002) 1.845 (0.006)
10 0.037 (<0.001<0.001) 0.051 (0.002) 1.34 (0.007) 0.019 (<0.001<0.001) 0.035 (0.001) 1.385 (0.005) 0.01 (<0.001<0.001) 0.038 (0.002) 1.358 (0.005)
6 1 0.057 (0.001) 0.054 (0.002) 8.954 (0.002) 0.023 (<0.001<0.001) 0.039 (0.002) 6.993 (0.002) 0.011 (<0.001<0.001) 0.034 (0.002) 5.009 (0.002)
2 0.057 (0.001) 0.056 (0.002) 8.162 (0.006) 0.023 (<0.001<0.001) 0.038 (0.002) 6.501 (0.004) 0.011 (<0.001<0.001) 0.034 (0.002) 4.735 (0.003)
5 0.057 (0.001) 0.054 (0.002) 5.015 (0.014) 0.023 (<0.001<0.001) 0.036 (0.001) 4.226 (0.01) 0.011 (<0.001<0.001) 0.033 (0.002) 3.274 (0.007)
10 0.057 (0.001) 0.046 (0.001) 2.152 (0.01) 0.023 (<0.001<0.001) 0.034 (0.001) 2.036 (0.007) 0.011 (<0.001<0.001) 0.028 (0.002) 1.92 (0.006)
Eigen Centrality 3 1 0.035 (0.002) 0.035 (0.002) 0.142 (0.002) 0.02 (0.001) 0.023 (0.002) 0.125 (0.002) 0.013 (0.001) 0.016 (0.001) 0.133 (0.001)
2 0.035 (0.002) 0.034 (0.002) 0.217 (0.001) 0.02 (0.001) 0.023 (0.001) 0.178 (0.001) 0.013 (0.001) 0.015 (0.001) 0.161 (0.001)
5 0.035 (0.002) 0.035 (0.002) 0.329 (0.002) 0.02 (0.001) 0.022 (0.001) 0.299 (0.001) 0.013 (0.001) 0.017 (0.001) 0.214 (0.001)
10 0.035 (0.002) 0.034 (0.002) 0.297 (0.003) 0.02 (0.001) 0.023 (0.002) 0.305 (0.002) 0.013 (0.001) 0.016 (0.001) 0.277 (0.002)
6 1 0.049 (0.002) 0.038 (0.002) 0.247 (0.003) 0.022 (0.001) 0.024 (0.002) 0.234 (0.003) 0.014 (0.001) 0.023 (0.002) 0.213 (0.003)
2 0.049 (0.002) 0.036 (0.002) 0.093 (0.002) 0.022 (0.001) 0.025 (0.002) 0.117 (0.001) 0.014 (0.001) 0.021 (0.002) 0.137 (0.002)
5 0.049 (0.002) 0.04 (0.003) 0.309 (0.002) 0.022 (0.001) 0.028 (0.002) 0.255 (0.002) 0.014 (0.001) 0.022 (0.002) 0.213 (0.001)
10 0.049 (0.002) 0.039 (0.002) 0.379 (0.002) 0.022 (0.001) 0.023 (0.002) 0.355 (0.002) 0.014 (0.001) 0.021 (0.002) 0.287 (0.002)
Harmonic Centrality 3 1 3.557 (0.062) 6.924 (0.415) 638.38 (2.005) 1.162 (0.018) 2.385 (0.11) 659.59 (1.106) 0.844 (0.01) 3.485 (0.202) 666.958 (0.861)
2 3.557 (0.062) 6.257 (0.351) 278.025 (2.062) 1.162 (0.018) 2.558 (0.123) 345.161 (1.525) 0.844 (0.01) 3.31 (0.175) 438.446 (1.367)
5 3.557 (0.062) 5.938 (0.351) 100.041 (1.132) 1.162 (0.018) 2.3 (0.097) 133.62 (0.664) 0.844 (0.01) 2.769 (0.122) 172.796 (0.717)
10 3.557 (0.062) 6.585 (0.336) 138.617 (1.075) 1.162 (0.018) 1.975 (0.076) 124.544 (0.664) 0.844 (0.01) 2.6 (0.116) 114.992 (0.443)
6 1 3.716 (0.052) 4.905 (0.21) 911.974 (0.594) 1.042 (0.011) 1.67 (0.058) 834.276 (0.415) 0.909 (0.012) 2.489 (0.113) 761.823 (0.306)
2 3.716 (0.052) 4.798 (0.252) 671.45 (1.439) 1.042 (0.011) 1.617 (0.057) 666.549 (1.167) 0.909 (0.012) 2.435 (0.114) 659.242 (0.767)
5 3.716 (0.052) 5.245 (0.284) 218.088 (1.249) 1.042 (0.011) 1.596 (0.051) 239.014 (1.211) 0.909 (0.012) 2.404 (0.106) 316.235 (1.292)
10 3.716 (0.052) 5.271 (0.241) 74.745 (0.494) 1.042 (0.011) 1.528 (0.063) 120.293 (0.574) 0.909 (0.012) 2.084 (0.109) 170.016 (0.804)
Figure 2 displays the resulting distributions of the five local statistics in four networks as introduced in Section 6: the original network (True), the network released from our method (GRAND), the network generated using the naive Laplace mechanism (Laplace) and the non-private network from standard model estimation (Hat). It can be seen that the privatized network from our method matches the true network well in all of the five metrics. It also substantially outperforms the naive Laplace method, which completely misses the pattern. Meanwhile, GRAND maintains a slightly deviated but similar performance compared to the non-private Hat network. In particular, the harmonic closeness of the original network exhibits a bi-modal pattern, with one tiny lower mode (on the left end of the figure panel). This subtle pattern is also well captured by our method.
Figure 2: The distributions of five local statistics of the privatized Caltech social network with privacy budget $\varepsilon=1$, compared with those in the true network and the non-private “Hat” network.
Statisticians’ collaboration network
The second example data set is the collaboration network between statistical researchers based on publication data in four statistical journals during the period of 2004–2013, originally collected by Ji and Jin (2016). Here we use the processed version of the data set analyzed in Li et al. (2020b). Each node in this network is a statistician and an edge indicates that the two statisticians coauthored at least one paper during the data collection period. We use the same procedure to process the data set as before, and the resulting network has 509 nodes with an average degree of 4.24. Note that this network is much sparser than the Caltech network, indicating a more difficult model fitting. We still hold out half of the nodes and privatize the network structure of the other half using privacy budget $\varepsilon=1$. The inner product latent space model is used and the latent dimension $d=4$ is selected by cross-validation on the hold-out data.
The distributions of the five local statistics in the resulting networks are shown in Figure 3. The advantage of our method is still evident. The released network matches the true network reasonably well, and maintains a performance that is very close to that of the Hat network, with a small deviation due to the incorporation of privacy guarantees. Similar to the previous Caltech network, the true network also exhibits a bi-modal pattern for the harmonic centrality. However, compared with the previous example, the two modes in this case are much closer to each other, which significantly increases the difficulty in preserving them when perturbation is introduced for privacy. The privatized network from our method, though does not perfectly recover the magnitudes, still captures the bi-modal pattern. We consider this as an impressive success, especially under such a small privacy budget.
Figure 3: The distributions of five local properties of the privatized statisticians’ coauthorship network with privacy budget $\varepsilon=1$, compared with those in the true network and the non-private “Hat” network.
Additional evaluations on social network data
We also evaluate our methods on 107 social networks from the CommunityFitNet data set introduced in Ghasemian et al. (2019). Our method demonstrates competitive performance, and uniformly outperforms the naive Laplace mechanism. For brevity, we include the summary results in Appendix G.
8 Discussion
This paper defines node-level differential privacy and introduces a novel privatization mechanism, named GRAND, to achieve node-level differential privacy. Our work combines data privacy with a novel node-wise estimation method, yielding a mechanism to release network structures under node-level differential privacy while provably preserving original network properties asymptotically for a general class of network models. Notably, this is the first computationally feasible node DP mechanism for network release that includes guarantees for property preservation. Numerically, we demonstrate its effectiveness in preserving the distributions of multiple local statistics in both synthetic and real-world data sets.
It is important to note that our mechanism offers a flexible framework, where various components can be substituted with alternative options. For instance, the choice of network estimation method for the hold-out network, the criterion for node-wise estimation, and the DIP mechanism for introducing noise can be adapted as appropriate. Thus, while laying a foundational cornerstone for solving the node DP problem, this framework also opens numerous avenues for future research and customization. One promising direction is to generalize the proposed method so that it can encode more systematic privacy protection for the hold-out network. In fact, as previously discussed, using such hold-out data currently appears unavoidable to achieve the desired privacy guarantees within the current scope of generality. One viable solution is to adopt a $(\varepsilon,\delta)$-differentially private holdout set (e.g., through kernel density estimation) with theoretical underpinnings and a transfer learning model as discussed in Section 4.2. Yet another potentially useful future direction is to extend the current framework to accommodate graphon models, in which $Z_{i}$’s are univariate with the known uniform distribution but $W$ is unknown.
Appendix
The appendix includes proofs of the theoretical results and additional numerical experiments for the paper.
Appendix A Proof of Theorem 1
Proof.
Due the node-wise estimation procedure, note that the change of any node $i\in[n]$ does not change the estimate of $\hat{Z}_{j},j\neq i,j\in[n]$. That means, the estimates $\hat{Z}_{i}$’s are separable. By applying Theorem 2 of Bi and Shen [2023], $\tilde{Z}_{i}$ acquired via Equation (8) is differentially private. Since the generative function $W$ is assumed to be known. We have each node of $\tilde{A}^{11}$ also being differentially private. ∎
Appendix B Proof of Theorem 2 (Marginal Convergence)
The DIP procedure, given $\hat{Z}_{1}$, is to apply univariate DIP transformation by columns. In each step, we use the approximated (conditional) CDF, $\hat{F}_{m}^{j|1:(j-1)}$ of the $j$th variable, conditioning on the previous various j1j-1 variables, for privatization: for each $i$, we generate $e_{i}$ from $\text{Laplace}(0,1/\epsilon)$ independently, and then compute
$\tilde{Z}_{1ij}=(\hat{F}_{m}^{j|1:(j-1)})^{-1}(G(\hat{F}_{m}^{j|1:(j-1)}(\hat{Z}_{1ij})+e_{i})).$
To prove the consistency for the latent distribution, we need the following key steps: 1) Prove the uniform convergence of the estimated $\hat{F}$ and its inverse based on $\{\hat{Z}_{i}\}_{i>n}$; 2) Prove the convergence of marginal distribution of each $\tilde{Z}_{i}$. The potential challenges are from the fact that $\{\hat{Z}_{i}\}_{i>n}$ are not precisely $\{Z_{i}\}_{i>n}$ and they are not independent. Moreover, each $\hat{Z}_{i},i\leq n$ is also dependent on $\{\hat{Z}_{i}\}_{i>n}$. Therefore, the analysis of the convergence of these quantities have to take these dependence into consideration.
B.1 Crucial tools
The following lemma provides a crucial tool in our main proof.
Lemma B.1 (Uniform Convergence of Empirical CDF with Perturbed Data).
Let $Q=\{Q_{i.}\}_{i=1}^{m}$ be i.i.d. random variables from a distribution $F$ on $\bf R$}, and let $F_{m}(x)$ be the empirical CDF computed from $Q$. Let $\hat{Q}=\{\hat{Q}_{i.}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $\hat{F}_{m}(x)$ denote the empirical CDF based on the perturbed data $\hat{Q}$.
Assume the following regularity conditions:
  • The distribution $F$ is a continuous distribution with density $f$, where $f$ is bounded by a constant $C_{\mathrm{up}}>0$, i.e., $f(y)\leq C_{\mathrm{up}}$.
  • The perturbed data satisfies $\sup_{1\leq i\leq m}|\hat{Q}_{i.}-Q_{i.}|\leq\delta_{m}$ for some $\delta_{m}$ that is associated with $m$.
Then, the empirical CDF $\hat{F}_{m}(x)$, satisfies the following uniform convergence bound:
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq O_{P}\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{1}{m}}\right).$
Proof.
We decompose the total error as follows:
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq\underbrace{\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F_{m}(x)|}_{\text{Perturbation Error}}+\underbrace{\sup_{x\in{\bf R}}|F_{m}(x)-F(x)|}_{\text{Sampling Error}}.$
From Glivenko-Cantelli, we have:
$\sup_{x\in{\bf R}}|F_{m}(x)-F(x)|=O_{P}\left(\sqrt{\frac{1}{m}}\right).$
For any $x\in{\bf R}$, the empirical CDFs are defined as:
$F_{m}(x)=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}_{\{Q_{i}\leq x\}},\quad\hat{F}_{m}(x)=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i}\leq x\}}.$
The difference between $\hat{F}_{m}(x)$ and $F_{m}(x)$ arises when $\mathbf{1}_{\{\hat{Q}_{i}\leq x\}}\neq\mathbf{1}_{\{Q_{i}\leq x\}}$. Define
$D_{i}(x)=\left|\mathbf{1}_{\{\hat{Q}_{i}\leq x\}}-\mathbf{1}_{\{Q_{i}\leq x\}}\right|.$
We have
$|\hat{F}_{m}(x)-F_{m}(x)|\leq\frac{1}{m}\sum_{i=1}^{m}D_{i}(x).$
Since $|\hat{Q}_{i}-Q_{i}|\leq\delta_{m}$, $D_{i}(x)$ can be non-zero only if $Q_{i}\in H(x)$, where $H(x)$ is the interval:
$H(x)=\left\{y\in{\bf R}:|y-x|\leq\delta_{m}\right\}.$
Therefore,
$D_{i}(x)\leq\mathbf{1}_{H(x)}(Q_{i}).$
We have
$E\left[D_{i}(x)\right]\leq P\left(Q_{i}\in H(x)\right)\leq C_{\mathrm{up}}\cdot(2\delta_{m})$
which is uniform in $x$, leading to
$\sup_{x\in{\bf R}}E\left[D_{i}(x)\right]\leq C_{\mathrm{up}}\cdot(2\delta_{m}).$
For a universal control, we call the results from empirical process theory for VC classes. That is, we have:
$E\left[\sup_{x\in{\bf R}}\left|\frac{1}{m}\sum_{i=1}^{m}D_{i}(x)-E[D_{i}(x)]\right|\right]\leq K\sqrt{\frac{1}{m}}.$
Thus,
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F_{m}(x)|\leq\sup_{x}E[D_{i}(x)]+O_{P}\left(\sqrt{\frac{1}{m}}\right)=O\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{p}{m}}\right).$
Therefore,
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq O\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{1}{m}}\right).$
Corollary B.1.
Under the conditions of Lemma B.1, if $\delta_{m}\to 0$ as $m\to 0$, then the approximate $\hat{F}_{m}$ based on $\hat{Q}_{1},\cdots,\hat{Q}_{m}$ in our problem satisfies that $\hat{F}_{m}$ converges to $F$ uniformly in probability:
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
The next tool we need is about a conditional distribution. In particular, suppose a $d$-dimensional random variable $X=(X_{1},\ldots,X_{d})^{\prime}\sim F$, we aim to estimate the conditional CDF
$F^{\,d\mid 1:(d-1)}(x\,\big{|}\,x_{1},\dots,x_{d-1})\;=\;P\!\bigl{(}X_{d}\leq x\,\big{|}\,X_{1}=x_{1},\dots,X_{d-1}=x_{d-1}\bigr{)}.$
This is because we have to estimate the conditional CDF when applying the DIP procedure by following the probability chain rule. In the following result, we will use a kernel estimator. Let $K:{\bf R}^{d-1}\to[0,\infty)$ be a bounded, continuous kernel with $\int_{{\bf R}^{d-1}}K(u)\,du=1$, and let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\,\big{|}\,x_{1},\dots,x_{d-1})\;:=\;\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\Bigr{)}}.$
We assume that the denominator is nonzero, or use a small regularizing constant if needed. Then the following result can be seen as a generalization of Lemma B.1.
Lemma B.2 (Uniform Convergence of Kernel-based Conditional Empirical CDF with Perturbed Data).
Let $Q=\{Q_{i}\}_{i=1}^{m}$ be i.i.d. $d$-dimensional random vectors from a distribution $F$ on ${\bf R}^{d}$, where $Q_{i}=(Q_{i,1},\dots,Q_{i,d})$. Let $\hat{Q}=\{\hat{Q}_{i}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $K:{\bf R}^{d-1}\to[0,\infty)$ be a bounded (by a constant $K_{\max}$) and Lipschitz continuous kernel with $\int_{{\bf R}^{d-1}}K(u)\,du=1$. Let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\,\big{|}\,u)\;:=\;\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}},$
where $u=(x_{1},\dots,x_{d-1})$ and $\hat{Q}_{i,1:(d-1)}=(\hat{Q}_{i,1},\dots,\hat{Q}_{i,d-1})$. We assume that the denominator is nonzero for any $u$ in the domain of interest, or use a small regularizing constant if needed.
Assume the following regularity conditions:
  • The joint density $f(x_{1},\dots,x_{d})$ of $F$ is continuous and bounded by a constant $C_{\mathrm{up}}>0$.
  • The marginal density of the conditioning variables, $f_{1:(d-1)}(u)$, is continuous and bounded. Furthermore, there exists a compact set $\mathcal{U}\subset{\bf R}^{d-1}$ such that $\inf_{u\in\mathcal{U}}f_{1:(d-1)}(u)>0$.
  • The conditional CDF $F^{\,d\mid 1:(d-1)}(x\mid u)$ is Lipschitz continuous with respect to $u=(x_{1},\dots,x_{d-1})$.
  • The perturbed data satisfies $\sup_{1\leq i\leq m}\|\hat{Q}_{i}-Q_{i}\|\leq\delta_{m}$ where $\|{\cdot}\|$ is the Euclidean norm.
If $\delta_{m}=o((\log m)^{-c})$ for some constant $c>0$ and we choose $h=(\log m)^{-c}$, we have
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\,\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;-\;F^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}\;=\;o_{p}(1).$
Hence $\widehat{F}_{m}^{\,d\mid 1:(d-1)}$ converges uniformly to $F^{\,d\mid 1:(d-1)}$ in probability over $\mathcal{U}\times{\bf R}$.
Proof.
The perturbation control strategy we use in this proof is very similar to the control of indicator functions in the proof of Lemma B.1. The main difference is that we have additional smoother kernels (which are indeed easier to control).
Define
$\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid x_{1},\dots,x_{d-1})$
to be the same kernel estimator but using $\{Q_{i}\}$, i.e.,
$\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid x_{1},\dots,x_{d-1})\;=\;\frac{\sum_{i=1}^{m}\mathbf{1}\{Q_{i,d}\leq x\}\;K\Bigl{(}\tfrac{x_{1}-Q_{i,1}}{h},\dots,\tfrac{x_{d-1}-Q_{i,d-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\Bigl{(}\tfrac{x_{1}-Q_{i,1}}{h},\dots,\tfrac{x_{d-1}-Q_{i,d-1}}{h}\Bigr{)}}.$
Then for any $(x_{1},\dots,x_{d-1},x)$,
$\bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}-F^{\,d\mid 1:(d-1)}\bigr{|}\;\leq\;\underbrace{\bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}-\widetilde{F}_{m}^{\,d\mid 1:(d-1)}\bigr{|}}_{(\text{I})}\;+\;\underbrace{\bigl{|}\widetilde{F}_{m}^{\,d\mid 1:(d-1)}-F^{\,d\mid 1:(d-1)}\bigr{|}}_{(\text{II})}.$
Next, we discuss the convergence bound for the two terms separately.
Term (II):
Term $(\text{II})$ represents the uniform error of a standard kernel-based conditional CDF estimator. Under the assumptions that $f_{1:(d-1)}(u)$ is continuous and bounded below by a positive constant on $\mathcal{U}$, and $F^{\,d\mid 1:(d-1)}(x\mid u)$ is Lipschitz continuous with respect to $u$, along with the chosen bandwidth $h$ such that $h\to 0$ and $mh^{d-1}\to\infty$, standard nonparametric estimation results (e.g., from Wasserman [2006] for rates of kernel regression) give the following uniform convergence bound:
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\bigl{|}\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)-F^{\,d\mid 1:(d-1)}(x\mid u)\bigr{|}\;=\;O_{P}\!\Bigl{(}(m\,h^{\,d-1})^{-\tfrac{1}{2}}+h\Bigr{)}.$
Term (I):
We now compare the estimator with perturbed data $\{\hat{Q}_{i}\}$ to the same estimator with unperturbed data $\{Q_{i}\}$. For any fixed $(x,u)$, write
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;=\;\frac{A(x,u)}{B(u)},\quad\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;=\;\frac{C(x,u)}{D(u)},$
where
$A(x,u)\;=\;\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\,K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)},\quad B(u)\;=\;\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)},$
$C(x,u)\;=\;\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}Q_{i,d}\leq x\bigr{\}}\,K\!\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)},\quad D(u)\;=\;\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)}.$
Their difference is
$\frac{A(x,u)}{B(u)}-\frac{C(x,u)}{D(u)}\;=\;\frac{\,A(x,u)D(u)-C(x,u)B(u)\,}{\,B(u)D(u)\,}.$
We will show that the numerator |A(x,u)D(u)C(x,u)B(u)||A(x,u)D(u)-C(x,u)B(u)| is of order $O_{P}(m(\delta_{m}+\delta_{m}/h)+\sqrt{m})$ and $B(u),\,D(u)=\Omega_{P}(m)$, giving
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\frac{A(x,u)}{B(u)}-\frac{C(x,u)}{D(u)}\Bigr{|}\;=\;O_{P}\!\Bigl{(}\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\Bigr{)}.$
We analyze the terms:
  • Denominator Control: For $u\in\mathcal{U}$, we are given that $\inf_{u\in\mathcal{U}}f_{1:(d-1)}(u)>0$. By the uniform consistency of kernel density estimators (under the continuity of $f_{1:(d-1)}$, boundedness of $K$, and $h\to 0$, $mh^{d-1}\to\infty$), we have that $D(u)/m\to f_{1:(d-1)}(u)$ uniformly in probability over $\mathcal{U}$. Thus, $D(u)=\Omega_{P}(m)$ uniformly over $\mathcal{U}$. Similarly, $B(u)=\Omega_{P}(m)$ uniformly over $\mathcal{U}$ (as the perturbation $\delta_{m}$ is small, it does not change the order of the sum). Therefore, $B(u)D(u)=\Omega_{P}(m^{2})$.
  • Numerator Control: We write the numerator as:
    A(x,u)D(u)C(x,u)B(u)=(A(x,u)C(x,u))D(u)C(x,u)(B(u)D(u)).A(x,u)D(u)-C(x,u)B(u)=(A(x,u)-C(x,u))D(u)-C(x,u)(B(u)-D(u)).
    We need to bound each part of this expression uniformly.
    $|B(u)-D(u)|=\left|\sum_{i=1}^{m}\left[K\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}-K\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)}\right]\right|.$
    Since $K$ is Lipschitz continuous with constant $L_{K}$, and $\|\hat{Q}_{i}-Q_{i}\|\leq\delta_{m}$: Each summand is bounded by $L_{K}\frac{\|\hat{Q}_{i,1:(d-1)}-Q_{i,1:(d-1)}\|}{h}\leq L_{K}\frac{\|\hat{Q}_{i}-Q_{i}\|}{h}\leq L_{K}\frac{\delta_{m}}{h}$. Summing over $m$ terms, we get:
    $\sup_{u\in\mathcal{U}}|B(u)-D(u)|\leq mL_{K}\frac{\delta_{m}}{h}=O(m\delta_{m}/h).$
    This is a deterministic uniform bound.
    $A(x,u)-C(x,u)=\sum_{i=1}^{m}\left[\mathbf{1}\{\hat{Q}_{i,d}\leq x\}K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})-\mathbf{1}\{Q_{i,d}\leq x\}K(\tfrac{u-Q_{i,1:(d-1)}}{h})\right].$
    Let $\Delta_{i}(x,u)$ be the $i$-th summand. We use the triangle inequality:
    $|\Delta_{i}(x,u)|$ $\leq$ $\underbrace{\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}\,K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})}_{\text{indicator contribution}}$
    $\;+\;\underbrace{\mathbf{1}\{Q_{i,d}\leq x\}\,\bigl{|}K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})-K(\tfrac{u-Q_{i,1:(d-1)}}{h})\bigr{|}}_{\text{kernel contribution}}.$
    For the kernel part: Each term is bounded by $1\cdot L_{K}\frac{\delta_{m}}{h}$. Summing over $m$ terms, this part contributes $O(m\delta_{m}/h)$ to $\sup_{u,x}|A(x,u)-C(x,u)|$.
    For the indicator part: Since $K(\cdot)\leq K_{\max}$, we have
    $\sum_{i=1}^{m}\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}\,K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})\leq K_{\max}\sum_{i=1}^{m}\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}.$
    Let $D_{i}(x)=\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}$. The sum $\sum_{i=1}^{m}D_{i}(x)$ is a random variable that depends only on $x$ and the $d$-th components of the data. From Lemma B.1 (univariate case, taking $d=1$ and $p=1$), we know that and since $\sup_{x\in{\bf R}}E[D_{i}(x)]\leq C_{\mathrm{up}}\cdot(2\delta_{m})=O(\delta_{m})$, we have
    $\sup_{x\in{\bf R}}\left|\sum_{i=1}^{m}D_{i}(x)\right|=O(m\delta_{m})+O_{P}(\sqrt{m}).$
    So the indicator part is bounded by $O_{P}(m\delta_{m}+\sqrt{m})$.
    Combining the two parts, we have
    $\sup_{u\in\mathcal{U},\,x\in{\bf R}}|A(x,u)-C(x,u)|=O_{P}(m\delta_{m}+\sqrt{m}+m\delta_{m}/h).$
    Note that we can similalry see $\sup_{u,x}|C(x,u)|=O_{P}(m)$. Therefore, we get
    $\sup_{u\in\mathcal{U},\,x\in{\bf R}}|A(x,u)D(u)-C(x,u)B(u)|$
    $\leq$ $\sup_{u,x}|A(x,u)-C(x,u)|\sup_{u}|D(u)|+\sup_{u,x}|C(x,u)|\sup_{u}|B(u)-D(u)|$
    $=$ $O_{P}(m^{2}\delta_{m}+m\sqrt{m}+m^{2}\delta_{m}/h).$
Combining the bounds for the numerator and denominator:
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)-\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}=O_{P}\left(\frac{m^{2}\delta_{m}+m\sqrt{m}+m^{2}\delta_{m}/h}{m^{2}}\right)$
$O_{P}\left(\delta_{m}+\frac{\sqrt{m}}{m}+\frac{\delta_{m}}{h}\right)=O_{P}\left(\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\right).$
Given the conditions $\delta_{m}=o(h)$ and $\delta_{m}\to 0$, and $1/\sqrt{m}\to 0$, this term vanishes in probability, i.e., $o_{P}(1)$.
Combining the vanishing rates of Term (I) and Term (II):
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\,\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;-\;F^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}\;=\;O_{P}\!\Bigl{(}(m\,h^{\,d-1})^{-\tfrac{1}{2}}+h\Bigr{)}+O_{P}\left(\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\right).$
Picking $h=(\log m)^{-c}$ and $\delta_{m}=o((\log m)^{-c})$ ensures the vanishing error.
Theorem B.1.
Let $F$ be a continuous and strictly increasing cumulative distribution function (CDF) (or conditional CDF) on $\bf R$}, and let $F^{-1}$ denote its inverse function. Assume the conditions of Corollary B.1 hold. Let $\hat{F}_{m}$ be the smoothed CDF estimator of Corollary B.1. Then, for any closed interval $a,b]\subset(0,1)$, the inverse functions $\hat{F}_{m}^{-1}$ converge uniformly in probability to $F^{-1}$ on $[a,b]$; that is,
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
Moreover, for any closed interval $a,b]\subset(0,1)$, the sequence $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on $[a,b]$; that is, for any $\varepsilon,\eta>0$, there exist $\delta>0$ and $M\in\mathbb{N}$ such that for all $m\geq M$,
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)<\eta.$
Proof.
We aim to show that for any $\varepsilon>0$,
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
Since $F^{-1}$ is continuous and strictly increasing on $(0,1)$, it is uniformly continuous on the closed interval $[a,b]$. Therefore, for any $\varepsilon>0$, there exists $\delta>0$ such that for all $u,v\in[a,b]$,
$|u-v|<\delta\implies|F^{-1}(u)-F^{-1}(v)|<\frac{\varepsilon}{2}.$
Define the event
$\mathscr{A}_{m}=\left\{\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq\delta\right\}.$
Since $\hat{F}_{m}$ converges to $F$ uniformly in probability, we have $P(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On the event $\mathscr{A}_{m}$, for all $x\in{\bf R}$,
$|\hat{F}_{m}(x)-F(x)|\leq\delta.$
In particular, at $x=F^{-1}(u)$,
$|\hat{F}_{m}(F^{-1}(u))-F(F^{-1}(u))|=|\hat{F}_{m}(F^{-1}(u))-u|\leq\delta,$
which implies $\hat{F}_{m}(F^{-1}(u))\in[u-\delta,u+\delta]$.
For $x\leq F^{-1}(u-\delta)$, since $F$ is strictly increasing,
$F(x)\leq u-\delta\implies\hat{F}_{m}(x)\leq F(x)+\delta\leq u-\delta+\delta=u.$
Thus, $\hat{F}_{m}(x)\leq u$ for all $x\leq F^{-1}(u-\delta)$.
Similarly, for $x\geq F^{-1}(u+\delta)$,
$F(x)\geq u+\delta\implies\hat{F}_{m}(x)\geq F(x)-\delta\geq u+\delta-\delta=u.$
Therefore, $\hat{F}_{m}(x)\geq u$ for all $x\geq F^{-1}(u+\delta)$.
By the definition of the quantile function,
$\hat{F}_{m}^{-1}(u)=\inf\{x\in{\bf R}:\hat{F}_{m}(x)\geq u\}.$
From the observations above, it follows that
$\hat{F}_{m}^{-1}(u)\in[F^{-1}(u-\delta),\ F^{-1}(u+\delta)].$
Therefore,
$|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\leq\max\left\{|F^{-1}(u-\delta)-F^{-1}(u)|,\ |F^{-1}(u+\delta)-F^{-1}(u)|\right\}<\frac{\varepsilon}{2}.$
On the event $\mathscr{A}_{m}$, this holds uniformly for all $u\in[a,b]$, so
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\frac{\varepsilon}{2}.$
Thus,
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right)\leq P(A_{m}^{c})\xrightarrow{m\to\infty}0.$
This concludes the proof that
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
For the second part, let $\varepsilon,\eta>0$ be given. Since $F^{-1}$ is continuous on the closed interval $[a,b]$, it is uniformly continuous. Therefore, there exists $\delta>0$ such that for all $u,v\in[a,b]$,
$|u-v|<\delta\implies|F^{-1}(u)-F^{-1}(v)|<\frac{\varepsilon}{3}.$ (11)
From the uniform convergence in probability of $\hat{F}_{m}^{-1}$ to $F^{-1}$ established earlier, we have:
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
This means that for the given $\varepsilon$ and $\eta$, there exists $M\in\mathbb{N}$ such that for all $m\geq M$,
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (12)
Define the event:
$\mathscr{B}_{m}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\frac{\varepsilon}{3}\right\}.$
Then, $P(\mathscr{B}_{m})>1-\frac{\eta}{2}$ for all $m\geq M$.
On the event $\mathscr{B}_{m}$, for any $u,v\in[a,b]$ with $|u-v|<\delta$, we have:
$|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|$ $\leq|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|+|F^{-1}(u)-F^{-1}(v)|+|F^{-1}(v)-\hat{F}_{m}^{-1}(v)|$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
Here, the first and third terms are bounded by $\frac{\varepsilon}{3}$ due to (15), and the middle term is bounded by $\frac{\varepsilon}{3}$ due to the uniform continuity of $F^{-1}$ in (11).
Therefore, on the event $\mathscr{B}_{m}$,
$\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon.$
Thus,
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)\leq P(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
This completes the proof that $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on $[a,b]$. ∎
Next, we introduce the conditional counterpart of Theorem B.1.
Theorem B.2.
Under the conditions of Lemma B.2, assume additionally that:
  • For each fixed $u\in\mathcal{U}$, the conditional CDF $F^{\,d\mid 1:(d-1)}(\cdot|u)$ is strictly increasing.
  • The conditional CDF $F^{\,d\mid 1:(d-1)}(x|u)$ is jointly continuous in $(x,u)$ for $x\in{\bf R}$ and $u\in\mathcal{U}$.
  • Suppose $\delta_{m}=o((\log m)^{-c})$ for some constant $c>0$ and we choose $h=(\log m)^{-c}$ in the kernel.
Then for any closed interval $a,b]\subset(0,1)$, and any compact set $K\subset{\bf R}^{d-1}$, we have the following conclusions.
  1. 1.
    Uniform convergence:
    $\sup_{u_{0}\in K}\sup_{q\in[a,b]}|(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})-(F^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})|\xrightarrow{P}0\quad\text{as }m\to\infty.$
  2. 2.
    Uniform equicontinuity in conditioning variables: For any $\varepsilon,\eta>0$, there exists $\delta>0$ and $M\in\mathbb{N}$ such that for all $m\geq M$, for any $u_{0},u^{\prime}_{0}\in K$ with $\|u_{0}-u^{\prime}_{0}\|<\delta$:
    $P\left(\sup_{q\in[a,b]}|(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})-(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u^{\prime}_{0})|\geq\varepsilon\right)<\eta.$
Proof.
We first decompose the proof into several steps. For notational simplicity, we write $F(\cdot|u_{0})$ for $F^{\,d\mid 1:(d-1)}(\cdot|u_{0})$ and $\hat{F}_{m}(\cdot|u_{0})$ for $\hat{F}_{m}^{\,d\mid 1:(d-1)}(\cdot|u_{0})$. Also, let $F^{-1}(\cdot|u_{0})$ denote $(F^{\,d\mid 1:(d-1)})^{-1}(\cdot|u_{0})$ and $\hat{F}_{m}^{-1}(\cdot|u_{0})$ denote $(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(\cdot|u_{0})$.
Part 1: Uniform Convergence
We aim to show that for any $\varepsilon>0$,
$P\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
Since $F(\cdot|u_{0})$ is strictly increasing for each fixed $u_{0}$, its inverse $F^{-1}(\cdot|u_{0})$ is well-defined. By the joint continuity of $F(x|u_{0})$ in $(x,u_{0})$ and the compactness of $K\times[a,b]$ (when $F^{-1}$ maps this to a compact set), the function $F^{-1}(q|u_{0})$ is uniformly continuous on $a,b]\times K$. Therefore, for any $\varepsilon>0$, there exists $\delta_{q}>0$ such that for all $q_{1},q_{2}\in[a,b]$ and $u_{0}\in K$,
$|q_{1}-q_{2}|<\delta_{q}\implies|F^{-1}(q_{1}|u_{0})-F^{-1}(q_{2}|u_{0})|<\frac{\varepsilon}{2}.$
Define the event
$\mathscr{A}_{m}=\left\{\sup_{u_{0}\in K}\sup_{x\in{\bf R}}|\hat{F}_{m}(x|u_{0})-F(x|u_{0})|\leq\delta_{q}\right\}.$
From Lemma B.2, we have $P(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On $\mathscr{A}_{m}$, for all $u_{0}\in K$ and $x\in{\bf R}$,
$|\hat{F}_{m}(x|u_{0})-F(x|u_{0})|\leq\delta_{q}.$
Taking $x=F^{-1}(q|u_{0})$, for any $q\in[a,b]$,
$|\hat{F}_{m}(F^{-1}(q|u_{0})|u_{0})-F(F^{-1}(q|u_{0})|u_{0})|=|\hat{F}_{m}(F^{-1}(q|u_{0})|u_{0})-q|\leq\delta_{q},$
which implies $\hat{F}_{m}(F^{-1}(q|u_{0})|u_{0})\in[q-\delta_{q},q+\delta_{q}]$.
For all $x\leq F^{-1}(q-\delta_{q}|u_{0})$, since $F(\cdot|u_{0})$ is strictly increasing,
$F(x|u_{0})\leq q-\delta_{q}\implies\hat{F}_{m}(x|u_{0})\leq F(x|u_{0})+\delta_{q}\leq q-\delta_{q}+\delta_{q}=q.$
Similarly, for $x\geq F^{-1}(q+\delta_{q}|u_{0})$,
$F(x|u_{0})\geq q+\delta_{q}\implies\hat{F}_{m}(x|u_{0})\geq F(x|u_{0})-\delta_{q}\geq q+\delta_{q}-\delta_{q}=q.$
Therefore, $\hat{F}_{m}(x|u_{0})\geq q$ for all $x\geq F^{-1}(q+\delta_{q}|u_{0})$.
Recall that
$\hat{F}_{m}^{-1}(q|u_{0})=\inf\{x\in{\bf R}:\hat{F}_{m}(x|u_{0})\geq q\}.$
We can see
$\hat{F}_{m}^{-1}(q|u_{0})\in[F^{-1}(q-\delta_{q}|u_{0}),\ F^{-1}(q+\delta_{q}|u_{0})].$
Therefore,
$|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\leq\max\left\{|F^{-1}(q-\delta_{q}|u_{0})-F^{-1}(q|u_{0})|,\ |F^{-1}(q+\delta_{q}|u_{0})-F^{-1}(q|u_{0})|\right\}<\frac{\varepsilon}{2}.$
On the event $\mathscr{A}_{m}$, this holds uniformly for all $u_{0}\in K$ and $q\in[a,b]$, so
$\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|<\frac{\varepsilon}{2}.$
Part 2: Uniform Equicontinuity
For any given $\varepsilon,\eta>0$, since $F^{-1}(q|u_{0})$ is jointly continuous in $(q,u_{0})$ on the compact set $a,b]\times K$, it is uniformly continuous on this domain. Therefore, there exists $\delta_{u}>0$ such that for any $u_{0},u^{\prime}_{0}\in K$ with $\|u_{0}-u^{\prime}_{0}\|<\delta_{u}$, and any $q\in[a,b]$:
$|F^{-1}(q|u_{0})-F^{-1}(q|u^{\prime}_{0})|<\frac{\varepsilon}{3}.$ (13)
From the uniform convergence in probability of $\hat{F}_{m}^{-1}(\cdot|\cdot)$ to $F^{-1}(\cdot|\cdot)$ in the previous step, we know
$\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\xrightarrow{P}0\quad\text{as }m\to\infty.$
This means that for the given $\varepsilon$ and $\eta$, there exists $M\in\mathbb{N}$ such that for all $m\geq M$,
$P\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (14)
Define the event:
$\mathscr{B}_{m}=\left\{\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|<\frac{\varepsilon}{3}\right\}.$
Then, $P(\mathscr{B}_{m})>1-\frac{\eta}{2}$ for all $m\geq M$.
On the event $\mathscr{B}_{m}$, for any $u_{0},u^{\prime}_{0}\in K$ with $\|u_{0}-u^{\prime}_{0}\|<\delta_{u}$, and for any $q\in[a,b]$, we have:
$|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|$ $\leq|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|$
$\quad+|F^{-1}(q|u_{0})-F^{-1}(q|u^{\prime}_{0})|$
$\quad+|F^{-1}(q|u^{\prime}_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
Here, the first and third terms are bounded by $\frac{\varepsilon}{3}$ due to (14), and the middle term is bounded by $\frac{\varepsilon}{3}$ due to the uniform continuity of $F^{-1}(\cdot|\cdot)$ in (13).
Therefore, on the event $\mathscr{B}_{m}$,
$\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|<\varepsilon.$
Thus,
$P\left(\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|\geq\varepsilon\right)\leq P(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
This completes the proof. ∎
B.2 Latent distribution consistency: one dimensional case
We first consider the one dimensional GRAND when all latent vectors are univariate ($d=1$). We assume that $\hat{F}_{m}$ which satisfies Corollary B.1 and Theorem B.1 is available, and we focus on the $n$ i.i.d latent variables $Z_{i}\sim F$, $i=1,\ldots,n$, for the released network, and $\hat{Z}_{i}$’s are the approximations of $Z_{i}$’s satisfying $\max_{i}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}$ for $\delta_{m}=o(m^{-\tfrac{2\alpha}{2\alpha+(d-1)}})$. Let $\hat{F}_{m}$ be a kernel-smoothed CDF estimator following Lemma B.2 for $d=1$.
Recall that the privatized latent vector in this case is
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)\right)$
where $e_{i}$ is a Laplace random variable for the privacy budget, independent of everything else, and $G$ is the CDF of the distribution of $U(0,1)+\text{Laplace}$. We want to show that $\tilde{Z}_{i}$ follows $F$ asymptotically.
Our proving strategy takes an intermediate random variable
$\bar{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right),$
involving the same $e_{i}$. In our proof, we will first show that $\bar{Z}_{i}$ weakly converges to the correct distribution first, and then prove that $\tilde{Z}_{i}-\bar{Z}_{i}$ converges to zero in probability.
Lemma B.3 (Marginal convergence of $\bar{Z}_{i}$ to $F$).
Under Assumptions A2A4, then for any given $i$ such that $1\leq i\leq n$, we have
$\bar{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty.$
Proof.
First, by Corollary B.1, $\hat{F}_{m}$ converges uniformly to $F$ in probability:
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
Next, for $Z_{i}\sim F$ that is independent of $\hat{F}_{m}$, we have $F(Z_{i})\sim\text{Uniform}(0,1)$. Because $\hat{F}_{m}$ converges uniformly to $F$ in probability, we know that, as $m\to\infty$,
$\hat{F}_{m}(Z_{i})\xrightarrow{P}F(Z_{i})=U_{i}\sim U(0,1).$
Since $e_{i}$ is independent of $\hat{F}_{m}(Z_{i})$, we have
$\hat{F}_{m}(Z_{i})+e_{i}\xrightarrow{d}U_{i}+e_{i}\sim U(0,1)+\mathrm{Laplace}(1/\eta).$
Therefore, the Continuous Mapping Theorem indicates
$G(\hat{F}_{m}(Z_{i})+e_{i})\xrightarrow{d}G(U_{i}+e_{i})\sim U(0,1)$
as $m\to\infty$.
Finally, we want to show that
$\bar{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty,$
where $Y$ has cumulative distribution function $F$.
For this part, we need the result of Theorem B.1, the uniform convergence of $\hat{F}_{m}^{-1}$ to $F^{-1}$ in probability. Note the function $\hat{F}_{m}^{-1}$ and the variable $G(\hat{F}_{m}(Z_{i})+e_{i})$ both depend on $\hat{F}_{m}$, introducing dependence between them. To deal with it, we will show that $\bar{Z}_{i}$ converges in distribution to $Y\sim F$ by introducing an intermediate term $T_{i}=F^{-1}(G(F(Z_{i})+e_{i}))$, which has a distribution $F$. Our strategy is to show that $\bar{Z}_{i}$ is close to $T_{i}$ in probability. We can write:
$\bar{Z}_{i}-T_{i}$ $=$ $\left[\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right]$
$+\left[\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right].$
Denote:
$A_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right),$
$B_{i}=\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right).$
Then,
$\bar{Z}_{i}-T_{i}=A_{i}+B_{i}.$
We aim to prove that
$|A_{i}|=\left|\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right|\xrightarrow{P}0.$
Since $\hat{F}_{m}$ converges to $F$ uniformly in probability, and $Z_{i}$ is independent of $\hat{F}_{m}$, it follows that
$\hat{F}_{m}(Z_{i})\xrightarrow{P}F(Z_{i}).$
That is, for any $\delta>0$,
$P\left(|\hat{F}_{m}(Z_{i})-F(Z_{i})|\geq\delta\right)\xrightarrow{m\to\infty}0.$
As $G$ is continuous on $[0,1]$, it is uniformly continuous. Therefore, for any $\varepsilon^{\prime}>0$, there exists $\delta>0$ such that for all $u,v\in[0,1]$:
$|u-v|<\delta\implies|G(u)-G(v)|<\varepsilon^{\prime}.$
Thus, when $|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta$, we have:
$|G(\hat{F}_{m}(Z_{i})+e_{i})-G(F(Z_{i})+e_{i})|<\varepsilon^{\prime}.$
From Theorem B.1, the sequence $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on any closed interval $a,b]\subset(0,1)$. Specifically, for any $\varepsilon>0$ and $\eta>0$, there exists $\delta^{\prime}>0$ and $M\in\mathbb{N}$ such that for all $m\geq M$:
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)<\eta.$
Given an arbitrary $\varepsilon>0$, choose $\varepsilon^{\prime}=\delta^{\prime}$ corresponding to $\varepsilon$ in the uniform equicontinuity condition of $\hat{F}_{m}^{-1}$, and select $\delta$ corresponding to $\varepsilon^{\prime}$ in the uniform continuity of $G$.
Define the event:
$\mathscr{E}_{m}=\left\{|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon\right\}.$
Since $\hat{F}_{m}(Z_{i})\xrightarrow{P}F(Z_{i})$ and the second event occurs with high probability, $P(\mathscr{E}_{m})\xrightarrow{m\to\infty}1$. Note that we can make this event even stronger as
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon\right\}$
because the convergence of $\hat{F}_{m}$ is uniform.
On the event $\mathscr{E}_{m}$:
$|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\implies|G(u_{1})-G(u_{2})|<\varepsilon^{\prime}.$
Then, since $|G(u_{1})-G(u_{2})|<\varepsilon^{\prime}=\delta^{\prime}$, the uniform equicontinuity of $\hat{F}_{m}^{-1}$ implies:
$|\hat{F}_{m}^{-1}(G(u_{1}))-\hat{F}_{m}^{-1}(G(u_{2}))|<\varepsilon.$
Therefore, on $\mathscr{E}_{m}$, we have, for all $i$ simultaneously,
$|A_{i}|=\left|\hat{F}_{m}^{-1}(G(u_{1}))-\hat{F}_{m}^{-1}(G(u_{2}))\right|<\varepsilon.$
[…]
Next, we prove that
$|B_{i}|=\left|\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right|\xrightarrow{P}0.$
From our earlier result, for any $\varepsilon>0$ and $\eta^{\prime}>0$, there exists $M\in\mathbb{N}$ such that for all $m\geq M$:
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right)<\eta^{\prime}.$ (15)
We know that $U_{i}=G\left(F(Z_{i})+e_{i}\right)\in[0,1]$ and it is independent of $\hat{F}_{m}$.
Select a closed interval $a,b]\subset(0,1)$ such that
$P\left(U_{i}\in[a,b]\right)>1-\eta^{\prime}.$ (16)
On the event $\mathscr{E}_{m}^{\prime}=\left\{U_{i}\in[a,b]\right\}\cap\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\varepsilon\right\}$, we have
$|B_{i}|=\left|\hat{F}_{m}^{-1}(U_{i})-F^{-1}(U_{i})\right|<\varepsilon.$
And therefore,
$P\left(|B_{i}|\geq\varepsilon\right)\leq P(\mathscr{E}_{m}^{{}^{\prime}c})\leq P\left(U_{i}\notin[a,b]\right)+P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right).$
Using (15) and (16), we obtain
$P\left(|B_{i}|\geq\varepsilon\right)<\eta^{\prime}+\eta^{\prime}=2\eta^{\prime}.$
Since $\eta^{\prime}>0$ is arbitrary, we can make $P\left(|B_{i}|\geq\varepsilon\right)$ as small as desired by choosing $\eta^{\prime}$ appropriately and ensuring $m\geq M$. Therefore,
$|B_{i}|\xrightarrow{P}0.$
Combining the bounds on $A_{i}$ and $B_{i}$, we have:
$\bar{Z}_{i}-T_{i}\xrightarrow{P}0\quad\text{as }m\to\infty.$
Lastly, it is easy to see that
$T_{i}=F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\sim F^{-1}(U^{\prime})\sim F,$
where $U^{\prime}\sim\text{Uniform}(0,1)$. Since $\bar{Z}_{i}-T_{i}\xrightarrow{P}0$ and $T_{i}\xrightarrow{d}Y\sim F$, by Slutsky’s theorem,
$\bar{Z}_{i}=T_{i}+(\bar{Z}_{i}-T_{i})\xrightarrow{d}Y\sim F.$
Next, recall again that
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)\right).$
Our goal is to show that $\tilde{Z}_{i}\to F$ in distribution.
Lemma B.4 (Weak Convergence of $\tilde{Z}_{i}$ to $F$).
Suppose the conditions of Lemma B.3 hold. In addition, assume that the perturbation size $\delta_{m}$ satisfies $\delta_{m}\to 0$ as $m\to\infty$, and for the current set of observations $\{Z_{j}\}_{j=1}^{n}$ and their perturbed versions $\{\hat{Z}_{j}\}_{j=1}^{n}$, we have $\sup_{1\leq j\leq n}|\hat{Z}_{j}-Z_{j}|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$. Then for any given $i\in[n]$, $\tilde{Z}_{i}$ weakly converges to $F$ as $m\to\infty$.
Proof.
Our goal is to show that $\tilde{Z}_{i}\xrightarrow{d}Y\sim F$. We achieve this by leveraging Lemma B.3, which states that $\bar{Z}_{i}\xrightarrow{d}Y\sim F$. The remaining task is to prove that $\tilde{Z}_{i}-\bar{Z}_{i}\xrightarrow{P}0$. Specifically, we want to show that for any $\varepsilon>0$, $P\left(|\tilde{Z}_{i}-\bar{Z}_{i}|\geq\varepsilon\right)\xrightarrow{m,n\to\infty}0$.
Let’s define the arguments for $\hat{F}_{m}^{-1}$:
$u_{i}=G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right),\quad v_{i}=G\left(\hat{F}_{m}(Z_{i})+e_{i}\right).$
Then, $|\tilde{Z}_{i}-\bar{Z}_{i}|=|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|$. Our strategy is to show that $|u_{i}-v_{i}|$ is small in probability, and then use the uniform equicontinuity of $\hat{F}_{m}^{-1}$ from Theorem B.1.
We first analyze $|u_{i}-v_{i}|=\left|G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)-G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right|$. Since $G$ is a CDF of a continuous distribution (specifically, $U(0,1)+\text{Laplace}$), it is uniformly continuous on $\bf R$}. For any $\gamma>0$, there exists $\delta_{G,\gamma}>0$ such that for any $x,y\in{\bf R}$:
$|x-y|<\delta_{G,\gamma}\implies|G(x)-G(y)|<\gamma.$ (17)
Now, let’s bound the difference in the arguments of $G$: $|(\hat{F}_{m}(\hat{Z}_{i})+e_{i})-(\hat{F}_{m}(Z_{i})+e_{i})|=|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|$. Since the kernel $K$ is bounded and Lipschitz continuous, $\hat{F}_{m}$ itself is Lipschitz continuous. Let $L_{\hat{F}_{m}}$ be its Lipschitz constant (proportional to $K_{\max}/h$). Then, on the event where $\sup_{1\leq j\leq n}|\hat{Z}_{j}-Z_{j}|\leq\delta_{m}$ (which holds with probability tending to 1 by assumption), we have:
$|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\leq L_{\hat{F}_{m}}|\hat{Z}_{i}-Z_{i}|\leq L_{\hat{F}_{m}}\delta_{m}.$
As $\delta_{m}\to 0$, it follows that $|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\xrightarrow{P}0$. This means that for any $\delta_{G,\gamma}>0$,
$P\left(|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\geq\delta_{G,\gamma}\right)\xrightarrow{m\to\infty}0.$
By (17), this implies $|u_{i}-v_{i}|\xrightarrow{P}0$.
Next, we handle the mapping of $\hat{F}_{m}^{-1}$.
From Theorem B.1, the sequence $\{\hat{F}_{m}^{-1}\}$ is uniformly equicontinuous in probability on any closed interval $a,b]\subset(0,1)$. Specifically, for any $\varepsilon>0$ and $\eta>0$, there exist $\delta^{\prime}>0$ and $M_{1}\in\mathbb{N}$ such that for all $m\geq M_{1}$:
$P\left(\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})|\geq\varepsilon\right)<\eta/3.$ (18)
The arguments to $\hat{F}_{m}^{-1}$ are $u_{i}=G(\hat{F}_{m}(\hat{Z}_{i})+e_{i})$ and $v_{i}=G(\hat{F}_{m}(Z_{i})+e_{i})$. Both $u_{i}$ and $v_{i}$ are random variables in $[0,1]$. We know $v_{i}\xrightarrow{d}U(0,1)$. Therefore, for any $\eta_{0}>0$, we can choose a closed interval $a,b]\subset(0,1)$ such that $P(v_{i}\in[a,b])\geq 1-\eta_{0}/2$. Since $|u_{i}-v_{i}|\xrightarrow{P}0$ (from Step 1), it also follows that $u_{i}\xrightarrow{P}v_{i}$. Consequently, $P(u_{i}\in[a,b])\geq 1-\eta_{0}/2$ for sufficiently large $m$. Thus, for any $\eta_{0}>0$, there exists an interval $[a,b]$ such that for sufficiently large $m$,
$P(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta_{0}.$ (19)
Now, let $\varepsilon>0$ and $\eta>0$ be arbitrary. We conduct the following steps.
  1. 1.
    Choose $\delta^{\prime}$ and $M_{1}$ from (18) for this $\varepsilon$ and $\eta/3$.
  2. 2.
    Choose $\gamma=\delta^{\prime}$ for the uniform continuity of $G$ in (17).
  3. 3.
    Choose $\delta_{G,\gamma}$ corresponding to this $\gamma$.
  4. 4.
    Since $|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\xrightarrow{P}0$, there exists $M_{2}\in\mathbb{N}$ such that for $m\geq M_{2}$,
    $P\left(|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|\geq\delta_{G,\gamma}\right)<\eta/3.$
  5. 5.
    Choose $\eta_{0}=\eta/3$ for (19), which defines the interval $[a,b]$ and ensures $P(u_{i}\in[a,b]\text{ and }v_{i}\in[a,b])\geq 1-\eta/3$ for $m\geq M_{3}$.
Define the event $\mathscr{E}_{m}$ as the intersection of three high-probability events:
$\mathscr{E}_{m}=\left\{|\hat{F}_{m}(\hat{Z}_{i})-\hat{F}_{m}(Z_{i})|<\delta_{G,\gamma}\right\}\cap\left\{u_{i}\in[a,b]\text{ and }v_{i}\in[a,b]\right\}\cap\left\{\sup_{\begin{subarray}{c}q_{1},q_{2}\in[a,b]\\ |q_{1}-q_{2}|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(q_{1})-\hat{F}_{m}^{-1}(q_{2})|<\varepsilon\right\}.$
For $m\geq\max(M_{1},M_{2},M_{3})$, we have $P(\mathscr{E}_{m}^{c})\leq\eta/3+\eta/3+\eta/3=\eta$. On the event $\mathscr{E}_{m}$:
  • We have $|u_{i}-v_{i}|<\gamma=\delta^{\prime}$.
  • Both $u_{i}$ and $v_{i}$ are in $[a,b]$.
  • The uniform equicontinuity of $\hat{F}_{m}^{-1}$ in (18) applies.
Therefore, on $\mathscr{E}_{m}$, we have:
$|\tilde{Z}_{i}-\bar{Z}_{i}|=|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|<\varepsilon.$
Since $P(\mathscr{E}_{m}^{c})\xrightarrow{m\to\infty}0$, we conclude that $|\tilde{Z}_{i}-\bar{Z}_{i}|\xrightarrow{P}0$.
This completes the proof.
B.3 Latent distribution consistency: multidimensional case
Having established the asymptotic distribution of $\tilde{Z}_{i}$ in the case of $d=1$, now we proceed to prove the weak convergence in the multidimensional case. The proof for the multidimensional case is essentially applying the univariate proofs sequentially across variables, following the same DIP procedures using conditional distributions. Therefore, we will explain the details using the two dimensional case.
Theorem B.3.
Let $Z_{i}=(Z_{i1},Z_{i2})$ be i.i.d. $2$-dimensional random vectors from a distribution $F$ on ${\bf R}^{2}$. Assume that the true distribution $F$ has compact support $S\subset{\bf R}^{2}$. Assume the following conditions hold:
  • The conditions of Lemma B.2 apply to both the marginal CDF $F_{1}$ (for $\hat{F}_{1,m}$) and the conditional CDF $F_{2|1}(\cdot|z_{1})$ (for $\hat{F}_{2|1,m}$). This includes properties of relevant densities, kernels (bounded and Lipschitz).
  • The conditions of Theorem B.2 apply to the inverse functions $F_{1}^{-1}$ and $(F_{2|1})^{-1}$, and their estimators. This implies $F_{1}$ and $F_{2|1}(\cdot|z_{1})$ are strictly increasing, and $F_{2|1}(x|z_{1})$ is jointly continuous in $(x,z_{1})$.
  • The perturbation size $\delta_{m}$ satisfies $\delta_{m}=o((\log m)^{-c})$, and $\sup_{1\leq j\leq n}\|\hat{Z}_{j}-Z_{j}\|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$, for some constant $c>0$.
  • For each $k=1,2$, the bandwidth $h_{m}^{(k)}$ for $\hat{F}_{k|1:(k-1),m}$ is chosen as $h_{m}=h_{m}^{(k)}=(\log m)^{-c}$ for the same constant $c>0$.
Then for any given $i\in[n]$, the privatized latent vector $\tilde{Z}_{i}=(\tilde{Z}_{i1},\tilde{Z}_{i2})$ weakly converges to $F$ as $m\to\infty$; that is, $\tilde{Z}_{i}\xrightarrow{d}Y\sim F$.
Proof.
Our goal is to show that the privatized latent vector $\tilde{Z}_{i}=(\tilde{Z}_{i1},\tilde{Z}_{i2})$ weakly converges to the true latent distribution $F$. That is, $\tilde{Z}_{i}\xrightarrow{d}Y\sim F$, where $Y=(Y_{1},Y_{2})$ is a random vector with CDF $F$. We’ll achieve this by demonstrating the convergence of each component in sequence and then combining them for joint convergence.
For notational simplicity within this proof, let $F_{k}(\cdot)$ denote the marginal CDF of $Z_{k}$, and $F_{k|1:(k-1)}(\cdot|u)$ denote the true conditional CDF for the $k$-th dimension given $u=(z_{1},\dots,z_{k-1})$. Similarly, $\hat{F}_{k,m}$ and $\hat{F}_{k|1:(k-1),m}(\cdot|u)$ are their estimators. Inverse functions are denoted with 1-1. Let $S_{k}$ represent the compact support of $Z_{k}$ (derived from the compact support $S$ of $F$), and $S_{1:(k-1)}$ be the compact support for the conditioning variables $(Z_{1},\dots,Z_{k-1})$. Because $F$ has compact support $S\subset{\bf R}^{d}$, all $Z_{ik}$ and $(Z_{i1},\dots,Z_{i,k-1})$ almost surely lie within their compact projected supports. This is crucial for applying uniform convergence results from previous lemmas, which hold over compact domains.
Step 1 – convergence of the first coordinate $\tilde{Z}_{i1}$:
The first component, $\tilde{Z}_{i1}$, is defined as:
$\tilde{Z}_{i1}=\hat{F}_{1,m}^{-1}\left(G\left(\hat{F}_{1,m}(\hat{Z}_{i1})+e_{i}^{(1)}\right)\right)$
Directly from Lemma B.4, we have:
$\tilde{Z}_{i1}\xrightarrow{d}Y_{1}\sim F_{1}\quad\text{as }m\to\infty$
Step 2 – convergence of the second coordinate $\tilde{Z}_{i12}$ (conditioning on $\tilde{Z}_{i1}$):
We now analyze the conditional distribution of $\tilde{Z}_{i2}$ given $\tilde{Z}_{i1}=u_{1}$, for any fixed $u_{1}\in S_{1}$ (the compact support of $Y_{1}$). The quantity of interest is $P(\tilde{Z}_{i2}\leq x_{2}|\tilde{Z}_{i1}=u_{1})$. For this analysis, we treat $u_{1}$ as a deterministic conditioning variable. The definition of $\tilde{Z}_{i2}$ involves conditioning on $\tilde{Z}_{i1}$, so we consider the variable:
$\tilde{Z}_{i2}(u_{1})=\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
Note that $\hat{Z}_{i1}$ is still a random variable in the argument of $\hat{F}_{2|1,m}$. To properly apply the logic of Lemma B.4 for fixed conditioning, we will use the user-proposed intermediate variables where the outer conditioning is fixed to $u_{1}$, but the inner conditioning remains consistent with the true variables $Z_{i1}$. Let’s define two intermediate variables for comparison, where $u_{1}$ is the fixed outer conditioning value:
$\bar{Z}_{i2}(u_{1})=\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
$T_{i2}(u_{1})=F_{2|1}^{-1}\left(G\left(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}$
Our goal is to show $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{P}0$ and $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$.
To show $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$, we directly apply the logic from Lemma B.4 to the conditional setting. For a fixed $u_{1}$, the functions $\hat{F}_{2|1,m}(\cdot|u_{1})$ and $F_{2|1}(\cdot|u_{1})$ are well-defined. Let $q_{C}=G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)})$ and $q_{D}=G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)})$. We want to show $|\hat{F}_{2|1,m}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{D}|u_{1})|\xrightarrow{P}0$. This splits into:
$|\hat{F}_{2|1,m}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{C}|u_{1})|+|F_{2|1}^{-1}(q_{C}|u_{1})-F_{2|1}^{-1}(q_{D}|u_{1})|.$
  • Term 1: By Theorem B.2, $\sup_{u\in S_{1},q\in[a,b]}|\hat{F}_{2|1,m}^{-1}(q|u)-F_{2|1}^{-1}(q|u)|\xrightarrow{P}0$. Since $u_{1}$ is fixed in $S_{1}$ and $q_{C}$ is a random variable in $[0,1]$, this term goes to 0 in probability, using the same type of proof as in Theorem B.2.
  • Term 2: By uniform continuity of $F_{2|1}^{-1}(\cdot|u_{1})$ w.r.t. its first argument, this term goes to 0 in probability if $|q_{C}-q_{D}|\xrightarrow{P}0$. $|q_{C}-q_{D}|=|G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)})-G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i}^{(2)})|$. By uniform continuity of $G$, this goes to 0 if $|\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})-F_{2|1}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$. This is true by Lemma B.2.
Combining the previous two results, we have $\bar{Z}_{i2}(u_{1})-T_{i2}(u_{1})\xrightarrow{P}0$.
On the other hand, the term $\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})\xrightarrow{P}0$ involves changes in the inner argument’s conditioning variable ($\hat{Z}_{i1}$ to $Z_{i1}$) and the value being fed into $G$ ($\hat{Z}_{i2}$ to $Z_{i2}$).
$|\tilde{Z}_{i2}(u_{1})-\bar{Z}_{i2}(u_{1})|=\left|\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}-\hat{F}_{2|1,m}^{-1}\left(G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|u_{1}\Big{)}\right|$
Let $q_{A}=G(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)})$ and $q_{B}=G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)})$. So we want to control
$|\hat{F}_{2|1,m}^{-1}(q_{A}|u_{1})-\hat{F}_{2|1,m}^{-1}(q_{B}|u_{1})|.$
By the uniform equicontinuity of $\hat{F}_{2|1,m}^{-1}(\cdot|u_{1})$ with respect to its first argument from Theorem B.2, this term will converge to 0 in probability if $|q_{A}-q_{B}|\xrightarrow{P}0$.
To see this, note that
$|q_{A}-q_{B}|=\left|G\left(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i}^{(2)}\right)-G\left(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i}^{(2)}\right)\right|$
By uniform continuity of $G$, it is sufficient to show $|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$. By plugging in an intermediate term and calling the triangle inequality, we have
$|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\leq|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})|+|\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|$
  • Term 1: This term relies on $\hat{F}_{2|1,m}(\cdot|u)$ being Lipschitz continuous with respect to its first argument. So,
    $|\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})|\leq L_{\hat{F}}|\hat{Z}_{i2}-Z_{i2}|=o_{P}(1).$
  • Term 2: This term relies on $\hat{F}_{2|1,m}(x|u)$ being uniformly equicontinuous with respect to its conditioning argument $u$. So, $|\hat{F}_{2|1,m}(Z_{i2}|\hat{Z}_{i1})-\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})|\xrightarrow{P}0$ because $\|\hat{Z}_{i1}-Z_{i1}\|\leq\delta_{m}\to 0$.
These indicate $|q_{A}-q_{B}|\xrightarrow{P}0$.
As discussed, this proves the convergence of $\tilde{Z}_{i12}$ (conditioning on $\tilde{Z}_{i1}$).
Having obtained the convergence in both coordinates, by noticing that the joint PDF is the product of the two CDFs, we get the claimed consistency.
Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
Next, we will show that the empirical CDF of $\tilde{Z}_{i}$’s uniformly converges to $F$ as the proof of Theorem 3.
Again, we first work on the univariate version, $d=1$. And we start with the intermediate random vector $\bar{Z}_{i}$ again. Suppose $d=1$. Define the empirical CDF of $\bar{Z}_{i}$’s as
$\bar{F}_{n}(x)=\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}$
and
$\tilde{F}_{n}(x)=\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}$
Lemma C.1.
Under the assumptions of Lemma B.3, we have
$\sup_{x}|\bar{F}_{n}(x)-F(x)|\xrightarrow{P}0$
as $m,n\to\infty$.
Proof.
The proof will be based on expanding the proof of Lemma B.1 and the proof of Lemma B.3. Use $T_{i}$ as in the proof of Lemma B.3 and define the empirical CDF of $T_{i}$’s as
$F_{n}(x)=\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{T_{i}\leq x\}}.$
The uniform convergence of $F_{n}$ to $F$ is already known. Hence we only focus on controlling the difference between $F_{n}$ and $\bar{F}_{n}$.
$|\bar{F}_{n}(x)-F_{n}(x)|$ $=|\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\frac{1}{n}\sum_{1\leq i\leq n}\mathbf{1}_{\{T_{i}\leq x\}}|$
$\leq\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|.$
Similar to the proof outlined for Lemma B.1, note that $\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}$ and $\mathbf{1}_{\{T_{i}\leq x\}}$ are different only if $x$ lies between $\bar{Z}_{i}$ and $T_{i}$. So for any $\varepsilon>0$, if $|\bar{Z}_{i}-T_{i}|<\varepsilon$, that indicates $x-\varepsilon<T_{i}<x+\varepsilon$, whose probability can be control by $F$ since $T_{i}\sim F$. Specifically, we have
$\sup_{x}|\bar{F}_{n}(x)-F_{n}(x)|$ $\leq\sup_{x}\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{T_{i}\leq x\}}|\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-T_{i}|<\varepsilon}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i}\mathbf{1}_{\{T_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\#\{i:|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}\big{\}}$
$\leq O_{P,n}(\varepsilon)+\frac{1}{n}\sum_{i}\mathbf{1}_{|\{\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}.$ (20)
where $O_{P,n}$ denotes that quantity order with high probability in $n$.
To control the second term, we go back to the proof of Lemma B.3 again in which we have already defined
$A_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right),$
$B_{i}=\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right).$
and thus
$\bar{Z}_{i}-T_{i}=A_{i}+B_{i}.$
Recall that in that context, we take an interval $[a,b]$, and here we specify $a,b$ to be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$. We have defined
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon/2\right\}$
We have seen that $P(\mathscr{E}_{m})\to 1$ and on $\mathscr{E}_{m}$, we have $\sup_{i}|A_{i}|<\varepsilon/2$.
Unfortunately, the other event in the proof, $\mathscr{E}_{m}^{\prime}$ depends on $i$, so we could not achieve the uniform control on $B_{i}$’s. Instead, let us define $\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\varepsilon/2\right\}$. Again, $P(\mathscr{E}_{m}^{\prime})\to 1$ as shown in Lemma B.3.
Therefore, on the event $\mathscr{E}_{m}\cap\mathscr{E}_{m}^{\prime}$, we have
$\frac{1}{n}\sum_{i}\mathbf{1}_{|\{\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i}\notin[a,b]\}}.$
By Hoefding’s inequality, we can further define an event $\mathscr{E}_{m}^{{}^{\prime\prime}}$ such that
$\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i}\notin[a,b]\}}\leq\frac{\varepsilon}{2}$
with probability tending to 1 (with respect to $n$). By union probability, we can see that the event of $\mathscr{E}_{m}\cap\mathscr{E}_{m}^{\prime}\cap\mathscr{E}_{m}^{{}^{\prime\prime}}$ happens with probability going to 1.
Since $\varepsilon$ is arbitrary, because of (C), we have
$\sup_{x}|\bar{F}_{n}(x)-F_{x}(x)|\xrightarrow{P}0.$
Then this convergence can be transferred to $\tilde{Z}_{i}$ with a similar control.
Lemma C.2.
Under the assumptions of Lemma C.1, we further have
$\sup_{x}|\tilde{F}_{n}(x)-F(x)|\xrightarrow{P}0$
as $m,n\to\infty$.
Proof.
With Lemma C.1, we only have to show that
$\sup_{x}|\tilde{F}_{n}(x)-\bar{F}_{n}(x)|\xrightarrow{P}0.$
Calling a similar comparison as before, for any sufficiently small $\varepsilon$,
$\sup_{x}|\bar{F}_{n}(x)-\tilde{F}_{n}(x)|$ $\leq\sup_{x}\frac{1}{n}\sum_{i}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i}\leq x\}}-\mathbf{1}_{\{\tilde{Z}_{i}\leq x\}}|\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i:|\bar{Z}_{i}-\tilde{Z}_{i}|<\varepsilon}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\big{\}}$
$\leq\sup_{x}\big{\{}\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\}}\big{\}}.$ (21)
Note that $\tilde{Z}_{i},i\in[n]$ are independent, because $\hat{F}_{m}$ is based only on $Z_{i},i>n$. For the first term $\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}$, note that $P(Z_{i}\in B(x,\varepsilon))=O(\varepsilon)$ in the current assumption and because of the uniform convergence of $\bar{F}_{n}$ and the independence between $\tilde{Z}_{i}$’s, we have
$\sup_{x}\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i}\in B(x,\varepsilon)\}}=O_{P,n}(\varepsilon).$
To control the second term, we will reuse the quantities in the proof of Lemma B.4. Recall that
$u_{i}=G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right),\quad v_{i}=G\left(\hat{F}_{m}(Z_{i})+e_{i}\right).$
For the current $\epsilon$, pick $a,b$ such that $P(U\sim[a,b])\geq 1-\varepsilon/4$. We have
$\big{\{}i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon\big{\}}$
$\subset$ $\big{\{}i:|\bar{Z}_{i}-\tilde{Z}_{i}|\geq\varepsilon,v_{i},u_{i}\in[a,b]\big{\}}\cup\big{\{}i:v_{i}\notin[a,b]\big{\}}\cup\big{\{}i:u_{i}\notin[a,b]\big{\}}$
$=$ $\big{\{}i:|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|\geq\varepsilon,v_{i},u_{i}\in[a,b]\big{\}}\cup\big{\{}i:v_{i}\notin[a,b]\big{\}}\cup\big{\{}i:u_{i}\notin[a,b]\big{\}}$
Therefore, we have
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}\leq\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\hat{F}_{m}^{-1}(u_{i})-\hat{F}_{m}^{-1}(v_{i})|\geq\varepsilon,v_{i},u_{i}\in[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{v_{i}\notin[a,b]}+\frac{1}{n}\sum_{i}\mathbf{1}_{u_{i}\notin[a,b]}.$
As the proof of Lemma B.4, Theorem B.1 indicates that there exists $\delta>0$ and $M>0$, such that for any $m>M$,
$\sup_{u,v\in[a,b],|u-v|<\delta}|\hat{F}^{-1}_{m}(u)-\hat{F}^{-1}_{m}(v)|<\epsilon$
with probability larger than $1-\epsilon/4.$ Define the event $\mathscr{E}^{{}^{\prime\prime\prime}}$ to be the intersect of this event and the event $\sup_{i}|G(u_{i})-G(v_{i})|<\delta$. Under the current assumptions of $\hat{Z}_{i}$’s concentration and Lemma B.1, we can see that $P(\mathscr{E}^{{}^{\prime\prime\prime}})\to 1.$
The last two terms in the bound would be trivial, as they do not involve the inverse CDF estimate. Specifically, we already know the uniform convergence of $\hat{F}_{m}$ to $F$ under the current assumptions by Lemma B.4. We know that uniformly, we have
$u_{i}-G(F(Z_{i}))\xrightarrow{P}0,v_{i}-G(F(Z_{i}))\xrightarrow{P}0.$
Additionally, since $G(F(Z_{i}))\sim U(0,1)$, we have
$\frac{1}{n}\sum_{i}\mathbf{1}_{v_{i}\notin[a,b]}=O_{P,m,n}(\epsilon),\quad\frac{1}{n}\sum_{i}\mathbf{1}_{u_{i}\notin[a,b]}=O_{P,m,n}(\epsilon).$
Combining these, we have $\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\bar{Z}_{i}-T_{i}|\geq\varepsilon\}}=O(P,m,n)(\epsilon)$.
Because $\epsilon$ is arbitrary, we complete the proof.
Lemma C.3 (Two-dimensional intermediate convergence).
Under the conditions of Lemma C.1 and Lemma B.2,
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\bar{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0$
where
$\bar{F}_{n}(x_{1},x_{2})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}.$
Proof.
Define reference variables $T_{i}=(T_{i1},T_{i2})$ where
$T_{i1}=F_{1}^{-1}(G(F_{1}(Z_{i1})+e_{i1}))$
and
$T_{i2}=F_{2|1}^{-1}(G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i2})|T_{i1})$
Let $F_{n}$ be their empirical CDF:
$F_{n}(x_{1},x_{2})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}.$
By construction, $T_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}F$ , so by the multivariate Glivenko-Cantelli theorem:
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|F_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0.$
Thus we only need to show:
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})|\xrightarrow{P}0.$
For any $(x_{1},x_{2})$:
$|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})|$ $=\Big{|}\frac{1}{n}\sum_{i=1}^{n}[\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}]\Big{|}$
For any $\varepsilon>0$, we can bound this by:
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1$
For any $(x_{1},x_{2})$, when $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we can bound: Note that $|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|\leq|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|$. Moreover, by moving the sup operator from outside the sum to inside the sum, we have:
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})}[|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}[\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$=\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}[\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:|\bar{Z}_{i1}-T_{i1}|<\varepsilon}\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i2}-T_{i2}|<\varepsilon}\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|.$
For each coordinate $j=1,2$, the term $\sup_{x_{j}}|\mathbf{1}_{\{\bar{Z}_{ij}\leq x_{j}\}}-\mathbf{1}_{\{T_{ij}\leq x_{j}\}}|$ equals 1 only if there exists some $x_{j}$ between $\bar{Z}_{ij}$ and $T_{ij}$. When $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we have $|\bar{Z}_{ij}-T_{ij}|<\varepsilon$ for both $j=1,2$.
Using our assumptions about the true $F$, with the same type of probability concentration we have used in previous lemmas, we can see the above bound leads to
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|=O_{P,n}(\varepsilon).$
For the second term, we need to control $\|\bar{Z}_{i}-T_{i}\|$. Note that
$\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\iff|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}$
For the first coordinate, we can directly use the events in Lemma C.1. Recall
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i1})-F_{1}(Z_{i1})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
and
$\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F_{1}^{-1}(u)|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
where $\delta,\delta^{\prime}$ are chosen as in Lemma C.1, and $[a,b]$ contains the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$.
For the second coordinate, conditioning on $\bar{Z}_{i1}$ and $T_{i1}$, define analogous events. Specifically, let $a,b$ be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$. Define
$\mathscr{E}_{m,2}=\left\{\sup_{x_{1}}\sup_{1\leq i\leq n}|\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})-F_{2|1}(Z_{i2}|Z_{i1})|<\delta\right\}\cap\left\{\sup_{x_{1}}\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(v|x_{1})|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
From Lemma B.2 we know that $P(\mathscr{E}_{m,2})\to 1$. Similarly, define
$\mathscr{E}_{m,2}^{\prime}=\left\{\sup_{x_{1}}\sup_{u\in[a,b]}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-F_{2|1}^{-1}(u|x_{1})|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
On the intersection of events $\mathscr{E}_{m}\cap\mathscr{E}_{m}^{\prime}\cap\mathscr{E}_{m,2}\cap\mathscr{E}_{m,2}^{\prime}$, we have for all $i$:
$\{\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\}$ $\subseteq\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\}$
$\subseteq\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\}\cup\{|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\}$
For the first coordinate, on event $\mathscr{E}_{m}\cap\mathscr{E}_{m}^{\prime}$, we have shown in Lemma C.1 that $|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}$ can happen only if $U_{i1}\notin[a,b]$, where $U_{i1}=G(F_{1}(Z_{i1})+e_{i1})$.
For the second coordinate, on event $\mathscr{E}_{m,2}\cap\mathscr{E}_{m,2}^{\prime}$, similarly $|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}$ can happen only if $U_{i2}\notin[a,b]$, where $U_{i2}=G(F_{2|1}(Z_{i2}|Z_{i1})+e_{i2})$. By Hoeffding’s inequality:
$P\left(\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i1}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\to 0$
$P\left(\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i2}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\to 0$
Therefore, with additional intersection of the two events from the Hoeffding’s inequality,
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i}[\mathbf{1}_{|\{U_{i1}\notin[a,b]\}}+\mathbf{1}_{|\{U_{i2}\notin[a,b]\}}]$
$\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\frac{\varepsilon}{2}.$
By union bound, the intersection of all these events has probability approaching 1 as $m,n\to\infty$. Since $\varepsilon$ is arbitrary, we conclude:
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\bar{F}_{n}(x_{1},x_{2})-F_{n}(x_{1},x_{2})|\xrightarrow{P}0$
Lemma C.4 (Two-dimensional privatized convergence).
Under the conditions of Lemma C.3, we have
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\tilde{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0$
where
$\tilde{F}_{n}(x_{1},x_{2})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}.$
Proof.
With Lemma C.3, we only need to show that
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})|\xrightarrow{P}0.$
For any $(x_{1},x_{2})$:
$|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})|$ $=\Big{|}\frac{1}{n}\sum_{i=1}^{n}[\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}]\Big{|}$
For any $\varepsilon>0$, we can bound the difference by:
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}|$
$+\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1$
For the first term, when $\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon$, similar to Lemma C.3, we use the following decomposition
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}[\sup_{x_{1}}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}|+\sup_{x_{2}}|\mathbf{1}_{\{\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}|]$
Using our assumptions about the true $F$ and because of the convergence indicated by Lemma C.3, this term is $O_{P,n}(\varepsilon)$.
For the second term, note that
$\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon\iff|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}$
For the first coordinate, from Lemma C.2, we know that for any $\gamma>0$, there exist events with probability approaching 1 such that
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}\leq\gamma$
For the second coordinate, recall the construction:
$\tilde{Z}_{i2}$ $=\hat{F}_{2|1,m}^{-1}(G(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i2})|\tilde{Z}_{i1})$
$\bar{Z}_{i2}$ $=\hat{F}_{2|1,m}^{-1}(G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i2})|\bar{Z}_{i1})$
Define
$u_{i}=G(\hat{F}_{2|1,m}(\hat{Z}_{i2}|\hat{Z}_{i1})+e_{i2}),\quad v_{i}=G(\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})+e_{i2})$
For controlling $|\tilde{Z}_{i2}-\bar{Z}_{i2}|$, by triangle inequality,
$|\tilde{Z}_{i2}-\bar{Z}_{i2}|$ $=|\hat{F}_{2|1,m}^{-1}(u_{i}|\tilde{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(v_{i}|\bar{Z}_{i1})|$
$\leq|\hat{F}_{2|1,m}^{-1}(u_{i}|\tilde{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(u_{i}|\bar{Z}_{i1})|+|\hat{F}_{2|1,m}^{-1}(u_{i}|\bar{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(v_{i}|\bar{Z}_{i1})|$
For any $\gamma$, since we have already seen the convergence in the first dimension as well as Lemma B.3) and Lemma B.4, we know there exists a compact set K such that:
$P(\frac{1}{n}\sum_{i}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}>\gamma/4)\to 0,\quad P(\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}>\gamma/4)\to 0.$
For $\tilde{Z}_{i1},\bar{Z}_{i1}\in K$, by Theorem B.2, for our fixed $\varepsilon$, there exists $\delta>0$ such that with probability approaching 1:
$\sup_{x_{1}\in K}\sup_{\begin{subarray}{c}y_{1}\in K\\ |x_{1}-y_{1}|<\delta\end{subarray}}\sup_{u\in[a,b]}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(u|y_{1})|<\varepsilon/2\sqrt{2}$
For the difference in probability levels $(u_{i},v_{i})$, by the same theorem on compact set K:
$\sup_{x_{1}\in K}\sup_{u,v\in[a,b],|u-v|<\delta_{1}}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(v|x_{1})|<\varepsilon/2\sqrt{2}$
Therefore, when $\tilde{Z}_{i1},\bar{Z}_{i1}\in K$:
$\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}\subseteq\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\}\cup\{u_{i}\notin[a,b]\}\cup\{v_{i}\notin[a,b]\}\cup\{|u_{i}-v_{i}|\geq\delta_{1}\}$
Combining all terms and using Hoeffding’s inequality, we have
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}$ $\leq\frac{1}{n}\sum_{i}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\}}$
$+\frac{1}{n}\sum_{i}\mathbf{1}_{\{u_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{v_{i}\notin[a,b]\}}+\frac{1}{n}\sum_{i}\mathbf{1}_{\{|u_{i}-v_{i}|\geq\delta_{1}\}}$
$\leq\gamma$
with probability approaching 1 as $m,n\to\infty$.
Combining the above results and picking $\gamma=O(\varepsilon)$, we have
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i}[\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}+\mathbf{1}_{\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}}]=O_{P,m,n}(\varepsilon).$
Because $\varepsilon$ is arbitrary, we complete the proof. ∎
Combining Lemmas C.3 and C.4 gives the claimed result of Theorem 3 in the two dimensional case. Note that the arguments of Lemmas C.3 and C.4 can carried over to any fixed number of dimension $d$, which gives the final theorem.
Appendix D Proof of Theorem 4
D.1 Consistency under inner product latent space models
In this section, we give the proof of the needed concentration bound of $\hat{Z}_{i}$’s. We will prove a high probability error bound for each individual $\hat{Z}_{i}$, $i=1,\cdots,n$. Recall that $\hat{Z}_{i}$’s are identifiable only up to an orthogonal transformation. This indeterminacy does not affect the results of our study, so we ignore it for notational simplicity.
The estimation of $\hat{Z}_{i}$, $i=1,\cdots,n$ is done by solving the problem
$\text{minimize}_{Z_{i},\alpha_{i}}~~\frac{1}{m}\sum_{j=n+1}^{m}-\big{[}A_{ij}\log(\sigma(Z_{i}^{T}\hat{Z}_{j}+\alpha_{i}+\alpha_{j}))+(1-A_{ij})\log(1-\sigma(Z_{i}^{T}\hat{Z}_{j}+\alpha_{i}+\alpha_{j})))\big{]}.$
As discussed, this problem is indeed a logistic regression estimation problem with measurement errors and one offset variable. Therefore, the proof is based on the theory we can get from studying logistic regression.
We first introduce the result of logistic regression as the tool. For clarity, below we first set up a logistic regression problem for our discussion. We consider i.i.d. data $\{(x_{i},y_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$, generated according to the logistic model:
$\operatorname{P}(y_{i}=1\mid x_{i})=\sigma(\beta^{*\top}x_{i}),\quad\text{where }\sigma(z)=\frac{1}{1+e^{-z}}.$
Here, $\beta^{*}\in\mathbb{R}^{d}$ is the true parameter vector. It uniquely minimizes the population (expected) negative log-likelihood:
$\mathcal{L}(\beta)=\mathbb{E}_{(x,y)}[\ell(\beta;x,y)],$
where $\ell(\beta;x,y)$ is the negative log-likelihood for a single data point (using the identity $\log(\sigma(z))=-\log(1+e^{-z})$ and $\log(1-\sigma(z))=-z-\log(1+e^{-z})$):
$\ell(\beta;x,y)$ $=-[y\log(\sigma(\beta^{\top}x))+(1-y)\log(1-\sigma(\beta^{\top}x))]$
$=-[y(\beta^{\top}x)-\log(1+e^{\beta^{\top}x})]$
$=\log(1+e^{\beta^{\top}x})-y(\beta^{\top}x).$ (22)
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta$. Following equation (D.1), we define two empirical average negative log-likelihood functions based on the sample $\{(x_{i},\tilde{x}_{i},y_{i})\}_{i=1}^{n}$:
  • The loss based on the true (unobserved) predictors $x_{i}$:
    $L(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;\,x_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}x_{i}})-y_{i}(\beta^{\top}x_{i})].$
  • The loss based on the observed perturbed predictors $\tilde{x}_{i}$:
    $\widetilde{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\beta;\,\tilde{x}_{i},y_{i})=\frac{1}{n}\sum_{i=1}^{n}[\log(1+e^{\beta^{\top}\tilde{x}_{i}})-y_{i}(\beta^{\top}\tilde{x}_{i})].$
The corresponding estimators (Maximum Likelihood Estimates based on these empirical losses) are:
  • The ”ideal” estimator (uncomputable in practice as $x_{i}$ are unknown):
    $\hat{\beta}^{\,(true)}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}L(\beta).$
  • The actual estimator computed from observed data $(y_{i},\tilde{x}_{i})$:
    $\hat{\beta}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}\widetilde{L}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|_{2}$ of the actual estimator relative to the true parameter $\beta^{*}$. Our theoretical analysis follows the strategy of Wang et al. [2024], with necessary modifications.
First, we need the following regularity conditions.
  1. (C1)
    $x_{i}$ are i.i.d. sub-exponential vectors with bounded Orlicz norm and $E\|{x_{i}}\|^{4}<\infty$.
  2. (C2)
    $\|\beta^{*}\|_{2}\leq R_{\beta}(n)=C_{\beta}\log m$ (constant $R_{\beta}>0$) for another (potentially large integer $m$).
  3. (C3)
    $\mathcal{L}(\beta)$ minimized uniquely at $\beta^{*}$. Hessian $H^{*}:=\nabla^{2}\mathcal{L}(\beta^{*})\succeq\mu I$ ($\mu>0$). Exists fixed $r_{1}>0$ s.t. $\nabla^{2}\mathcal{L}(\beta)\succeq\mu/2I$ for $\beta\in B_{r_{1}}(\beta^{*})$.
  4. (C4)
    $\|\tilde{x}_{i}-x_{i}\|_{2}\leq\delta$.
  5. (C5)
    $1/\sqrt{n}\ll\delta\ll 1$; $\delta=o\big{(}\frac{1}{\log n\log m}\big{)}$.
Lemma D.1.
Under the previous setup and conditions (C1)–(C5), for any desired polynomial decay rate $k>2$, there exists a sufficiently large constant $C_{1}$ and the corresponding $c_{1}$, such that the event $\mathcal{E}_{1}$: $\|\nabla L(\beta^{*})\|_{2}\leq C_{1}\sqrt{(\log n)/n}$ holds with probability at least $1-c_{1}n^{-(k+1)}$ for sufficiently large $n$.
Proof.
Let $g=\nabla L(\beta^{*})\in\mathbb{R}^{d}$. The $j$-th component is $g_{j}=\frac{1}{n}\sum_{i=1}^{n}Z_{ij}$, where $Z_{ij}=\epsilon_{i}x_{ij}=(\sigma(\beta^{*\top}x_{i})-y_{i})x_{ij}$. Since $x_{i}$ is sub-exponential and $\epsilon_{i}$ is bounded ($|\epsilon_{i}|\leq 1$), each $Z_{ij}$ is a mean-zero sub-exponential random variable. Let $K=\sup_{i,j}\|Z_{ij}\|_{\psi_{1}}$ be the sub-exponential parameter (Orlicz $\psi_{1}$-norm), which is $O(1)$. Let $\sigma^{2}=\sup_{i,j}\mathbb{E}[Z_{ij}^{2}]\leq\sup_{i,j}\frac{1}{4}\mathbb{E}[x_{ij}^{2}]=O(1)$.
We apply the scalar Bernstein inequality to each coordinate sum $g_{j}$. There exists a universal constant $C_{B}$ such that for $t>0$:
$\operatorname{P}\left(|g_{j}|\geq t\right)\leq 2\exp\left(-C_{B}n\min\left(\frac{t^{2}}{\sigma^{2}},\frac{t}{K}\right)\right)\leq 2\exp\left(-c^{\prime}n\min(t^{2},t)\right),$
for some constant $c^{\prime}$ depending on $\sigma^{2}$ and $K$. We want a bound $t_{n}$ such that $|g_{j}|\leq t_{n}$ holds for all $j=1,\dots,d$ simultaneously with probability $p_{n}^{\prime}\leq O(n^{-(k+1)})$ for the target $k>2$. Using a union bound:
$\operatorname{P}\left(\max_{j=1,\dots,d}|g_{j}|\geq t_{n}\right)\leq\sum_{j=1}^{d}\operatorname{P}\left(|g_{j}|\geq t_{n}\right)\leq d\cdot 2\exp\left(-c^{\prime}n\min(t_{n}^{2},t_{n})\right).$
Let $C_{1}^{\prime}=\sqrt{(k+1)/c^{\prime}}$. Choosing $t_{n}=C_{1}^{\prime}\sqrt{(\log n)/n}$ ensures the failure probability for a single coordinate is $O(n^{-(k+1)})$. This deviation $t_{n}\to 0$, justifying the use of $\min(t_{n}^{2},t_{n})=t_{n}^{2}$. By the union bound, the event $\mathcal{E}^{\prime}_{1}:=\{\max_{j}|g_{j}|\leq t_{n}\}$ holds with probability $P(\mathcal{E}^{\prime}_{1})\geq 1-O(n^{-(k+1)})$.
Now we bound the L2 norm on this event $\mathcal{E}^{\prime}_{1}$:
$\|\nabla L(\beta^{*})\|_{2}^{2}=\sum_{j=1}^{d}g_{j}^{2}\leq\sum_{j=1}^{d}t_{n}^{2}=d\cdot t_{n}^{2}=d(C_{1}^{\prime})^{2}\frac{\log n}{n}.$
Taking the square root:
$\|\nabla L(\beta^{*})\|_{2}\leq\sqrt{d}C_{1}^{\prime}\sqrt{\frac{\log n}{n}}.$
Let $C_{1}=\sqrt{d}C_{1}^{\prime}$. Thus, event $\mathcal{E}_{1}$ holds with probability $P(\mathcal{E}_{1})\geq 1-O(n^{-(k+1)})$.
Lemma D.2.
Under Assumptions (C1)–(C5), define event $\mathcal{E}_{2}$: $\nabla^{2}L(\beta)\succeq\frac{\mu}{4}I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. For some constant $c_{3}>0$, we have
$P(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$
for sufficiently large $n$.
Proof.
Fix any $\beta$ with $\|\beta-\beta^{*}\|\leq r_{1}$. Write
$X_{i}(\beta)\;=\;\nabla^{2}\ell(\beta;x_{i},y_{i})\;-\;\mathbb{E}\bigl{[}\nabla^{2}\ell(\beta;x_{i},y_{i})\bigr{]}.$
Then $\{X_{i}(\beta)\}_{i=1}^{n}$ are independent, mean‐zero, symmetric $d\times d$ matrices. Moreover
$\nabla^{2}\ell(\beta;x,y)=\sigma^{\prime}(x^{\top}\beta)\,x\,x^{\top},\qquad\sigma^{\prime}(z)\;\leq\;\tfrac{1}{4},$
so
$\|X_{i}(\beta)\|_{\rm op}\;\leq\;\bigl{\|}\nabla^{2}\ell(\beta;x_{i},y_{i})\bigr{\|}_{\rm op}+\bigl{\|}\mathbb{E}[\nabla^{2}\ell(\beta;x_{i},y_{i})]\bigr{\|}_{\rm op}\;=\;O(\|x_{i}\|^{2}).$
Under the sub‐exponential assumption on each coordinate of $x_{i}$, one shows $\|X_{i}(\beta)\|_{\psi_{1}}\leq K$ for some constant $K$. Likewise, defining
$\sigma^{2}(\beta)=\Bigl{\|}\sum_{i=1}^{n}\mathbb{E}\bigl{[}X_{i}(\beta)^{2}\bigr{]}\Bigr{\|}_{\rm op}=O(n).$
Hence by the matrix‐Bernstein inequality (e.g. Vershynin 2018, Thm 6.2.1) there exist constants $c_{1},c_{2}>0$ such that for any $t>0$,
$\Pr\Bigl{(}\bigl{\|}\tfrac{1}{n}\sum_{i=1}^{n}X_{i}(\beta)\bigr{\|}_{\rm op}>t\Bigr{)}\;\leq\;2\,d\;\exp\!\Bigl{(}-c_{1}\,n\,\min\{t^{2}/K^{2},\;t/K\}\Bigr{)}.$
In particular, taking $t=\mu/4$ gives
$\Pr\Bigl{(}\|\nabla^{2}L(\beta)-\mathbb{E}\nabla^{2}\ell(\beta)\|_{\rm op}>\tfrac{\mu}{4}\Bigr{)}\;\leq\;2\,d\,\exp(-c_{2}\,n).$
Now, we proceed to extend the result from a fixed $\beta$ to uniform over the ball using the $\varepsilon$-net construction. Let
$B\;=\;\{\beta\in\mathbb{R}^{d}:\|\beta-\beta^{*}\|\leq r_{1}\}.$
For $\varepsilon>0$ (to be specified), choose an $\varepsilon$-net $\mathcal{N}\subset B$ of size
$|\mathcal{N}|\;\leq\;\Bigl{(}\frac{3r_{1}}{\varepsilon}\Bigr{)}^{d}.$
For each $\beta_{0}\in\mathcal{N}$, we already have
$\Pr\Bigl{(}\bigl{\|}\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\bigr{\|}_{\rm op}>\tfrac{\mu}{4}\Bigr{)}\;\leq\;2\,d\,e^{-c_{2}n}.$
A union bound over $\mathcal{N}$ yields
$\Pr\bigl{(}\exists\,\beta_{0}\in\mathcal{N}:\|\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\|_{\rm op}>\tfrac{\mu}{4}\bigr{)}\;\leq\;|\mathcal{N}|\;2\,d\,e^{-c_{2}n}\;=\;2\,d\,(3r_{1}/\varepsilon)^{d}\,e^{-c_{2}n}.$
We have
$\nabla^{2}\ell(\beta;x,y)=\sigma^{\prime}(x^{\top}\beta)\,x\,x^{\top},$
so for any two parameters $\beta,\beta^{\prime}$,
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=\bigl{[}\sigma^{\prime}(x^{\top}\beta)-\sigma^{\prime}(x^{\top}\beta^{\prime})\bigr{]}\,x\,x^{\top}.$
By the mean‐value theorem, there exists some scalar $\xi$ on the line segment between $x^{\top}\beta$ and $x^{\top}\beta^{\prime}$, such that
$\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)=\sigma^{\prime\prime}(\xi)\,\bigl{(}x^{\top}(\beta-\beta^{\prime})\bigr{)}\,x\,x^{\top}.$
Taking operator norms and using $\|x\,x^{\top}\|_{\rm op}=\|x\|^{2}$, we get
$\bigl{\|}\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)\bigr{\|}_{\rm op}\;\leq\;\bigl{|}\sigma^{\prime\prime}(\xi)\bigr{|}\;\bigl{|}x^{\top}(\beta-\beta^{\prime})\bigr{|}\;\|x\|^{2}\;\leq\;\sup_{u}\bigl{|}\sigma^{\prime\prime}(u)\bigr{|}\;\|x\|\;\|\beta-\beta^{\prime}\|\;\|x\|^{2}.$
Since on $\|\beta-\beta^{*}\|\leq r_{1}$ one has $\xi$ ranging over a compact interval, $\sup_{u}|\sigma^{\prime\prime}(u)|$ is finite. Therefore we can set
$L\;=\;\Bigl{(}\sup_{u}|\sigma^{\prime\prime}(u)|\Bigr{)}\,\|x\|^{3}\;=\;O\bigl{(}\|x\|^{3}\bigr{)},$
and conclude
$\|\nabla^{2}\ell(\beta;x,y)-\nabla^{2}\ell(\beta^{\prime};x,y)\|_{\rm op}\;\leq\;L\,\|\beta-\beta^{\prime}\|.$
With probability at least $1-e^{-c_{3}^{\prime}n}$, we have $\max_{i\leq n}\|x_{i}\|\leq C\log n$. We can take
$L=C^{\prime}(\log n)^{3}$
for some constant $C^{\prime}$. Set
$\varepsilon=\min\!\Bigl{\{}\frac{\mu}{16\,L},\,r_{1}\Bigr{\}}.$
Then for any $\beta\in B$, there exists $\beta_{0}\in\mathcal{N}$ with $\|\beta-\beta_{0}\|\leq\varepsilon$, and
$\|\nabla^{2}L(\beta)-\nabla^{2}L(\beta_{0})\|_{\rm op}\;\leq\;L\,\|\beta-\beta_{0}\|\;\leq\;\frac{\mu}{16}.$
Also note that with this choice of $\varepsilon$, the probability in (i) holds with $O(e^{-c_{3}n})$.
On the intersection of the high‐probability events from (i) and the bound $\max_{i}\|x_{i}\|\leq C\log n$, we have for every $\beta\in B$:
$\|\nabla^{2}L(\beta)-\mathbb{E}\nabla^{2}\ell(\beta)\|_{\rm op}$ $\leq\|\nabla^{2}L(\beta_{0})-\mathbb{E}\nabla^{2}\ell(\beta_{0})\|_{\rm op}+\|\nabla^{2}L(\beta)-\nabla^{2}L(\beta_{0})\|_{\rm op}$
$\leq\frac{\mu}{4}+\frac{\mu}{16}\;<\;\frac{\mu}{2}.$
Since $\mathbb{E}\nabla^{2}\ell(\beta)=\nabla^{2}\mathcal{L}(\beta)\succeq\mu I$, it follows that $\nabla^{2}L(\beta)\succeq\mu I-\tfrac{\mu}{2}I=\tfrac{\mu}{2}I>\tfrac{\mu}{4}I$ uniformly over $B$. Hence
$\Pr(\mathcal{E}_{2}^{c})\;=\;O(e^{-c_{3}n}).$
Lemma D.3.
Under Assumptions (C1)–(C5), for any desired polynomial decay rate $k>2$, there exist constants $C,C_{p}>0$ (depending on $\sigma_{x},R_{\beta},\mu,d,k$), such that for sufficiently large $n$:
$\operatorname{P}\left(\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\delta(\log n)(\log m),\frac{\log n}{\sqrt{n}}\right\}\right)\geq 1-C_{p}n^{-(k+1)}.$
Proof.
Let $k>2$ be the target polynomial decay exponent for the probability. Define high-probability events:
  • $\mathcal{E}_{0}$: Event where $\|x_{i}\|_{2}\leq C_{R}\log n=:R^{\prime}(n)$ for all $i=1..n$. $P(\mathcal{E}_{0})\geq 1-n^{-(k+1)}$.
  • $\mathcal{E}_{1}$: Event where $\|\nabla L(\beta^{*})\|_{2}\leq C_{1}\sqrt{(\log n)/n}$. From Lemma D.1, $P(\mathcal{E}_{1})\geq 1-c_{1}n^{-(k+1)}$.
  • $\mathcal{E}_{2}$: $\nabla^{2}L(\beta)\succeq\mu/4I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$. From Lemma D.2, $P(\mathcal{E}_{2})\geq 1-e^{-c_{3}n}$.
  • $\mathcal{E}_{3}$: $\nabla^{2}\widetilde{L}(\beta)\succeq\mu/8I=:\tilde{\mu}I$ uniformly for $\beta\in B_{r_{1}}(\beta^{*})$.
Independent of any randomness, a simple perturbation check shows
$\bigl{\|}\nabla^{2}\widetilde{L}(\beta)-\nabla^{2}L(\beta)\bigr{\|}_{\rm op}\;\leq\;K\,\delta,$
where $K$ depends only on the $\|x_{i}\|$‐bounds (and thus is $O((\log n)^{2})$ on $\mathcal{E}_{0}$). Hence
$\nabla^{2}\widetilde{L}(\beta)\;\succeq\;\nabla^{2}L(\beta)\;-\;K\delta\,I\;\succeq\;\frac{\mu}{4}\,I\;-\;K\delta\,I.$
By Assumption (C5) we have $K\delta\leq\mu/8$ for large $n$, so
$\nabla^{2}\widetilde{L}(\beta)\;\succeq\;\frac{\mu}{8}\,I\quad\mbox{ for }\forall\|\beta-\beta^{*}\|\leq r_{1},$
which is exactly the event $\mathcal{E}_{3}$. Thus
$\mathcal{E}_{2}\;\subset\;\mathcal{E}_{3}\quad\Longrightarrow\quad P(\mathcal{E}_{3})\;\geq\;P(\mathcal{E}_{2}).$
Let $\mathcal{E}=\mathcal{E}_{0}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$, and we have $P(\mathcal{E})\geq 1-c_{5}n^{-(k+1)}$.
Next, we work under the conditions from $\mathcal{E}$. We want to use Lemma F.3 to show the unique solution. Set $\phi(\beta)=\nabla\widetilde{L}(\beta)$, $x^{*}=\beta^{*}$, $y^{*}=\nabla\widetilde{L}(\beta^{*})$, $y=0$, $\delta_{1}=r_{1}$.
We first verify the conditions of Lemma F.3.
  • Lower bound: For $\|{\beta-\beta^{*}}\|=r_{1}$. Let $H_{avg}(\beta)=\int_{0}^{1}\nabla^{2}\widetilde{L}(\beta^{*}+t(\beta-\beta^{*}))dt$. The segment $\beta^{*},\beta]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}$, $\nabla^{2}\widetilde{L}(\cdot)\succeq\tilde{\mu}I$ in $B_{r_{1}}(\beta^{*})$, thus $H_{avg}(\beta)\succeq\tilde{\mu}I$. Then, $\|{\phi(\beta)-y^{*}}\|=\|{H_{avg}(\beta)(\beta-\beta^{*})}\|\geq\lambda_{\min}(H_{avg}(\beta))\|{\beta-\beta^{*}}\|\geq\tilde{\mu}r_{1}$. Set $\rho=\tilde{\mu}r_{1}$. The condition $\min_{\|{\beta-\beta^{*}}\|=r_{1}}\|{\phi(\beta)-y^{*}}\|\geq\rho$ holds on $\mathcal{E}_{3}$.
  • Upper bound: We need to show $\|{y-y^{*}}\|=\|{0-\nabla\widetilde{L}(\beta^{*})}\|=\|{\nabla\widetilde{L}(\beta^{*})}\|\leq\rho=\tilde{\mu}r_{1}$. We decompose the gradient using the triangle inequality:
    $\|{\nabla\widetilde{L}(\beta^{*})}\|_{2}$ $=\|{\nabla L(\beta^{*})+\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}$
    $\leq\|{\nabla L(\beta^{*})}\|_{2}+\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}.$
    On event $\mathcal{E}_{1}$, the first term is bounded:
    $\|{\nabla L(\beta^{*})}\|_{2}\leq C_{1}\frac{\sqrt{\log n}}{\sqrt{n}}.$
    For the second term, we analyze $\nabla(\widetilde{L}-L)(\beta^{*})=\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})$, where
    $T_{i}(\beta^{*})=-y_{i}(\tilde{x}_{i}-x_{i})+\sigma(\beta^{*\top}\tilde{x}_{i})(\tilde{x}_{i}-x_{i})+[\sigma(\beta^{*\top}\tilde{x}_{i})-\sigma(\beta^{*\top}x_{i})]x_{i}.$
    We bound the norm of $T_{i}(\beta^{*})$ using the triangle inequality:
    $\|T_{i}(\beta^{*})\|_{2}\leq|y_{i}|\|{\tilde{x}_{i}-x_{i}}\|_{2}+|\sigma(\beta^{*\top}\tilde{x}_{i})|\|{\tilde{x}_{i}-x_{i}}\|_{2}+|\sigma(\beta^{*\top}\tilde{x}_{i})-\sigma(\beta^{*\top}x_{i})|\|{x_{i}}\|_{2}.$
    Using bounds $|y_{i}|\leq 1$, $|\sigma(\cdot)|\leq 1$, $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta$ (which is Assumption C4), and the $\frac{1}{4}$-Lipschitz property of $\sigma$:
    $\|T_{i}(\beta^{*})\|_{2}\leq 1\cdot\delta+1\cdot\delta+\frac{1}{4}|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\|{x_{i}}\|_{2}.$
    Applying Cauchy-Schwarz to $|\beta^{*\top}(\tilde{x}_{i}-x_{i})|\leq\|{\beta^{*}}\|_{2}\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\|{\beta^{*}}\|_{2}\delta$, under $\mathcal{E}_{0}$ and C2, we have
    $\|T_{i}(\beta^{*})\|_{2}$ $\leq 2\delta+\frac{1}{4}\|{\beta^{*}}\|_{2}\delta\|{x_{i}}\|_{2}=\left(2+\frac{1}{4}\|{\beta^{*}}\|_{2}\|{x_{i}}\|_{2}\right)$
    $\leq\left(2+\frac{1}{4}R_{\beta}(n)R^{\prime}(n)\right)\delta$
    $\leq\left(2+\frac{1}{4}(C_{\beta}\log m)(C_{R}\log n)\right)\delta$
    $=\left(2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)\right)\delta.$
    Therefore, this leads to
    $\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}$ $=\|{\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta^{*})}\|_{2}\leq\frac{1}{n}\sum_{i=1}^{n}\|T_{i}(\beta^{*})\|_{2}=\left(2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)\right)\delta.$
    Let $C_{G}^{*}(n)=2+\frac{C_{\beta}C_{R}}{4}(\log n)(\log m)=O((\log n)(\log m))$. We have:
    $\|{\nabla(\widetilde{L}-L)(\beta^{*})}\|_{2}\leq C_{G}^{*}(n)\delta.$
    Combining with the bound for $\|{\nabla L(\beta^{*})}\|$ on event $\mathcal{E}_{1}$:
    $\rho_{n}:=\|{\nabla\widetilde{L}(\beta^{*})}\|\leq\frac{C_{1}\log n}{\sqrt{n}}+C_{G}^{*}(n)\delta\quad(\text{holds on }\mathcal{E}_{0}\cap\mathcal{E}_{1}).$
    The condition required for Lemma D.2 is $\rho_{n}\leq\rho=\tilde{\mu}r_{1}$. Substituting the orders:
    $O\left(\frac{\log n}{\sqrt{n}}\right)+O(\delta(\log n)(\log m))\leq\frac{\mu r_{1}}{8}.$
    Since $\delta=o(\frac{1}{(\log n)(\log m)})$ by C5, the inequality holds for sufficiently large $n$.
  • Injectivity: On $\mathcal{E}_{3}$, Jacobian $J_{\phi}(\beta)=\nabla^{2}\widetilde{L}(\beta)\succeq\tilde{\mu}I\succ 0$ in $B_{r_{1}}(\beta^{*})$, implying $\phi$ is injective on $B_{r_{1}}(\beta^{*})$.
All conditions hold on $\mathcal{E}$ for sufficiently large $n$. Therefore, we can apply Lemma F.3 for Localization. On $\mathcal{E}$, Lemma F.3 guarantees the existence of $\hat{\beta}_{soln}\in B_{r_{1}}(\beta^{*})$, such that $\nabla\widetilde{L}(\hat{\beta}_{soln})=0$. Since $\widetilde{L}(\beta)$ is strictly convex on $B_{r_{1}}(\beta^{*})$ (as $\nabla^{2}\widetilde{L}\succeq\tilde{\mu}I>0$ on $\mathcal{E}_{3}$), $\hat{\beta}_{soln}$ is the unique minimizer in $B_{r_{1}}(\beta^{*})$. Global convexity implies that $\hat{\beta}_{soln}$ is the unique global minimum $\hat{\beta}$. Therefore, on $\mathcal{E}$, we have that $\hat{\beta}$ exists, is unique, and satisfies
$\|\hat{\beta}-\beta^{*}\|_{2}\leq r_{1}.$
Next, we derive a bound on $\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}$. We will first show that $\hat{\beta}^{\,(true)}$ must also be in $B_{r_{1}}(\beta^{*})$, under $\mathcal{E}$. Recall that that
$\nabla^{2}L(\beta)\;\succeq\;\frac{\mu}{4}\,I\quad\text{for all }\|\beta-\beta^{*}\|\leq r_{1},\quad\text{and}\quad\|\nabla L(\beta^{*})\|_{2}\;\leq\;C_{1}\sqrt{\frac{\log n}{n}}\,.$
On the event $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, for sufficiently large $n$, we have
$\nabla^{2}L(\beta)\;\succeq\;\frac{\mu}{4}\,I\quad\forall\;\|\beta-\beta^{*}\|\leq r_{1},\qquad\|\nabla L(\beta^{*})\|_{2}\;\leq\;\frac{\mu}{8}\,r_{1}.$
Let $\beta$ satisfy $\|\beta-\beta^{*}\|\geq r_{1}$, and define
$\beta_{0}\;=\;\beta^{*}\;+\;\frac{r_{1}}{\|\beta-\beta^{*}\|}\,(\beta-\beta^{*}),$
so that $\|\beta_{0}-\beta^{*}\|=r_{1}$. By Taylor’s theorem along the segment from $\beta^{*}$ to $\beta_{0}$,
$L(\beta_{0})=L(\beta^{*})+\nabla L(\beta^{*})^{\top}(\beta_{0}-\beta^{*})+\frac{1}{2}\,(\beta_{0}-\beta^{*})^{\top}\Bigl{[}\int_{0}^{1}\nabla^{2}L\bigl{(}\beta^{*}+s(\beta_{0}-\beta^{*})\bigr{)}\,ds\Bigr{]}\,(\beta_{0}-\beta^{*}).$
On the event $\mathcal{E}_{2}$, for every point $\xi$ on that segment we have $\nabla^{2}L(\xi)\succeq\frac{\mu}{4}\,I$, so the averaged Hessian $H_{\rm avg}=\int_{0}^{1}\nabla^{2}L(\beta^{*}+s(\beta_{0}-\beta^{*}))\,ds$ also satisfies $H_{\rm avg}\succeq\frac{\mu}{4}\,I$. Hence
$L(\beta_{0})-L(\beta^{*})\;\geq\;-\|\nabla L(\beta^{*})\|_{2}\,r_{1}\;+\;\tfrac{1}{2}\cdot\tfrac{\mu}{4}\,r_{1}^{2}\;=\;\frac{\mu}{8}\,r_{1}^{2}\;-\;\|\nabla L(\beta^{*})\|_{2}\,r_{1}.$
Since on $\mathcal{E}_{1}$ we have $\|\nabla L(\beta^{*})\|_{2}\leq\frac{\mu}{8}\,r_{1}$, it follows that
$L(\beta_{0})\;\geq\;L(\beta^{*}).$
Define the “outward” vector for $\beta$ (recall that $\|\beta-\beta^{*}\|\geq r_{1}$)
$w\;=\;\beta-\beta_{0}\;=\;\beta^{*}+\frac{\|\beta-\beta^{*}\|}{r_{1}}\,(\beta_{0}-\beta^{*})\;-\;\beta_{0}\;=\;\frac{\|\beta-\beta^{*}\|-r_{1}}{r_{1}}\,(\beta_{0}-\beta^{*}).$
Notice $w$ and $(\beta_{0}-\beta^{*})$ point in exactly the same direction.
Since $L$ is convex and differentiable everywhere, we have
$L(\beta)=L(\beta_{0}+w)\;\geq\;L(\beta_{0})\;+\;\nabla L(\beta_{0})^{\top}w.$
using the supporting hyperplane property of convex functions.
Note that for $v=(\beta_{0}-\beta^{*})/r_{1}$, using the support hyperplane property again, we have
$v^{\top}\nabla L(\beta_{0})\;>\;0.$
But
$w\;=\;\frac{\|\beta-\beta^{*}\|-r_{1}}{r_{1}}\;(\beta_{0}-\beta^{*})\;=\;(\|\beta-\beta^{*}\|-r_{1})\,v,$
so
$\nabla L(\beta_{0})^{\top}w\;=\;(\|\beta-\beta^{*}\|-r_{1})\;v^{\top}\nabla L(\beta_{0})\;>\;0.$
Hence we have
$L(\beta)\;\geq\;L(\beta_{0})\;+\;\nabla L(\beta_{0})^{\top}w\;\;>\;\;L(\beta_{0}).$
Thus no minimizer can lie on or outside the sphere of radius $r_{1}$.
Now, since $\hat{\beta}^{(\mathrm{true})}\in B_{r_{1}}(\beta^{*})$ and $\nabla L(\hat{\beta}^{(\mathrm{true})})=0$, using Taylor expansion of the gradient around $\beta^{*}$, we have:
$0=\nabla L\bigl{(}\hat{\beta}^{(\mathrm{true})}\bigr{)}=\nabla L(\beta^{*})\;+\;\underbrace{\int_{0}^{1}\nabla^{2}L\bigl{(}\beta^{*}+t(\hat{\beta}^{(\mathrm{true})}-\beta^{*})\bigr{)}\,dt}_{=:H_{\rm avg}}\;\bigl{(}\hat{\beta}^{(\mathrm{true})}-\beta^{*}\bigr{)}.$
By construction $H_{\rm avg}\succeq\frac{\mu}{4}I$. Therefore
$\frac{\mu}{4}\;\bigl{\|}\hat{\beta}^{(\mathrm{true})}-\beta^{*}\bigr{\|}_{2}\;\leq\;\bigl{\|}H_{\rm avg}\,(\hat{\beta}^{(\mathrm{true})}-\beta^{*})\bigr{\|}_{2}=\|\nabla L(\beta^{*})\|_{2}\;\leq\;C_{0}\sqrt{\frac{\log n}{n}}.$
Next, we get the bound for $\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}$. Let $\Delta=\hat{\beta}-\hat{\beta}^{\,(true)}$. On $\mathcal{E}$, $\hat{\beta},\hat{\beta}^{\,(true)}\in B_{r_{1}}(\beta^{*})$, so the segment $\hat{\beta}^{\,(true)},\hat{\beta}]\subset B_{r_{1}}(\beta^{*})$. On $\mathcal{E}_{3}\subset\mathcal{E}$, $\nabla^{2}\widetilde{L}(\beta)\succeq\tilde{\mu}I$ on this segment. Let $(\Delta^{\prime})=\widetilde{L}(\hat{\beta}^{\,(true)}+\Delta^{\prime}). ff$ is $\tilde{\mu}-$strongly convex along $0,\Delta]$. Gradient monotonicity gives $\langle\nabla f(\Delta)-\nabla f(0),\Delta-0\rangle\geq\tilde{\mu}\|\Delta\|_{2}^{2}$. With $\nabla f(\Delta)=0$ and $\nabla f(0)=\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})$:
$\langle-\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)}),\Delta\rangle\geq\tilde{\mu}\|\Delta\|_{2}^{2}.$
By Cauchy-Schwarz:
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\|\Delta\|_{2}\geq\tilde{\mu}\|\Delta\|_{2}^{2}.$
If $\|\Delta\|_{2}\neq 0$:
$\|\Delta\|_{2}\leq\frac{1}{\tilde{\mu}}\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}.$
We use the similar derivations we had before to bound the gradient differences. Let $B(n)=\|\hat{\beta}^{\,(true)}\|_{2}=O(\log n)$ on $\mathcal{E}$. Then
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq\left(2+O((\log n)(\log m))\right)\delta.$
Let $C_{G}(n)=(2+O((\log n)(\log m)))$. Then $\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq C_{G}(n)\delta$. Substituting:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}\leq\frac{C_{G}(n)}{\tilde{\mu}}\delta=\frac{8C_{G}(n)}{\mu}\delta.$
Let $C_{\Delta}(n)=\frac{8C_{G}(n)}{\mu}=O((\log n)(\log m))$. Then on $\mathcal{E}$:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}\leq C_{\Delta}(n)\delta=O(\delta(\log n)(\log m)).$
In summary, under event $\mathcal{E}$:
$\|\hat{\beta}-\beta^{*}\|_{2}$ $\leq\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}+\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}$
$\leq O\left(\max\left\{\delta(\log n)(\log m),\frac{\log n}{\sqrt{n}}\right\}\right).$
This holds with probability at least $1-c_{5}n^{-(k+1)}$. This completes the proof.
Before concluding the proof, we list a few side notes here.
  • If we know that all $x_{i}$’s are bounded and $\beta$ is also bounded, following the same derivations above would drop the $\log n$ terms involved because of the increasing norm. Also, we no longer need the polynomial probability from $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$. And the result error bound would be
    $\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max(\delta,\sqrt{\frac{\log n}{n}})$
    with probability at least $1-e^{-c_{5}n}$ as long as $\delta\gg 1/\sqrt{n}$.
  • If we have one offset covariate, the proof and the result still hold. The only difference is that we exclude on coordinate from $\beta$ in the estimation.
With the preparation of Lemma D.3, we are ready to prove the result for $\hat{Z}_{i}$’s.
Theorem 6.
Under the inner product latent space model model. Let the distribution $F$ satisfy the following two properties:
  1. 1.
    The support of $F$ is contained within a ball of radius $R$ for some $R>0$
  2. 2.
    Let $\Sigma=\mathbb{E}[ZZ^{\top}]$. There exists a constant $\mu>0$ such that:
    $\Sigma\succeq\mu I_{d}$
    where $I_{d}$ is the $d\times d$ identity matrix.
Moreover, on the hold-out data, suppose that with probability tending to 1, we can achieve
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}=o(1).$
Let $\hat{Z_{i}}$, $1\leq i\leq n$, be the estimated latent space vectors from our node-wide maximum likelihood estimation procedure (Algorithm 1). There exists positive constants $c,C$ and $C^{\prime}$ such that for sufficiently large $n$, we have
$\max_{1\leq i\leq n}\|{\hat{Z}_{i}-Z_{i}}\|\leq C\max\left\{\delta_{m},\frac{\sqrt{\log m+\log n}}{\sqrt{m}}\right\}$
with probability approaching 1, for $m\gg\log n$.
Proof of Theorem 6.
In the node-wise estimation procedure, for any fixed node $i$, we can treat $(\hat{Z}_{i},\hat{\alpha}_{i})$ as the $\hat{\beta}$ in a logistic regression problem with true covariates $Z_{j}$, $n+1\leq j\leq n+m$, and perturbed covariates $\hat{Z}_{j},n+1\leq j\leq n+m$. We apply Lemma D.3 to all $1\leq i\leq n$ to get the result. Notice that in handling the success probability, the event involving $\hat{Z}_{i}$, $n+1\leq i\leq n+m$, is indeed shared across all $1\leq i\leq n$, for which the probability control can be improved.
To apply Lemma D.3 to all $\hat{Z}_{i},1\leq i\leq n$, the only requirement is that the positive definite assumption on the population (expected) Hessian in Lemma D.3 holds uniformly for all $1\leq i\leq n$. To see this, let $\beta$ be an arbitrary latent vector from the distribution $F$. The population Hessian for the logistic loss is defined as the expectation of the single-point Hessian over the distribution of the covariates $x\sim F$:
$H(Z_{i})=\mathbb{E}_{Z_{j},A_{ij}}\left[\nabla^{2}\ell(Z_{i};Z_{j},A_{ij})\right]=\mathbb{E}\left[\sigma^{\prime}(Z_{i}^{\top}Z_{j})Z_{j}Z_{j}^{\top}\right]$ (23)
where $\sigma^{\prime}(z)=\frac{e^{z}}{(1+e^{z})^{2}}$ is the derivative of the standard sigmoid function.
To find a uniform lower bound for the smallest eigenvalue of $H(\beta)$, it is sufficient to find a $\mu>0$ such that for any unit vector $v\in\mathbb{R}^{d}$ (i.e., $\|v\|_{2}=1$), we have $v^{\top}H(Z_{i})v\geq\mu$, for any $Z_{i}$ in the domain.
By the Cauchy-Schwarz, we have
$|Z_{i}^{\top}Z_{j}|\leq\|Z_{i}\|_{2}\|Z_{j}\|_{2}\leq R^{2}$
The minimum value of $\sigma^{\prime}$ on the interval $R^{2},R^{2}]$ occurs at the endpoints $\sigma^{\prime}(R^{2})$. Let us define this positive constant as $w_{\min}:=\sigma^{\prime}(R^{2})>0$. Since $|Z_{i}^{\top}Z_{j}|\leq R^{2}$ for any draws, we have
$\sigma^{\prime}(\beta^{\top}x)\geq w_{\min}.$
From (23), we can write the quadratic form as:
$v^{\top}H(Z_{i})v$ $==\mathbb{E}\left[\sigma^{\prime}(Z_{i}^{\top}Z_{j})(v^{\top}Z_{j})^{2}\right]\geq w_{\min}\mathbb{E}\left[(v^{\top}Z_{j})^{2}\right]$
$w_{\min}v^{\top}\left(\mathbb{E}[Z_{j}Z_{j}^{\top}]\right)v.$ (24)
That means,
$v^{\top}H(Z_{i})v\geq w_{\min}(v^{\top}\Sigma v)$
for any $Z_{i}$ in the domain.
Therefore, as long as we have $\Sigma\succeq\mu I$ for some $\mu>0$, the requirement hold for all $Z_{i},i\in[n]$.
D.2 Consistency under the RDPG
Similar to the inner product model scenario, for the RDPG case, the etimation is essentially an ordinary least square (OLS) problem with measurement errors. Thus we first set up a generic lienar regression problem and study the crucial properties.
We consider i.i.d. data $\{(x_{i},y_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$, generated conditional on true predictors $x_{i}$ with probability $p_{i}=\mathbb{E}[y_{i}|x_{i}]=x_{i}^{\top}\beta^{*}$, where $\beta^{*}\in\mathbb{R}^{d}$ is the true parameter vector. The squared error loss for a single data point is $\ell_{sq}(\beta;x,y)=(y-x^{\top}\beta)^{2}$.
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta_{n}$. The average empirical losses based on true and perturbed data can hence be represented as:
$L(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-x_{i}^{\top}\beta)^{2}.$
$\widetilde{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\tilde{x}_{i}^{\top}\beta)^{2}.$
The population loss is $\mathcal{L}(\beta)=\mathbb{E}[(y-x^{\top}\beta)^{2}]$. The true parameter $\beta^{*}$ minimizes $\mathcal{L}(\beta)$.
The corresponding OLS estimators are:
$\hat{\beta}^{\,(true)}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}L(\beta).$
$\hat{\beta}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}\widetilde{L}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|_{2}$. We impose the following regularity conditions.
  1. (B1)
    (Model): $y_{i}\sim\text{Bernoulli}(x_{i}^{\top}\beta^{*})$ independently.
  2. (B2)
    (Boundedness): $\|x_{i}\|_{2}\leq R_{x}$ a.s. and $\|\beta^{*}\|_{2}\leq R_{\beta}$ for constants $R_{x},R_{\beta}>0$. We also assume $0\leq x_{i}^{\top}\beta^{*}\leq 1$ for all $i$.
  3. (B3)
    (Expected Design Matrix): $\Sigma_{x}=\mathbb{E}[x_{i}x_{i}^{\top}]\succeq\mu I$ for some fixed $\mu>0$. Let $\hat{\Sigma}_{x}=\frac{1}{n}\sum x_{i}x_{i}^{\top}$.
  4. (B4)
    (Perturbation): $\|\tilde{x}_{i}-x_{i}\|_{2}\leq\delta_{n}$, with $\delta_{n}\to 0$ as $n\to\infty$.
Lemma D.4.
Under Assumptions (B1)-(B4), there exist constants $C,c_{1},c_{2}>0$ (depending on $R_{x},R_{\beta},\mu,d$) such that for sufficiently large $n$:
$\operatorname{P}\left(\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\sqrt{\frac{d}{n}},\delta_{n}\right\}\right)\geq 1-c_{1}e^{-c_{2}n}.$
Proof.
Let $\Delta_{\text{total}}=\hat{\beta}-\beta^{*}$. We use the triangle inequality:
$\|\hat{\beta}-\beta^{*}\|_{2}\leq\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}+\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}.$
Let $\mathcal{E}$ denote a high-probability event (specified later) where concentration bounds hold.
Let $\epsilon_{i}=y_{i}-\mathbb{E}[y_{i}|x_{i}]=y_{i}-x_{i}^{\top}\beta^{*}$. $\mathbb{E}[\epsilon_{i}|x_{i}]=0$. The error of the ideal OLS estimator is given by:
$\hat{\beta}^{\,(true)}-\beta^{*}=\hat{\Sigma}_{x}^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}\right).$
Taking norms:
$\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}.$
We bound the two terms on the right. Let $\epsilon_{i}=y_{i}-\mathbb{E}[y_{i}|x_{i}]=y_{i}-x_{i}^{\top}\beta^{*}$. $\mathbb{E}[\epsilon_{i}|x_{i}]=0$. The error of the ideal OLS estimator is given by:
$\hat{\beta}^{\,(true)}-\beta^{*}=\hat{\Sigma}_{x}^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}\right),$
where $\hat{\Sigma}_{x}=\frac{1}{n}\sum x_{i}x_{i}^{\top}$. Taking norms:
$\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}.$
We bound the two terms on the right using concentration inequalities.
First, the term $\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}$ involves the sample covariance concentration for i.i.d random vectors. By standard concentration result (e.g., Vershynin [2018]), we have
$\operatorname{P}\left(\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\geq\mu/2\right)\leq 2d\exp\left(\frac{-n(\mu/2)^{2}}{8R_{x}^{4}}\right)=2d\exp\left(\frac{-n\mu^{2}}{32R_{x}^{4}}\right).$
Define $\mathcal{E}_{A}$ to be the event that $\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\leq\mu/2$. Then $P(\mathcal{E}_{A})\geq 1-2de^{-c^{\prime}n}$ for $c^{\prime}=\mu^{2}/(32R_{x}^{4})$. On event $\mathcal{E}_{A}$, using Weyl’s inequality:
$\lambda_{\min}(\hat{\Sigma}_{x})\geq\lambda_{\min}(\Sigma_{x})-\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\geq\mu-\mu/2=\mu/2,$
which leads to
$\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{x})}\leq\frac{2}{\mu}.$
For the second term, let $S=\frac{1}{n}\sum Z_{i}$ where $Z_{i}=x_{i}\epsilon_{i}$. As noted, $\mathbb{E}[Z_{i}]=0$ and $|\epsilon_{i}|\leq 1$. Thus, $\|Z_{i}\|_{2}=\|x_{i}\epsilon_{i}\|_{2}\leq\|x_{i}\|_{2}|\epsilon_{i}|\leq R_{x}$. The vectors $Z_{i}$ are bounded, independent, mean-zero random vectors. By Hoeffding inequality on bounded random vectors, we have
$\mathbb{P}(\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}\leq C_{S}\sqrt{\frac{d}{n}})\geq 1-c_{1}^{\prime}e^{-c_{2}^{\prime}n}.$
Define such an event as $\mathcal{E}_{B}$, we have
$P(\mathcal{E}_{B})\geq 1-c^{\prime}_{1}e^{-c^{\prime}_{2}n}.$ (25)
Let $\mathcal{E}_{true}=\mathcal{E}_{A}\cap\mathcal{E}_{B}$. Then $P(\mathcal{E}_{true})\geq 1-c^{\prime\prime}_{1}e^{-c^{\prime\prime}_{2}n}$. On event $\mathcal{E}_{true}$: $\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq C_{true}\sqrt{d/n}$ for constant $C_{true}>0$.
Next, we bound $\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}$.
Let $\Delta=\hat{\beta}-\hat{\beta}^{\,(true)}$. The first-order condition for $\hat{\beta}$ is $\nabla\widetilde{L}(\hat{\beta})=0$. The Hessian $\nabla^{2}\widetilde{L}(\beta)=2\hat{\Sigma}_{\tilde{x}}=\frac{2}{n}\sum\tilde{x}_{i}\tilde{x}_{i}^{\top}$ is constant w.r.t. $\beta$. Since $L$ is quadratic, we have
$0=\nabla\widetilde{L}(\hat{\beta})=\nabla\widetilde{L}(\hat{\beta}^{\,(true)})+2\hat{\Sigma}_{\tilde{x}}\Delta$
Substituting $\nabla L(\hat{\beta}^{\,(true)})=0$ into the equation above, we have
$\nabla\widetilde{L}(\hat{\beta}^{\,(true)})=\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)}).$
Assuming $\hat{\Sigma}_{\tilde{x}}$ is invertible (justified below), we get
$\Delta=-\frac{1}{2}\hat{\Sigma}_{\tilde{x}}^{-1}\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})$
and
$\|\Delta\|_{2}\leq\frac{1}{2}\left\|\hat{\Sigma}_{\tilde{x}}^{-1}\right\|_{\mathrm{op}}\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}.$
Again, we bound the two terms on the right.
For the first,
$\left\|\hat{\Sigma}_{\tilde{x}}-\hat{\Sigma}_{x}\right\|_{\mathrm{op}}\leq\frac{1}{n}\sum\left\|\tilde{x}_{i}\tilde{x}_{i}^{\top}-x_{i}x_{i}^{\top}\right\|_{\mathrm{op}}\leq(2R_{x}+\delta_{n})\delta_{n}.$
Note that $\delta_{n}\to 0$. On event $\mathcal{E}_{A}$ ($\lambda_{\min}(\hat{\Sigma}_{x})\geq\mu/2$) and for sufficiently large $n$, we get
$\lambda_{\min}(\hat{\Sigma}_{\tilde{x}})\geq\mu/4.$
Let $\mathcal{E}_{C}$ be this event ($P(\mathcal{E}_{C})\geq P(\mathcal{E}_{A})$).
For the second term of gradient difference: $\nabla(\widetilde{L}-L)(\beta)=\frac{1}{n}\sum T_{i}(\beta)$, where
$T_{i}(\beta)=(\tilde{x}_{i}^{\top}\beta-y_{i})\tilde{x}_{i}-(x_{i}^{\top}\beta-y_{i})x_{i}=\tilde{x}_{i}^{\top}\beta(\tilde{x}_{i}-x_{i})+(\tilde{x}_{i}-x_{i})^{\top}\beta x_{i}-y_{i}(\tilde{x}_{i}-x_{i}).$
Thus, on event $\mathcal{E}_{true}$, we have
$\|T_{i}(\hat{\beta}^{\,(true)})\|_{2}\leq(|\tilde{x}_{i}^{\top}\hat{\beta}^{\,(true)}|+|y_{i}|)\delta_{n}+|(\tilde{x}_{i}-x_{i})^{\top}\hat{\beta}^{\,(true)}|\|x_{i}\|_{2}=O(\delta_{n}).$
Therefore, we have
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq\frac{1}{n}\sum\|T_{i}(\hat{\beta}^{\,(true)})\|_{2}\leq O(\delta_{n}).$
Substituting these bounds into the inequality for $\|\Delta\|_{2}$, we get on event $\mathcal{E}=\mathcal{E}_{true}\cap\mathcal{E}_{C}$:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}=\|\Delta\|_{2}\leq\frac{1}{2}\left(\frac{4}{\mu}\right)(C_{G}\delta_{n})=\frac{2C_{G}}{\mu}\delta_{n}.$
Finally, let $\mathcal{E}=\mathcal{E}_{true}\cap\mathcal{E}_{C}$. Taking $d$ as a constant, using the two bounds, we get
$\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\sqrt{\frac{1}{n}},\delta_{n}\right\}$
with probability at least $1-c_{1}e^{-c_{2}n}$.
Similarly as in the inner product latent space model, for RDPG, by Lemma D.4, we get the following result.
Theorem 7.
Under the random dot product graph model, assume that the latent random vector $Z_{i}$’s are bounded. Suppose $\mathbb{E}[Z_{i}Z_{i}^{T}]\succeq\mu I_{d}$ for some $\mu>0$. Moreover, on the hold-out data, suppose that with probability tending to 1, we can achieve
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}=o(1).$
Let $\hat{Z_{i}},1\leq i\leq n$ be the estimated latent space vectors from our node-wide maximum likelihood estimation procedure (Algorithm 1). If $m\gg\log{n}$, there exists positive constants $c,C$ and $C^{\prime}$ such that for sufficiently large $n$, we have
$\max_{1\leq i\leq n}\|{\hat{Z}_{i}-Z_{i}}\|\leq C\max\left\{\delta_{m},\frac{1}{\sqrt{m}}\right\}$
with probability approaching 1 as $n,m\to\infty$.
Proof.
The proof is simply applying Lemma D.4 for each estimator $\hat{Z}_{i},1\leq i\leq n$, then taking the union bound. ∎
D.3 Corollary 1
Proof of Corollary 1.
The only thing we need to check is Assumption A4. Note that the latent vectors in both models are not identifiable up to an orthogonal transformation. Since we have already match the orientations in our algorithm, we do not have to worry about orthogonal transformation anymore.
For the inner product latent space model, Theorem 2.1 of Li et al. [2023a] shows that
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=O_{P}(\frac{1}{m^{1/2-c^{\prime}}})$
for some constant $c^{\prime}<1/2$.
For the RDPG model or gRDPG, Theorem 1 in Rubin-Delanchy et al. [2022] already gives that
$\max_{n+1\leq i\leq n+m}\|{\hat{Z}_{i}-Z_{i}}\|=O_{P}(\frac{(\log m)^{c^{\prime}}}{\sqrt{m}})$
for some constant $c^{\prime}>0$.
By applying Theorem 4, with $\delta_{m}=\frac{1}{m^{1/2-c^{\prime}}}$ and $\frac{(\log m)^{c^{\prime}}}{\sqrt{m}}$, respectively. We can see that Assumption A4 holds.
Appendix E Proof of Theorem 5 (Convergence of Network Moments)
The conditional expectation of network moments, given the latent vectors, are essentially U-statistics. Hence, our proof uses the conditional expectation as an intermediate quantity. To structure the proof, we begin by analyzing the distribution of a U-statistic with respect to the approximation of an empirical CDF.
Let $X_{1},X_{2},\dotsc,X_{n}$ be independent and identically distributed (i.i.d.) random variables from a continuous distribution $F$ on ${\bf R}^{d}$. Suppose we have approximate observations $\hat{X}_{1},\hat{X}_{2},\dotsc,\hat{X}_{n}$, which may be dependent, but whose empirical distribution function $\hat{F}_{n}$ converges uniformly in probability to $F$. We are interested in understanding the relationship between U-statistics computed from the exact data and those computed from the approximate data.
Let $h:({\bf R}^{d})^{r}\to{\bf R}$ be a bounded, Lipchitz continuous, and symmetric kernel function. We define the U-statistics:
$U_{n}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\dotsb<i_{r}\leq n}h(X_{i_{1}},\dotsc,X_{i_{r}}),$
$\hat{U}_{n}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\dotsb<i_{r}\leq n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}}).$
Let us first introduce some notation and preliminary results that will be used in the proof.
Empirical Measures on $r$-Tuples:
- The empirical measure over combinations without replacement from $X_{i}$ is:
$F_{n}^{[r]}=\frac{1}{\binom{n}{r}}\sum_{(i_{1},\dots,i_{m})\in I_{n}}\delta_{(X_{i_{1}},\dots,X_{i_{r}})},$
where $\delta_{(X_{i_{1}},\dots,X_{i_{r}})}$ is the Dirac measure at the point $(X_{i_{1}},\dots,X_{i_{r}})$ in $({\bf R}^{d})^{r}$.
- Similarly, for $\hat{X}_{i}$:
$\hat{F}_{n}^{[r]}=\frac{1}{\binom{n}{r}}\sum_{(i_{1},\dots,i_{m})\in I_{n}}\delta_{(\hat{X}_{i_{1}},\dots,\hat{X}_{i_{r}})}.$
Product Measures:
- The product measure of $F$ is $F^{\otimes r}$, defined on $({\bf R}^{d})^{r}$ as:
$F^{\otimes r}(A)=\int_{A}dF(x_{1})\cdots dF(x_{r}),$
for measurable sets $A\subset({\bf R}^{d})^{r}$.
Lemma E.1.
Let $F$ be a CDF on ${\bf R}^{d}$ and $\hat{F}_{n}$ be empirical distribution functions on ${\bf R}^{d}$ defined as:
$\hat{F}_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\hat{X}_{i}}(x),$
where:
  • $\hat{X}_{1},\hat{X}_{2},\dotsc,\hat{X}_{n}$ are random variables (not necessarily independent) such that:
    $\sup_{x\in{\bf R}^{d}}|\hat{F}_{n}(x)-F(x)|\xrightarrow{P}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $({\bf R}^{d})^{r}$, that is:
$\hat{F}_{n}^{\otimes r}$ $\xrightarrow{\text{w.P.}}F^{\otimes r},$
where $\xrightarrow{\text{w.P.}}$ denotes weak convergence in probability.
Lemma E.2.
Let $F_{n}$ and $\hat{F}_{n}$ be empirical distribution functions on ${\bf R}^{d}$ defined as:
$F_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}(x),\quad\hat{F}_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\delta_{\hat{X}_{i}}(x),$
where:
  • $X_{1},X_{2},\dotsc,X_{n}$ are independent and identically distributed (i.i.d.) random variables with distribution $F$ on ${\bf R}^{d}$.
  • $\hat{X}_{1},\hat{X}_{2},\dotsc,\hat{X}_{n}$ are random variables (not necessarily independent) such that:
    $\sup_{x\in{\bf R}^{d}}|\hat{F}_{n}(x)-F(x)|\xrightarrow{P}0\quad\text{as }n\to\infty.$
Then, for each $m\geq 1$, the product measures $F_{n}^{\otimes r}$ and $\hat{F}_{n}^{\otimes r}$ converge weakly to $F^{\otimes r}$ on $({\bf R}^{d})^{r}$, that is:
$F_{n}^{\otimes r}$ $\xrightarrow{\text{w.a.s.}}F^{\otimes r},$
$\hat{F}_{n}^{\otimes r}$ $\xrightarrow{\text{w.P.}}F^{\otimes r},$
where $\xrightarrow{\text{w.a.s.}}$ denotes weak convergence almost surely, and $\xrightarrow{\text{w.P.}}$ denotes weak convergence in probability.
Proof.
Since $X_{1},X_{2},\dotsc,X_{n}$ are i.i.d. with distribution $F$, by the Glivenko–Cantelli theorem, we have:
$\sup_{x\in{\bf R}^{d}}|F_{n}(x)-F(x)|\xrightarrow{\text{a.s.}}0\quad\text{as }n\to\infty.$
This implies that for any bounded, continuous function $g:{\bf R}^{d}\to{\bf R}$,
$\int g\,dF_{n}\xrightarrow{\text{a.s.}}\int g\,dF.$
Now, consider any bounded, continuous function $h:({\bf R}^{d})^{r}\to{\bf R}$. We need to show that:
$\int h\,dF_{n}^{\otimes r}\xrightarrow{\text{a.s.}}\int h\,dF^{\otimes r}.$
Since $F_{n}^{\otimes r}=F_{n}\otimes F_{n}\otimes\cdots\otimes F_{n}$ ($r$ times), we have:
$\int h(x_{1},x_{2},\dotsc,x_{r})\,dF_{n}^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,dF_{n}(x_{1})\cdots dF_{n}(x_{r}).$
Similarly, for $F^{\otimes r}$:
$\int h(x_{1},x_{2},\dotsc,x_{r})\,dF^{\otimes r}(x_{1},x_{2},\dotsc,x_{r})=\int\cdots\int h(x_{1},x_{2},\dotsc,x_{r})\,dF(x_{1})\cdots dF(x_{r}).$
Since $F_{n}\xrightarrow{\text{a.s.}}F$ uniformly, for each $j=1,2,\dotsc,m$, we have:
$\int g_{j}\,dF_{n}\xrightarrow{\text{a.s.}}\int g_{j}\,dF,$
for any bounded, continuous function $g_{j}:{\bf R}^{d}\to{\bf R}$.
By the weak convergence of product measures [Billingsley, 2013], it follows that:
$\int h\,dF_{n}^{\otimes r}\xrightarrow{\text{a.s.}}\int h\,dF^{\otimes r}.$
The convergence of $\hat{F}_{n}^{\otimes r}$ to $F^{\otimes r}$ can be proved with the same reasoning.
Lemma E.3.
Under the above assumptions of Lemma E.2, we have
$\hat{U}_{n}-U_{n}\xrightarrow{P}0\quad\text{as }n\to\infty.$
Proof.
We aim to show that $\hat{U}_{n}-U_{n}\xrightarrow{P}0$ as $n\to\infty$. We will use
$\theta=\int h\,dF^{\otimes r}=\mathbb{E}[h(X_{1},\dotsc,x_{r})]$
as the intermediate quantity for the convergence. To directly characterize the approximation error between the U-statistic and the integral over the product measure, we consider the difference between sampling without replacement (as in $U_{n},\hat{U}_{n}$) and sampling with replacement (as in $\int h\,d{F}_{n}^{\otimes r},\int h\,d\hat{F}_{n}^{\otimes r}$).
It is well-known that the U-statistic $U_{n}$ can be related to an integral over the product measure $F_{n}^{\otimes r}$, but with a subtle difference in normalization. Let
$S_{\text{distinct}}=\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}})$
be the sum over all distinct $r$-tuples from $\hat{X}_{i}$’s. Also define
$S_{\text{all}}=\sum_{i_{1}=1}^{n}\cdots\sum_{i_{r}=1}^{n}h(\hat{X}_{i_{1}},\dotsc,\hat{X}_{i_{r}}),$
the sum over all $r$-tuples. With these definitions, we have
$\hat{U}_{n}=\frac{S_{\text{distinct}}}{\binom{n}{r}},\quad\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{all}}}{n^{r}}.$
Now, taking the difference of the two terms as
$\varepsilon_{n}=\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{distinct}}}{\binom{n}{r}}-\frac{S_{\text{all}}}{n^{r}}.$
If we consider only tuples with distinct indices, there are $\binom{n}{r}r!$ such tuples. Therefore, we have
$S_{\text{all}}=r!S_{\text{distinct}}+R_{n},$
where $R_{n}$ accounts for the contributions from tuples that include repeated indices. The number of tuples with repeated indices is at most: $n^{r}-\binom{n}{r}r!.$ Note that from the definitions of combinatorial terms, we have
$\binom{n}{r}r!=n^{r}\left(1-O\left(\frac{1}{n}\right)\right),\quad n^{r}-\binom{n}{r}r!=n^{r}O\left(\frac{1}{n}\right)=O(n^{r-1}).$
From this, we get $\hat{U}_{n}=\frac{S_{\text{distinct}}}{\binom{n}{r}}=\frac{S_{\text{all}}-R_{n}}{r!\binom{n}{r}}$. Compare this with $\int h\,d\hat{F}_{n}^{\otimes r}$, we have
$\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=\frac{S_{\text{all}}-R_{n}}{r!\binom{n}{r}}-\frac{S_{\text{all}}}{n^{r}}=S_{\text{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)-\frac{R_{n}}{r!\binom{n}{r}}$.
With the boundedness assumption of $h$, we have $S_{\text{all}}\leq Mn^{r}$, thus:
$S_{\text{all}}\left(\frac{1}{r!\binom{n}{r}}-\frac{1}{n^{r}}\right)=Mn^{r}O\left(\frac{1}{n^{r+1}}\right)=O\left(\frac{1}{n}\right).$
Similarly, we have $|R_{n}|=MO(n^{r-1})$ and $\frac{1}{r!\binom{n}{r}}=O\left(\frac{1}{n}\right).$ Combining these leads to
$|\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}|\leq O\left(\frac{1}{n}\right)+O\left(\frac{1}{n}\right)=O\left(\frac{1}{n}\right).$
Thus $\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}=O(1/n)\to 0$ as $n\to\infty$.
Using the same reasoning, define $\varepsilon_{n}={U}_{n}-\int h\,d{F}_{n}^{\otimes r}$ and we also have $|{\varepsilon}_{n}|=O(n^{-1}).$
Now we have the decomposition
$\hat{U}_{n}-U_{n}$ $=\left(\hat{U}_{n}-\int h\,d\hat{F}_{n}^{\otimes r}\right)+\left(\int h\,d\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,dF_{n}^{\otimes r}\right)+\left(\int h\,dF_{n}^{\otimes r}-U_{n}\right)$
$=\hat{\varepsilon}_{n}+\left(\int h\,d\hat{F}_{n}^{\otimes r}-\theta\right)+\left(\theta-\int h\,dF_{n}^{\otimes r}\right)-\varepsilon_{n}.$
Therefore,
$|\hat{U}_{n}-U_{n}|\leq|\hat{\varepsilon}_{n}|+\left|\int h\,d\hat{F}_{n}^{\otimes r}-\int h\,dF^{\otimes r}\right|+\left|\int h\,dF^{\otimes r}-\int h\,dF_{n}^{\otimes r}\right|+|\varepsilon_{n}|.$
Because of the boundedness and Lipschitz continuity of $h$, by the previous results of $\varepsilon$ and $\hat{\varepsilon}$, as well as Lemma E.2, we have
$\hat{U}_{n}-U_{n}\xrightarrow{P}0\quad\text{as }n\to\infty.$
Next, we introduce the property that, conditioning on the latent spaces, the network moments concentrate around its conditional expectation.
Lemma E.4 (Conditional concentration of Network Moments).
Consider a random (simple, undirected) graph on $n$ vertices with adjacency matrix $A=(A_{ij})$, where $1\leq i<j\leq n$, and $A_{ij}\sim\text{Bernoulli}(P_{ij})$ are independent random variables. Let $m\geq 2$ be fixed and consider a bounded function $h$ on the induced subgraph of $m$ distinct vertices with $|h|\leq 1$. Define
$X(A)=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n}h(A_{[i_{1},i_{2},\ldots,i_{r}]}),$
where $A_{[i_{1},i_{2},\ldots,i_{r}]}$ denotes the induced subgraph on vertices $\{i_{1},\ldots,i_{r}\}$. For every $\varepsilon>0$, there exist positive constants $c(r)$ (not depending on $n$) such that
$\mathbb{P}\bigl{(}|X-\mathbb{E}[X]|\geq\varepsilon\bigr{)}\leq 2\exp(-c(r)\varepsilon^{2}n^{2}),$
and in particular,
$X\xrightarrow{p}\mathbb{E}[X]\quad\text{as }n\to\infty.$
Proof.
Consider the set of edges $\{A_{ij}:1\leq i<j\leq n\}$. There are $\binom{n}{2}$ such edges, each an independent Bernoulli random variable. The random variable $X$ is a function of these $\binom{n}{2}$ independent variables. We will apply McDiarmid’s inequality (Lemma F.2), a bounded-differences inequality, to show that $X$ is sharply concentrated around its expectation.
First, we need to bound the effect of changing one edge on the value of $X$. Fix an edge $(u,v)$ with $1\leq u<v\leq n$. Consider two graphs $A$ and $A^{\prime}$ that differ only on the edge $(u,v)$: $A^{\prime}=A$ on all entries, except for $(u,v)$, where $A^{\prime}_{uv}=1-A_{uv}$. We have
$X(A)=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(A[i_{1},\ldots,i_{r}]),\quad X(A^{\prime})=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(A^{\prime}[i_{1},\ldots,i_{r}]).$
The only $r$-tuples of vertices that can be affected by the change in the edge $(u,v)$ are those that contain both $u$ and $v$. The number of such subsets is $\binom{n-2}{r-2}$, since we choose the remaining m2m-2 vertices from the n2n-2 other vertices.
For each affected $r$-tuple, the value of $h$ can change by at most $2$ in absolute value (since $|h|\leq 1$). Therefore, the change in the numerator of $X$ is at most $2\binom{n-2}{r-2}$. The change in $X$ when flipping a single edge $(u,v)$ is bounded by
$\frac{2\binom{n-2}{r-2}}{\binom{n}{r}}$.
We use the combinatorial identity:
$\frac{\binom{n-2}{r-2}}{\binom{n}{r}}=\frac{r(r-1)}{n(n-1)}.$
Hence, the maximum change in $X$ due to flipping one edge is
$\Delta:=\frac{2r(r-1)}{n(n-1)}.$
As $n\to\infty$, $\Delta\approx\frac{2r(r-1)}{n^{2}}$, which vanishes.
Now we apply McDiarmid’s inequality (Lemma F.2). There are $M=\binom{n}{2}$ edges. Each edge affects $X$ by at most $\Delta$. Hence,
$\sum_{j=1}^{r}\Delta_{j}^{2}\leq M\Delta^{2}=\binom{n}{2}\left(\frac{2r(r-1)}{n(n-1)}\right)^{2}\leq\frac{C^{\prime}(r)}{n^{2}}$
for some constant $C^{\prime}(r)$ that depends only on $r$.
By McDiarmid’s inequality, for any $\varepsilon>0$,
$\mathbb{P}(|X-\mathbb{E}[X]|\geq\varepsilon)\leq 2\exp\left(-\frac{2\varepsilon^{2}}{\sum_{j=1}^{r}\Delta_{j}^{2}}\right).$
Since $\sum_{j=1}^{r}\Delta_{j}^{2}=O(1/n^{2})$, we have
$\mathbb{P}(|X-\mathbb{E}[X]|\geq\varepsilon)\leq 2\exp(-c(r)\varepsilon^{2}n^{2})$
for some constant $c(r)>0$. This shows exponential concentration of $X$ around its expectation as $n$ grows. In particular, this indicates that
$X\xrightarrow{p}\mathbb{E}[X]\quad\text{as }n\to\infty.$
Now combining Lemma E.3 and Lemma E.4, we can get the major claim about the network moments.
Proof of Theorem 5.
[…]
$U_{n}=\mathbb{E}\big{[}X_{R}(A)|\{Z_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}\mathbb{E}\big{[}{\mathbb{I}}(A_{[i_{1},\cdots,i_{r}]}\cong R)|\{Z_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(Z_{i_{1}},\cdots,Z_{i_{r}})$
and
$\hat{U}_{n}=\mathbb{E}\big{[}X_{R}(A)|\{\hat{Z}_{i}\}\big{]}=\frac{1}{\binom{n}{r}}\sum_{1\leq i_{1}<\cdots<i_{r}\leq n}h(\hat{Z}_{i_{1}},\cdots,\hat{Z}_{i_{r}}).$
We have the following decomposition
$).\displaystyle X_{R}(A)-X_{R}(\tilde{A})=\big{(}X_{R}(A)-U_{n}\big{)}+\big{(}U_{n}-\hat{U}_{n}\big{)}+\big{(}\hat{U}_{n}-X_{R}(\tilde{A}$$\big{)}$.
Conditioning on $\{Z_{i}\}$ and $\{\hat{Z}_{i}\}$, respectively, the first and third term vanish in probability from Lemma E.4. By Theorem 3, we know that the empirical CDF of $\tilde{Z}_{i}$’s satisfies the requirement of Lemma E.2. And then we call Lemma E.3, and the second term also vanishes in probability. ∎
Appendix F Supporting lemmas
Lemma F.1 (Telescoping Inequality).
Let $m\geq 1$ be an integer, and let $\{a_{i}\}_{i=1}^{r}$ and $\{b_{i}\}_{i=1}^{r}$ be two sequences of real numbers satisfying $0\leq a_{i},b_{i}\leq 1$ for all $i=1,2,\dots,m$. Then, the following inequality holds:
$\left|\prod_{i=1}^{r}a_{i}-\prod_{i=1}^{r}b_{i}\right|\leq\sum_{k=1}^{r}\left|a_{k}-b_{k}\right|.$
Lemma F.2 (McDiarmid’s Inequality).
Let $Y_{1},\ldots,Y_{M}$ be independent random variables taking values in arbitrary sets, and let $f(y_{1},\ldots,y_{M})$ be a function such that for all $j$ and for all $y_{1},\ldots,y_{M},y_{j}^{\prime}$, we have
$|f(y_{1},\ldots,y_{M})-f(y_{1},\ldots,y_{j-1},y_{j}^{\prime},y_{j+1},\ldots,y_{M})|\leq c_{j},$
for some constants $c_{j}$. Define $X=f(Y_{1},\ldots,Y_{M})$ and $\mu=\mathbb{E}[X]$. Then for any $\varepsilon>0$,
$\mathbb{P}(|X-\mu|\geq\varepsilon)\leq 2\exp\left(-\frac{2\varepsilon^{2}}{\sum_{j=1}^{r}c_{j}^{2}}\right).$
Lemma F.3 (Lemma 2 of Yin et al. [2006]).
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ smooth, injective on $B_{\delta_{1}}(x^{*})$. Let $\phi(x^{*})=y^{*}$. If for $\rho,\delta_{1}>0$, $\min_{\|{x-x^{*}}\|=\delta_{1}}\|{\phi(x)-y^{*}}\|\geq\rho$. Then $\forall y$ with $\|{y-y^{*}}\|\leq\rho$, $\exists x\in B_{\delta_{1}}(x^{*})$ s.t. $\phi(x)=y$.
Appendix G Additional evaluation on CommunityFitNet
The CommunitFitNet is a collection of network data sets introduced in Ghasemian et al. [2019], which is also used in Ghasemian et al. [2020] and Li and Le [2023]. In addition to the real data analyses in the main text, we also conducted a comprehensive evaluation of our method on this data set. In particular, since the privacy problem is more meaningful in the context of social networks, we focus on the 107 social networks from this collection with size larger than 200. We follow the same procedure of releasing the network of only half of nodes while using the other half as the hold-out data set for model selection and model fitting. The results are evaluated using the same five metrics as we used in the main text.
In Figure 4, we show the scatterplot of the Wasserstein distances between the privatized network and the true network for the distribution of local statistics. Both our method and the Laplace mechanism are evaluated for pairwise comparison. In particular, it can be seen that in all these networks, our method gives a better perservation of network properties (i.e., a much smaller Wasserstein distances across all datasets) than the naive Laplace method.
(a) Node degree
(b) Number of V-shapes
(c) Number of Triangles
(d) Eigen-centrality
(e) Harmonic closeness
Figure 4: The Wasserstein distances between the privatized distributions of given local statistics and those in the true network on 107 social networks from CommunityFitNet.